### v3

# appendixReferences

Abstract
We give an approach for characterizing interference by lower bounding the number of units whose outcome depends on certain groups of treated individuals, such as depending on the treatment of others, or others who are at least a certain distance away. The approach is applicable to randomized experiments with binary-valued outcomes. Asymptotically conservative point estimates and one-sided confidence intervals may be constructed with no assumptions beyond the known randomization design, allowing the approach to be used when interference is poorly understood, or when an observed network might only be a crude proxy for the underlying social mechanisms. Point estimates are equal to Hájek-weighted comparisons of units with differing levels of treatment exposure. Empirically, we find that the width of our interval estimates is competitive with (and often smaller than) those of the EATE, an assumption-lean treatment effect, suggesting that the proposed estimands may be intrinsically easier to estimate than treatment effects.
1 Introduction
There is growing interest in the analysis of randomized experiments where participants may affect each other (“interference between units”), for example by transmission of disease, sharing of information, peer influence, or economic competition. Such mechanisms often play fundamental roles in our understanding of social behaviors, and their empirical characterization may be of scientific interest. In this work, we give a general approach for characterizing the prevalence of treatment effects and interference, which can be used to evaluate whether interference-causing mechanisms are widely prevalent in an experimental setting. The approach is applicable to experiments with binary-valued outcomes. It can be used to explore questions such as
  1. 1.
    How many units are affected by any treatment (including their own)?
  2. 2.
    How many units are affected by the treatment of others? of distant others?
  3. 3.
    How many units are affected by the treatment of others, when their own treatment satisfies some condition?
For each of these questions, our approach gives conservative point estimates and one-sided confidence intervals, both of which lower bound the true value. Under reasonable experiment designs, our point estimates will converge to a lower bound on the true value, while the one-sided interval will have coverage probability converging to at least the desired level. These consistency and coverage properties are agnostic, in that they require no assumptions or restrictions on the nature of the interference, and rely only on the known experiment design. As a result, our estimates will remain valid even if an observed social network is only a crude proxy for the actual social mechanisms.
Existing work in this area produced very loose (though valid) lower bounds (Choi, , 2023). Our approach is significantly tighter, and may be more practical as a result. After Hájek normalization, our point estimates are equal to Hájek-weighted contrasts, such as comparisons of treated versus untreated, or comparisons of different levels of indirect exposure. Under stronger assumptions, contrasts arise naturally as estimates of treatment effects. Our results indicate that without assumptions, these contrasts may sometimes be re-interpreted more weakly as lower bounds on the number of units who are affected by the treatments in a particular manner, albeit with new confidence intervals that use a novel bound for the estimator variance. We find empirically that the uncertainty of our estimates (i.e., the interval widths) is competitive with, and often better than, that of the EATE, an assumption-lean treatment effect.
Our approach does not replace estimation of treatment effects, but may serve as a useful complement. In the presence of interference, estimates of treatment effects may come with considerable caveats or limitations – for example by requiring strong assumptions, or by estimating marginal effects rather than total ones. As a result, the usage of these estimates may require caution, judgement, and a reliance on prior beliefs. Agnostic information about the underlying system may help by informing and challenging these beliefs. If the decision maker is confronted by the results of an agnostic analysis, even one that does not directly answer the policy question of interest, then their beliefs may require recalibration in light of the findings, which in turn may affect downstream decisions. On the other hand, an assumption-laden analysis may carry less weight if caution is required to interpret its findings.
2 Related Work
Our work is closely related to hypothesis testing in the presence of interference, which has been studied in Aronow, (2012); Athey et al., (2018); Pouget-Abadie et al., (2019); Basse et al., (2019); Han et al., (2022); Puelz et al., (2022); Hoshino and Yanagi, (2023), for testing of null effects and also of hypotheses regarding interference structures, such as the null of no interference between units. In these settings, p-values generally cannot be inverted into confidence intervals for a treatment effect unless additional assumptions are added, and the connection between testing and estimation under interference is not fully understood. Our work complements this literature, as we invert tests of interference specifications to construct one-sided intervals without additional assumptions – however not for a treatment effect, but rather for the number of units who violate the specification.
The estimation of treatment effects in the presence of interference is an active area of research (Sävje et al., , 2021; Li et al., , 2019; Hu et al., , 2021; Li and Wager, , 2022; Leung, , 2022; Ogburn et al., , 2022; Harshaw et al., , 2022; Park and Kang, , 2023; Gao and Ding, , 2023). In many of these works, the target parameter describes the effect of a marginal departure from the experiment design. For such estimands, consistent estimates can generally be constructed in assumption-lean regimes, although additional assumptions are needed for coverage of interval estimates. Harshaw et al., (2021) focuses on the problem of variance estimation. Estimation of total effects (as opposed to marginal effects) assuming low order interactions between unit treatments is considered in Cortez et al., (2022) under staggered rollout, and in Cortez-Rodriguez et al., (2023) for known exposure neighborhoods. Blake and Coey, (2014) describe a particular form of interference arising in e-commerce settings, studied in Farias et al., (2022); Dhaouadi et al., (2023); Johari et al., (2024). Banerjee et al., (2024) discusses an empirical campaign for information sharing in which interference may be more subtle and complex than is captured by models of local exposure such as infection.
3 Methodology (basic case)
Our approach can be adapted to a variety of estimands that characterize interference, including the questions that were presented in the introduction. In this section, we give a basic example, in which the estimand is the number of units who are affected by any treatment. This quantity, which we denote by $\tau^{\textup{basic}}$, includes units who are affected by their own treatment as well as those who are affected by the treatment of others. Sections 3.1 through 3.6 respectively present the generative model, estimand, point estimation, Hájek normalization, inference, and computation. In Section 4, the approach is then generalized to cover a wider range of target parameters.
3.1 Preliminaries
We consider an experiment on $N$ units, with $X=(X_{1},\ldots,X_{N})$ denoting the binary-valued treatment of each unit, and $Y=(Y_{1},\ldots,Y_{N})$ denoting their binary outcomes. We allow for arbitrary interference, so that the outcome $Y_{i}$ of unit $i$ may potentially depend on all $N$ treatments,
$Y_{i}=f_{i}(X_{1},\ldots,X_{N}),\qquad i\in[N]$ (1)
where the potential outcome mappings $\{f_{i}\}_{i=1}^{N}$ may be arbitrary and unknown.
Given an observed network, we will use $\eta_{i}$ to denote the set consisting of unit $i$ and $i$’s “close” neighbors, and $\eta_{i}^{(2)}$ to denote “non-close” neighbors of unit $i$. For example, $\eta_{i}$ might denote unit $i$ and those units whose shortest path distance to $i$ is 1, while $\eta_{i}^{(2)}$ might denote those whose shortest path distance is 2. We will let $W_{i}$ denote the thresholded number of treated close neighbors for unit $i$,
$W_{i}=1\{\textup{if at least $t_{i}$ units in $\eta_{i}$, excluding unit $i$, are treated}\},$ (2)
and $W_{i}^{(2)}$ will denote the thresholded number of treated non-close neighbors,
$W_{i}^{(2)}=1\{\textup{ if at least $t_{i}^{(2)}$ units in $\eta_{i}^{(2)}$ are treated}\}$ (3)
for some choice of thresholds $t_{i}$ and $t_{i}^{(2)}$.
Given integer $n>0$, we use $[n]$ to denote the set $\{1,\ldots,n\}$. Given a set $S\subseteq[N]$, we use $X_{S}$ denote the subvector of $X$ corresponding to the members of $S$.
3.2 Estimand
Let $\phi^{*}=(\phi_{1}^{*},\ldots,\phi_{N}^{*})$ indicate whether each unit is unaffected by treatment,
$\phi_{i}^{*}=\begin{cases}1&\textup{if $f_{i}(X)$ is constant in $X$}\\ 0&\textup{otherwise}\end{cases}$
Equivalently, we may say that $Y_{i}$ is deterministic if $\phi_{i}^{*}=1$. Let $\Phi$ denote the number of units unaffected by treatment,
$\Phi=\sum_{i=1}^{N}\phi_{i}^{*}$
and let $\tau^{\textup{basic}}$ denote the complement of $\Phi$,
$\tau^{\textup{basic}}=N-\Phi$
which equals the total number of units who are affected by $X$. The target parameter of interest is $\tau^{\textup{basic}}$.
3.3 Point Estimation
To estimate $\tau^{\textup{basic}}$, we first consider two Horvitz-Thompson estimators for $\Phi$, which are not computable as they require knowledge of $\phi^{*}$:
$\hat{\Phi}_{1}$ $=\sum_{i:\phi_{i}^{*}=1}\frac{1\{X_{i}=Y_{i}\}}{{P}(X_{i}=Y_{i})}$ $\hat{\Phi}_{2}$ $=\sum_{i:\phi_{i}^{*}=1}\frac{1\{X_{i}\neq Y_{i}\}}{{P}(X_{i}\neq Y_{i})}$ (4)
To give a single formula for both estimators we may define $Y_{ki}$ for $k\in[2]$ and $i\in[N]$ as the estimator-specific outcome
$Y_{ki}=\begin{cases}Y_{i}&\textup{ if $k=1$}\\ (1-Y_{i})&\textup{ if $k=2$}\end{cases}$
and rewrite (4) as
$\hat{\Phi}_{k}$ $=\sum_{i:\phi_{i}^{*}=1}\frac{1\{X_{i}=Y_{ki}\}}{P(X_{i}=Y_{ki})}$ (5)
To evaluate the probability $P(X_{i}=Y_{ki})$, we observe that if $\phi_{i}^{*}=1$, then $Y_{ki}$ is deterministic, implying that
$P(X_{i}=Y_{ki})=Y_{ki}\cdot P(X_{i}=1)+(1-Y_{ki})\cdot P(X_{i}=0),\qquad\textup{ if $\phi_{i}^{*}=1$.}$ (6)
It follows that (5) is an unbiased estimator of $\Phi$ provided that $P(X_{i}=0)$ and $P(X_{i}=1)$ are non-zero for all $i\in[N]$.
Since only deterministic outcomes are involved, the estimators will often exhibit simple behavior. For example, if treatment is assigned by independent Bernoulli randomization, then $\hat{\Phi}_{k}$ is a sum of independent variables, regardless of any interference that may be present. Similarly, if the dependencies between the unit treatments are bounded, then $\hat{\Phi}_{k}$ is a sum of variables whose dependencies will be similarly bounded. For this reason, under a variety of designs we may expect $\hat{\Phi}_{k}$ to concentrate at its expectation and be asymptotically normal (while also being uncomputable as $\phi^{*}$ is unknown).
Let $\hat{\Phi}_{k}(\phi)$ denote $\hat{\Phi}_{k}$, evaluated under the hypothesis that $\phi^{*}=\phi$. To give a concrete expression, we may write
$\hat{\Phi}_{k}(\phi)=\sum_{i:\phi_{i}=1}\frac{1\{X_{i}=Y_{ki}\}}{Y_{ki}\cdot P(X_{i}=1)+(1-Y_{ki})\cdot P(X_{i}=0)}$ (7)
where the denominator term is due to application of (6).
To conservatively estimate $\tau^{\textup{basic}}=N-\Phi$, we will plug in for $\Phi$ the smaller of $\hat{\Phi}_{1}(\phi)$ and $\hat{\Phi}_{2}(\phi)$ after maximizing over $\phi\in\{0,1\}^{N}$, resulting in the estimator
$\hat{\tau}=N-\min\Big{(}\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{1}(\phi)\ ,\ \max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{2}(\phi)\Big{)}$ (8)
The main properties of $\hat{\tau}$ are the following:
  1. 1.
    If $\hat{\Phi}_{1}$ and $\hat{\Phi}_{2}$ are both consistent for $\Phi$, then $\min(\hat{\Phi}_{1},\hat{\Phi}_{2})$ will also be consistent for $\Phi$, and $\hat{\tau}$ will asymptotically lower bound $\tau^{\textup{basic}}$, as stated by Theorem 1 given below:
    Theorem 1.
    Let $\hat{\Phi}_{1}$ and $\hat{\Phi}_{2}$ be consistent for $\Phi$, so that
    $\hat{\Phi}_{k}=\Phi+o_{P}(N),\qquad k=1,2$ (9)
    Then $\hat{\tau}$ converges to a lower bound on $\tau^{\textup{basic}}$, so that
    $\hat{\tau}\leq\tau^{\textup{basic}}+o_{P}(N)$ (10)
  2. 2.
    Despite its conservative nature as a lower bound, $\hat{\tau}$ will take large values when the difference in average outcomes is large between treated and control. To illustrate, Figure 1 shows an experiment in which the control units have substantially higher outcomes than the treated. As can be seen, the large difference in outcomes forces the set $\{i:X_{i}=Y_{i}\}$ to be small in size, and equivalently |$|\{i:X_{i}=Y_{ki}\}|$ must be small for $k=1$. In turn, this forces $\hat{\Phi}_{1}(\phi)$ given by (7) to be small for all $\phi$, and hence forces $\hat{\tau}$ given by (8) to be large.
    In fact, under mild conditions $\hat{\tau}$ converges to the inverse propensity weighted (IPW) difference in average outcomes between treated and control, as stated by Theorem 2:
    Theorem 2.
    Let $X$ satisfy
    $\sum_{i=1}^{N}\frac{X_{i}}{P(X_{i}=1)}=N+o_{P}(N)\quad\textup{and}\quad\sum_{i=1}^{N}\frac{1-X_{i}}{P(X_{i}=0)}=N+o_{P}(N)$ (11)
    so that the total weights of the treated and control converge to their expectations. Then $\hat{\tau}$ is asymptotically the IPW difference in means
    $\hat{\tau}=\left|\sum_{i=1}^{N}\frac{X_{i}Y_{i}}{P(X_{i}=1)}-\sum_{i=1}^{N}\frac{(1-X_{i})Y_{i}}{P(X_{i}=0)}\right|+o_{P}(N)$ (12)
Figure 1: Experiment in which control units have substantially higher outcomes than treated, resulting in a small number of units for which $X_{i}=Y_{i}$ (black outlined rectangles)
Example 1.
Consider an experiment in which $100$ units are divided equally between treatment and control, where the average outcome for the treated and control groups are respectively $0.8$ (40/50) and $0.4$ (20/50). Units who are unaffected by treatment have fixed outcomes, and are equally likely to be assigned to treatment or control. It follows that up to sampling variation, such units should be evenly distributed between treatment and control even after stratifying by outcome. Therefore, out of the 40 treated units whose outcome equals 1, at most roughly 20 such units could be unaffected by treatment, if they are to be equally represented in the control group. Likewise out of the 30 control units whose outcome equals 0, at most roughly 10 could be unaffected by treatment, if they are to be equally represented in the treatment group. This yields an upper bound of $(20+10)\times 2=60$ units who are unaffected by treatment, or equivalently a lower bound of 40 units who are affected by treatment. The lower bound corresponds to the difference in means of 0.4 between treatment and control. The limiting quantities that were used to construct this bound, namely the 20 control units with outcome equal to 1, and the 10 treated units with outcome equal to 0, are equivalent to the set $\{i:Y_{i}\neq X_{i}\}$ utilized by the estimator $\hat{\Phi}_{2}$.
Continuing the example, we describe a generative model in which the contrast in outcomes is tight for the number of units affected by treatment. In this model, the effect on unit $i$ of $X_{i}$, the contrast variable appearing in (12), is monotone–that is, not positive for some units and negative for others. We then describe two models under which the lower bound is loose, and which suggest that non-vaccuous upper bounds are not identifiable.
  1. 1.
    (Nonnegative effects) Of the 100 units in the experiment, let 40 units have outcome 1 regardless of treatment, 20 units have outcome 0 regardless of treatment, and the remaining 40 units have outcome 1 if treated and 0 if control. Under this model, the expected outcomes for treated and control can be seen to be $0.8$ and $0.4$ as described above, and the empirical diff-in-means will equal (up to sampling variation) the fraction of units affected by treatment.
  2. 2.
    (Cancellation of effects) Of the 100 units in the experiment, let 20 units have outcome 1 regardless of treatment, 20 units have outcome 0 if treated and 1 if control, and the remaining 60 units have outcome 1 if treated and 0 if control. Under this model, direct treatment affects some units positively and others negatively. While the expected outcomes match the previous example, here the diff-in-means is a loose lower bound on the fraction of units affected by treatment, which is 0.8.
  3. 3.
    (Pathological interference) Let $Y_{i}$ equal
    $Y_{i}$ =$=\begin{cases}\alpha_{i1}X_{i}+\alpha_{i0}(1-X_{i})&\textup{ if $\sum_{i}X_{i}$ is even}\\ \beta_{i1}X_{i}+\beta_{i0}(1-X_{i})&\textup{ if $\sum_{i}X_{i}$ is odd}\end{cases}$ (13)
    where $(\alpha_{i1},\alpha_{i0})$ describe the potential outcomes of the first example, equaling $(1,1)$ for units 1-40, (0,0) for units 41-60, and $(1,0)$ for units 61-100, while $(\beta_{i1},\beta_{i0})$ takes values $(1,0)$ and $(0,1)$ in some arbitrary proportion. If half of the units are assigned to treatment under the experiment design, then the outcomes will be indistinguishable from those of the first example; however, unlike the first example, all 100 units are affected by treatment.
Using Network Information
To utilize $W$, the thresholded number of treated neighbors given by (2), we may let $\hat{\Phi}_{1}$ and $\hat{\Phi}_{2}$ equal
$\hat{\Phi}_{k}=\sum_{i:\phi_{i}^{*}=1}\frac{1\{X_{i}=W_{i}=Y_{ki}\}}{{P}(X_{i}=W_{i}=Y_{ki})},\qquad k=1,2.$ (14)
For this version of $\hat{\Phi}_{k}$, we may show that $\hat{\tau}$ remains a consistent lower bound on $\tau^{\textup{basic}}$ under the conditions of Theorem 2, but converges to a different contrast given by
$\hat{\tau}=\left|\sum_{i=1}^{N}\frac{X_{i}W_{i}}{P(X_{i}W_{i}=1)}Y_{i}-\sum_{i=1}^{N}\frac{(1-X_{i})(1-W_{i})}{P(X_{i}=0,W_{i}=0)}Y_{i}\right|+o_{P}(N)$ (15)
This is the IPW difference in outcomes between units whose exposure variable $(X_{i},W_{i})$ equals (1,1) versus (0,0). If the effects of direct and indirect treatment reinforce each other, this contrast may be larger than the comparison of treated and control given by (12), implying a larger and tighter lower bound on $\tau^{\textup{basic}}$.
To give a single formula that includes both (5) and (14) as special cases, let $\hat{\Phi}_{k}$ equal
$\hat{\Phi}_{k}=\sum_{i:\phi_{i}^{*}=1}\frac{1\{Z_{i}=Y_{ki}\}}{P(Z_{i}=Y_{ki})},$ (16)
where $Z_{i}\in\{1,0,-1\}$ divides the units into 3 groups, as a function of each unit’s direct and indirect treatment. Under (16), $\hat{\tau}$ will converge to an IPW contrast between the units for which $Z_{i}$ equals 1 and 0 (“effective treatment” versus “effective control”), while units for which $Z_{i}=-1$ are not used. Letting $Z_{i}=X_{i}$ causes (16) to be equivalent to (5), while (14) is recovered by letting $Z_{i}$ equal
$Z_{i}=\begin{cases}1&\textup{ if $X_{i}=W_{i}=1$}\\ 0&\textup{ if $X_{i}=W_{i}=0$}\\ -1&\textup{ if $X_{i}\neq W_{i}$}\end{cases}$
3.4 Hájek Normalization
To reduce the variance of our IPW-based estimators, we may use Hájek normalization (Aronow and Samii, , 2017; Gao and Ding, , 2023). To construct a Hájek-normalized version of $\hat{\Phi}_{k}$ given by (16), we first observe that it equals
$\hat{\Phi}_{k}=\sum_{i:\phi_{i}^{*}=1}\left(\frac{Z_{i}Y_{ki}}{P(Z_{i}=1)}+\frac{(1-Z_{i})(1-Y_{ki})}{P(Z_{i}=0)}\right)$ (17)
where we have divided the event $\{Z_{i}=Y_{ki}\}$ into sub-events $\{Z_{i}=Y_{ki}=1\}$ and $\{Z_{i}=Y_{ki}=0\}$, and replaced the propensity score $P(Z_{i}=Y_{ki})$ by $Y_{ki}P(Z_{i}=1)+(1-Y_{ki})P(Z_{i}=0)$, which can be done because $Y_{ki}$ is deterministic when $\phi_{i}^{*}=1$. Eq. (17) suggests a Hájek normalized version given by
$\hat{\Phi}_{k}^{{\textup{Haj}}}=\sum_{i:\phi_{i}^{*}=1}\left[\frac{N}{\hat{N}_{1}}\cdot\frac{Z_{i}}{P(Z_{i}=1)}\cdot Y_{ki}+\frac{N}{\hat{N}_{0}}\cdot\frac{1-Z_{i}}{P(Z_{i}=0)}\cdot(1-Y_{ki})\right],$ (18)
where $\hat{N}_{0}$ and $\hat{N}_{1}$ are estimates of $N$ given by
$\hat{N}_{0}$ $=\sum_{j=1}^{N}\frac{1\{Z_{j}=0\}}{P(Z_{j}=0)},$ $\hat{N}_{1}$ $=\sum_{j=1}^{N}\frac{1\{Z_{j}=1\}}{P(Z_{j}=1)}.$
To construct a Hájek-weighted point estimate for $\tau^{\textup{basic}}$, we can substitute $\hat{\Phi}_{k}^{{\textup{Haj}}}(\phi)$ for $\hat{\Phi}_{k}(\phi)$ in the expression for $\hat{\tau}$ given by (8). In Theorem 4, we show that the resulting $\hat{\tau}$ exactly equals a Hájek-weighted comparison of units, given by
$\hat{\tau}=\left|\sum_{i=1}^{N}\frac{N}{\hat{N}_{1}}\cdot\frac{Z_{i}}{P(Z_{i}=1)}Y_{i}-\sum_{i=1}^{N}\frac{N}{\hat{N}_{0}}\cdot\frac{(1-Z_{i})}{P(Z_{i}=0)}Y_{i}\right|.$
3.5 Inference and Variance Estimation
If $\hat{\Phi}_{1}^{\textup{Haj}}$ and $\hat{\Phi}_{2}^{\textup{Haj}}$ are asymptotically normal, with consistent variance estimators denoted by $V_{1}$ and $V_{2}$, then by combining 1-sided confidence intervals it holds with probability converging to at least $1-\alpha$ that
$\Phi$ $\leq\min\left\{\hat{\Phi}_{1}^{\textup{Haj}}+z_{1-\frac{\alpha}{2}}\sqrt{V_{1}},\,\hat{\Phi}_{2}^{\textup{Haj}}+z_{1-\frac{\alpha}{2}}\sqrt{V_{2}}\right\}.$ (19)
To find a one-sided confidence interval for $\tau^{\textup{basic}}=N-\Phi$, we will upper bound $\Phi$ by (19) maximized over hypotheses for the unknown $\phi^{*}$. This yields the confidence statement that with probability converging to at least $(1-\alpha)$, it holds that
$\tau^{\textup{basic}}\geq N-\min\Big{(}\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{1}^{\textup{Haj}}(\phi)+z_{1-\frac{\alpha}{2}}\sqrt{V_{1}(\phi)},\,\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{2}^{\textup{Haj}}(\phi)+z_{1-\frac{\alpha}{2}}\sqrt{V_{2}(\phi)}\,\Big{)}$ (20)
where $V_{k}(\phi)$ denotes $V_{k}$, evaluated under the hypothesis that $\phi^{*}=\phi$. We give an expression for $V_{k}$ next.
Variance Estimation
The expression for $\hat{\Phi}_{k}^{\textup{Haj}}$ given by (18) is equivalent to
$\hat{\Phi}_{k}^{{\textup{Haj}}}=\sum_{i:\phi_{i}^{*}=1}\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{Z_{i}=Y_{ki}\}}{P(Z_{i}=Y_{ki})}$
It follows that the variance of $\hat{\Phi}_{k}$ can be written as a sum of covariance terms,
${\operatorname{Var}\,}\hat{\Phi}_{k}^{\textup{Haj}}=\sum_{i:\phi_{i}^{*}=1}\sum_{j:\phi_{j}^{*}=1}{\operatorname{Cov}}\left(\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{Z_{i}=Y_{ki}\}}{P(Z_{i}=Y_{ki})}\ ,\ \frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{Z_{i}=Y_{ki}\}}{P(Z_{i}=Y_{ki})}\right)$ (21)
To construct the one-sided interval (20), we propose $V_{k}$ given by
$V_{k}=\sum_{i:\phi_{i}^{*}=1}\sum_{j:\phi_{j}^{*}=1}{\operatorname{Cov}}\left(\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{Z_{i}=Y_{ki}\}}{P(Z_{i}=Y_{ki})}\ ,\ \frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{Z_{i}=Y_{ki}\}}{P(Z_{i}=Y_{ki})}\right)\\ \cdot\frac{1\{Z_{i}=Y_{ki},Z_{j}=Y_{kj}\}}{P(Z_{i}=Y_{ki},Z_{j}=Y_{kj})}$ (22)
Comparing (21) and (22), it can be seen that $V_{k}$ is a Horvitz-Thompson estimator which samples the covariance terms appearing in $\operatorname{Var}\hat{\Phi}_{k}$. Hence it is unbiased provided that the sampling probabilities $P(Z_{i}=Y_{ki},Z_{j}=Y_{kj})$ are non-zero for all $i,j\in[N]^{2}$. Should this condition not hold, our variance estimators will be conservatively biased, as will be established in Theorem 5.
To evaluate (22), we may use the following expressions for the covariance and propensity values, which hold whenever $Y_{ki}$ and $Y_{kj}$ are deterministic, and hence for all $(i,j)$ such that $\phi_{i}^{*}\phi_{j}^{*}=1$,
${\operatorname{Cov}}\left(\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{Z_{i}=Y_{ki}\}}{P(Z_{i}=Y_{ki})}\ ,\ \frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{Z_{i}=Y_{ki}\}}{P(Z_{i}=Y_{ki})}\right)=\\ \sum_{a=0}^{1}\sum_{b=0}^{1}1\{Y_{ik}=a,Y_{jk}=b\}{\operatorname{Cov}}\left(\frac{N}{\hat{N}_{a}}\cdot\frac{1\{Z_{i}=a\}}{P(Z_{i}=a)}\ ,\ \frac{N}{\hat{N}_{b}}\cdot\frac{1\{Z_{i}=b\}}{P(Z_{i}=b)}\right)$ (23)
and
$P(Z_{i}=Y_{ki},Z_{j}=Y_{kj})=\sum_{a=0}^{1}\sum_{b=0}^{1}1\{Y_{ik}=a,Y_{jk}=b\}P(Z_{i}=a,Z_{j}=b).$ (24)
Using these expressions, we can evaluate $V_{1}(\phi)$, which equals $V_{1}$ under the hypothesis that $\phi^{*}=\phi$, for any $\phi\in\{0,1\}^{N}$.
The motivation for our sampling-based variance estimator is that it takes small value when the treated and control units have large differences in average outcome. Analogous to the intuition portrayed in Figure 1, when the diff-in-means is large, the number of pairs $(i,j)$ such that the event $\{Z_{i}=Y_{ki},Z_{j}=Y_{kj}\}$ occurs must be small for either $k=1$ or $k=2$, and hence $V_{1}(\phi)$ or $V_{2}(\phi)$ will be small, even if maximized over $\phi\in\{0,1\}^{N}$.
We discuss some conceptual differences between the variance estimator of our setting and that of treatment effect estimation. Generally, the variance of a treatment effect estimator involves products between pairs of potential outcomes that cannot be simultaneously observed, such as the outcome of the same unit under both treatment and control. Such product terms cannot be estimated in an unbiased manner, but rather must be conservatively bounded (Harshaw et al., , 2021). These difficulties do not arise in the variance expression (21), as it involves only the outcomes of units who are unaffected by treatment. As a result, $V_{k}$ will be unbiased under conditions to be described in Theorem 5, but must itself be bounded over the unknown set of units who are unaffected by treatment, as is done in (20).
3.6 Computation
To construct the one-sided interval (20), we must solve the optimization problem
$\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{k}(\phi)+z_{1-\frac{\alpha}{2}}\sqrt{V_{k}(\phi)},$ (25)
for $k=1,2$. We can write this problem as
$\max_{\phi\in\{0,1\}^{N}}v^{T}\phi+z_{1-\frac{\alpha}{2}}\sqrt{\phi^{T}Q\phi}$ (26)
where $\hat{\Phi}_{k}(\phi)=v^{T}\phi$ and $V_{k}(\phi)=\phi^{T}Q\phi$ for $v\in\mathbb{R}^{N}$ and $Q\in\mathbb{R}^{N\times N}$ given by
$v_{i}$ $=\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{Z_{i}=Y_{ki}\}}{Y_{ki}\cdot P(Z_{i}=1)+(1-Y_{ki})\cdot P(Z_{i}=0)}$ $i\in[N]$ (27)
$Q_{ij}$ $=C_{ij}\cdot\frac{1\{Z_{i}=Y_{ki},Z_{j}=Y_{kj}\}}{P_{ij}}$ $i,j\in[N]^{2}.$ (28)
in which $C_{ij}$ and $P_{ij}$ are given by the right hand sides of (23) and (24) respectively
To solve (26), we let $q_{i}$ denote the $i$th row of $Q$, let $B=\sum_{ij}|Q_{ij}|$, and reformulate the problem as
$\max_{\phi,u}$ $\ v^{T}\phi+z_{1-\frac{\alpha}{2}}\sqrt{\sum_{i=1}^{N}u_{i}}$
subject to $\ u_{i}\leq q_{i}^{T}\phi+B(1-\phi_{i})$ $\forall\ i\in[N]$
$\ u_{i}\geq q_{i}^{T}\phi-B(1-\phi_{i})$ $\forall\ i\in[N]$
$\ u_{i}\leq B\phi_{i}$ $\forall\ i\in[N]$
$\ u_{i}\geq B\phi_{i}$ $\forall\ i\in[N]$
$\ \phi\in\{0,1\}^{N}$
$\ u\in\mathbb{R}^{N}$
It can be seen that $u_{i}=0$ if $\phi_{i}=0$, and $u_{i}=q_{i}^{T}\phi$ if $\phi_{i}=1$, enforcing that $\sum_{i}u_{i}=\phi^{T}Q\phi$. We solve this binary program using Gurobi through YALMIP (Lofberg, , 2004). For the instances that we considered, involving 1000 or more units, Gurobi was generally able to solve the problem to acceptable tolerances. If stopped early, Gurobi returns an upper bound, retaining the coverage properties of the actual solution.
4 Generalization to other estimands
We may generalize the approach of Section 3 to other estimands such as
  • The number of units affected by the treatment of others
  • The number of units affected by the treatment of others, excluding close neighbors
  • The number of units affected by the treatment of others when their direct treatment satisfies some condition
These estimands and their estimators will follow a general recipe.
  1. 1.
    First we introduce a vector $\phi^{*}\in\{0,1\}^{N}$ that identifies the units who do not meet the criteria of the estimand, and let $\Phi=\sum_{i=1}^{N}\phi_{i}^{*}$, so that the estimand is given by
    $\tau=N-\Phi$
  2. 2.
    We then introduce a pair of estimators for $\Phi$, which we denote by $\hat{\Phi}_{k}$ for $k=1,2$, or by $\hat{\Phi}_{k}^{{\textup{Haj}}}$ for their Hájek-weighted versions. A consistent lower bound for the estimand $\tau$ is given by
    $\hat{\tau}=N-\min\left\{\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{1}^{\textup{Haj}}(\phi)\,,\,\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{2}^{\textup{Haj}}(\phi)\right\}$ (29)
    where $\hat{\Phi}_{k}^{\textup{Haj}}(\phi)$ denotes $\hat{\Phi}_{k}^{\textup{Haj}}$, evaluated under the hypothesis that $\phi^{*}=\phi$.
  3. 3.
    If $\hat{\Phi}_{k}^{\textup{Haj}}$ for $k=1,2$ is consistent and asymptotically normal, with variance consistently estimated by $V_{k}$, then we can construct an asymptotic $(1-\alpha)$ level one-sided confidence interval for $\tau$, with lower boundary given by
    $N-\min\left\{\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{1}^{\textup{Haj}}(\phi)+z_{1-\frac{\alpha}{2}}\sqrt{V_{1}(\phi)}\,,\,\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{2}^{\textup{Haj}}(\phi)+z_{1-\frac{\alpha}{2}}\sqrt{V_{2}(\phi)}\,\right\},$ (30)
    where $V_{k}(\phi)$ denotes $V_{k}$, evaluated under the hypothesis that $\phi^{*}=\phi$.
Following this recipe, we introduce the estimands $\tau^{\textup{indirect}}$, $\tau^{\textup{nonneighbors}}$, $\tau^{\textup{control}}$, and $\tau^{\textup{tr}}$, defining $(\phi^{*},\hat{\Phi}_{1},\hat{\Phi}_{2})$ for each one. After doing so, we present Hájek-normalized estimators and variance estimates in Sections 4.1 and 4.2.
The number of units affected by the treatment of others
Let $\phi_{i}^{*}\in\{0,1\}^{N}$ identify the units who are not affected by the treatment of others
$\phi_{i}^{*}=1\{\textup{if $Y_{i}$ is affected only by $X_{i}$, so that $Y_{i}=f_{i}(X_{i})$}\}$ (31)
and let $\tau^{\textup{indirect}}=N-\Phi$, the number of units affected by the treatment of others. To estimate $\tau^{\textup{indirect}}$, we use $\hat{\Phi}_{1}$ and $\hat{\Phi}_{2}$ given by
$\hat{\Phi}_{k}$ $=\sum_{i:\phi_{i}^{*}=1}\frac{1\{W_{i}=Y_{ki}\}}{{P}(W_{i}=Y_{ki}|X_{i})}$
Long-range interference
Let $\phi^{*}\in\{0,1\}^{N}$ identify the units who are affected only by their own treatment and that of their close neighbors,
$\phi_{i}^{*}=1\{\textup{if $Y_{i}$ is affected only by $X_{\eta_{i}}$, so that $Y_{i}=f_{i}(X_{\eta_{i}})$}\}$
and let $\tau^{\textup{nonneighbors}}$ equal $N-\Phi$, the number of units affected by the treatment of others excluding close neighbors. To estimate $\tau^{\textup{nonneighbors}}$, we use $\hat{\Phi}_{1}$ and $\hat{\Phi}_{2}$ given by
$\hat{\Phi}_{k}$ $=\sum_{i:\phi_{i}^{*}=1}\frac{1\{W_{i}^{(2)}=Y_{ki}\}}{{P}(W_{i}^{(2)}=Y_{ki}|X_{\eta_{i}})}$
where $W_{i}^{(2)}$ is given by (3), and measures treatment of non-close neighbors.
The number of units whose outcome under direct control (treatment) is affected by the treatment of others
Let $\phi^{*}\in\{0,1\}^{N}$ identify the units whose outcome, when receiving control themselves, does not depend on the treatment of others
$\phi_{i}^{*}=1\{\textup{ if $Y_{i}$ is constant for all $X$ such that $X_{i}=0$}\}$
in which case $\tau^{\textup{control}}=N-\Phi$ equals the number of units whose outcome under direct control is affected by the treatment of others. For example, if a vaccine grants immunity, then the disease status of a unit when receiving the placebo (but not necessarily when receiving the vaccine) might depend on the vaccination status of others.
To estimate $\tau^{\textup{control}}$, we consider $\hat{\Phi}_{1}$ and $\hat{\Phi}_{2}$ given by
$\hat{\Phi}_{k}$ $=\sum_{i:\phi_{i}^{*}=1}\frac{1\{X_{i}=0,W_{i}=Y_{ki}\}}{{P}(X_{i}=0,W_{i}=Y_{ki})}$
We may similarly define $\tau^{\textup{tr}}$, the number of units whose outcome under direct treatment is affected by the treatment of others, by defining $\phi^{*}$ to be
$\phi_{i}^{*}$ $=1\{\textup{ if $Y_{i}$ is constant for all $X$ such that $X_{i}=1$}\}$
and $\hat{\Phi}_{1}$ and $\hat{\Phi}_{2}$ to be
$\hat{\Phi}_{k}$ $=\sum_{i:\phi_{i}^{*}=1}\frac{1\{X_{i}=1,W_{i}=Y_{ki}\}}{{P}(X_{i}=1,W_{i}=Y_{ki})}$
4.1 General Functional Form
We give a general functional form for our estimands and estimators. In our estimands, $\phi^{*}\in\{0,1\}^{N}$ identifies the units whose outcome satisfies an exposure mapping condition for some subset of possible treatment assignments. We may write this as
$\phi_{i}^{*}=1\{\textup{ if $\exists\ \tilde{f}_{i}$ such that $f_{i}(X)=\tilde{f}_{i}(T_{i}(X))$ for all $X\in R_{i}$}\}$ (32)
where $T_{i}\equiv T_{i}(X)$ equals a vector of summary statistics describing unit $i$’s exposure to treatment. The set $R_{i}\subseteq\{0,1\}^{N}$ dictates when $T_{i}$ must be predictive of unit $i$’s outcome. For example, letting $T_{i}=X_{i}$ and $R_{i}=\{0,1\}^{N}$ induces $\phi^{*}$ given by (31).
The IPW estimators $\hat{\Phi}_{k}$ for $k=1,2$ are given by
$\hat{\Phi}_{k}=\sum_{i:\phi_{i}^{*}=1}\frac{1\{X_{i}\in R_{i},Z_{i}=Y_{ki}\}}{P(X_{i}\in R_{i},Z_{i}=Y_{ki}|T_{i})}$ (33)
for some choice of $Z_{i}\equiv Z_{i}(X)\in\{1,0,-1\}$ that maps unit $i$’s direct and indirect treatment to one of 3 groups (i.e., effective treatment, effective control, or not used).
Hájek-normalized versions of $\hat{\Phi}_{k}$ are given by
$\hat{\Phi}_{k}^{\textup{Haj}}=\sum_{i:\phi_{i}^{*}=1}\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{X\in R_{i},Z_{i}=Y_{ki}\}}{P(X\in R_{i},Z_{i}=Y_{ki}|T_{i})},$ (34)
where $\hat{N}_{1}$ and $\hat{N}_{0}$ denote Hájek normalization factors
$\hat{N}_{a}$ $=\sum_{j=1}^{N}\frac{1\{X\in R_{i},Z_{i}=a\}}{P(X\in R_{i},Z_{i}=a|T_{i})},\qquad a\in\{0,1\}.$ (35)
As will be shown in Theorem 4, using the Hájek-weighted estimates to construct $\hat{\tau}$ by (29) will result in a Hájek-weighted comparison of outcomes, given by
$\hat{\tau}=\left|\frac{N}{\hat{N}_{1}}\cdot\frac{1\{X\in R_{i},Z_{i}=1\}}{P(X\in R_{i},Z_{i}=1|T_{i})}Y_{i}-\sum_{i=1}^{N}\frac{N}{\hat{N}_{0}}\cdot\frac{1\{X\in R_{i},Z_{i}=0\}}{P(X\in R_{i},Z_{i}=0|T_{i})}Y_{i}\right|,$ (36)
Table 1 gives values for $\{T_{i},R_{i},Z_{i}\}_{i=1}^{N}$ corresponding to the estimands considered in this paper, while Table 2 gives expressions for the contrasts given by (36).
Estimand $T_{i}$ $R_{i}$ $Z_{i}$
$\tau^{\textup{basic}}$ 1 $\{0,1\}^{N}$ $X_{i}$
$\tau^{\textup{basic}}$ (w/network) 1 $\{0,1\}^{N}$ $1\{X_{i}=W_{i}=1\}-1\{X_{i}\neq W_{i}\}$
$\tau^{\textup{indirect}}$ $X_{i}$$\{0,1\}^{N}$ $W_{i}$
$\tau^{\textup{nonneighbors}}$ $X_{\eta_{i}}$$\{0,1\}^{N}$ $W_{i}^{(2)}$
$\tau^{\textup{control}}$ 1 $\{X\in\{0,1\}^{N}:X_{i}=0\}$ $W_{i}$
$\tau^{\textup{tr}}$ 1 $\{X\in\{0,1\}^{N}:X_{i}=1\}$ $W_{i}$
Table 1: Exposure mappings, subsets, and contrast mappings for various estimands. Note that letting $T_{i}=1$ for all $X$ enforces $Y_{i}$ to equal a constant for all $X\in R_{i}$, should $\phi_{i}^{*}=1$.
Estimand Hájek-normalized $\hat{\tau}$, expressed as contrast given by (36)
$\tau^{\textup{basic}}$ ${\left|\sum_{i=1}^{N}\frac{N}{\hat{N}_{1}}\frac{X_{i}}{P(X_{i}=1)}Y_{i}-\sum_{i=1}^{N}\frac{N}{\hat{N}_{0}}\frac{1-X_{i}}{P(X_{i}=0)}Y_{i}\right|}$
$\tau^{\textup{basic}}$ (w/network) ${\left|\sum_{i=1}^{N}\frac{N}{\hat{N}_{1}}\frac{1\{X_{i}=W_{i}=1\}}{P(X_{i}=W_{i}=1)}Y_{i}-\sum_{i=1}^{N}\frac{N}{\hat{N}_{0}}\frac{1\{X_{i}=W_{i}=0\}}{P(X_{i}=W_{i}=0)}Y_{i}\right|}$
$\tau^{\textup{indirect}}$ ${\left|\sum_{i=1}^{N}\frac{N}{\hat{N}_{1}}\frac{1\{W_{i}=1\}}{P(W_{i}=1|X_{i})}Y_{i}-\sum_{i=1}^{N}\frac{N}{\hat{N}_{0}}\frac{1\{W_{i}=0\}}{P(W_{i}=0|X_{i})}Y_{i}\right|}$
$\tau^{\textup{nonneighbors}}$ ${\left|\sum_{i=1}^{N}\frac{N}{\hat{N}_{1}}\frac{1\{W_{i}^{(2)}=1\}}{P(W_{i}^{(2)}=1|X_{\eta_{i}})}Y_{i}-\sum_{i=1}^{N}\frac{N}{\hat{N}_{0}}\frac{1\{W_{i}^{(2)}=0\}}{P(W_{i}^{(2)}=0|X_{\eta_{i}})}Y_{i}\right|}$
$\tau^{\textup{control}}$ ${\left|\sum_{i=1}^{N}\frac{N}{\hat{N}_{1}}\frac{1\{X_{i}=0,W_{i}=1\}}{P(X_{i}=0,W_{i}=1)}Y_{i}-\sum_{i=1}^{N}\frac{N}{\hat{N}_{0}}\frac{1\{X_{i}=0,W_{i}=0\}}{P(X_{i}=0,W_{i}=0)}Y_{i}\right|}$
$\tau^{\textup{tr}}$ ${\left|\sum_{i=1}^{N}\frac{N}{\hat{N}_{1}}\frac{1\{X_{i}=1,W_{i}=1\}}{P(X_{i}=1,W_{i}=1)}Y_{i}-\sum_{i=1}^{N}\frac{N}{\hat{N}_{0}}\frac{1\{X_{i}=1,W_{i}=0\}}{P(X_{i}=1,W_{i}=0)}Y_{i}\right|}$
Table 2: Hájek weighted point estimate $\hat{\tau}$, expressed as contrasts given by (36). The terms $\hat{N}_{1}$ and $\hat{N}_{0}$ are defined by (35).
4.2 Variance Estimation for General Case
Let $v_{i}$ denote
$v_{i}=\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{X\in R_{i},Z_{i}=Y_{ki}\}}{P(X\in R_{i},Z_{i}=Y_{ki}|T_{i})}$ (37)
so that $\hat{\Phi}_{k}^{\textup{Haj}}=\sum_{i:\phi_{i}^{*}=1}v_{i}$. It follows that the variance of $\hat{\Phi}_{k}^{\textup{Haj}}$ can be written as a sum of covariances
${\operatorname{Var}\,}\hat{\Phi}_{k}^{\textup{Haj}}=\sum_{i:\phi_{i}^{*}=1}\sum_{j:\phi_{j}^{*}=1}{\operatorname{Cov}}(v_{i},v_{j})$
Unlike the special case where $T_{i}$ is constant for all units, generally we cannot know or compute ${\operatorname{Cov}}(v_{i},v_{j})$ without knowledge of the outcome models $f_{i}$ and $f_{j}$, even if $\phi_{i}^{*}\phi_{j}^{*}=1$. As a result, the covariance terms must be estimated. We thus propose variance estimator $V_{k}$ given by
$V_{k}=\sum_{i:\phi_{i}^{*}=1}\sum_{j:\phi_{j}^{*}=1}\Bigg{\{}\mathbb{E}\left[v_{i}v_{j}|T_{i},T_{j}\right]-\mathbb{E}\left[v_{i}|T_{i}\right]\cdot\mathbb{E}\left[v_{j}|T_{j}\right]\cdot\\ \frac{P(T_{i})\cdot P(T_{j})}{P(T_{i},T_{j})}\Bigg{\}}\cdot\frac{1\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}}{P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})}$ (38)
where $\bar{Z}_{i}$ equals $Z_{i}$ if $X\in R_{i}$ and equals -1 otherwise,
$\bar{Z}_{i}=\begin{cases}Z_{i}&\textup{ if $X\in R_{i}$}\\ -1&\textup{ if $X\notin R_{i}$,}\end{cases}$ (39)
allowing the event $\{X\in R_{i},Z_{i}=Y_{ki}\}$ to be written more concisely as $\{\bar{Z}_{i}=Y_{ki}\}$ under binary outcomes.
We will show in Theorem 5 that (38) is conservatively biased, with the exception that if $T_{i}$ is a constant for each unit (as is the case for $\tau^{\textup{basic}}$, $\tau^{\textup{control}}$, and $\tau^{\textup{tr}}$), then the estimator simplifies to a weighted sample of covariance terms
$V_{k}=\sum_{i:\phi^{*}=1}\sum_{j:\phi^{*}=1}{\operatorname{Cov}}(v_{i},v_{j})\cdot\frac{1\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}}{P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj})},$
and is unbiased if the sampling probabilities $P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj})$ are non-zero, and conservative otherwise.
Theorem 3 gives computable expressions for the terms appearing in (38). Using these expressions, we may evaluate $V_{k}(\phi)$, the value of $V_{k}$ under the hypothesis that $\phi^{*}=\phi$, without knowledge of the true outcome models $\{f_{i}\}$.
Theorem 3.
If $\phi_{i}^{*}=1$ and $\bar{Z}_{i}=Y_{ki}$, then it holds that
$\mathbb{E}[v_{i}|T_{i}]$ $=\sum_{a=0}^{1}1\{Y_{ki}=a\}\mathbb{E}\left[\cdot\frac{1\{\bar{Z}_{i}=a\}N}{P(\bar{Z}_{i}=a|T_{i})\hat{N}_{a}}|T_{i}\right],$
If $\phi_{i}^{*}\phi_{j}^{*}=1$ and $11\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}=$ 1, then it holds that
$\mathbb{E}[v_{i}v_{j}|T_{i},T_{j}]=$ $\sum_{a,b=0}^{1}1\{Y_{ki}=a,Y_{kj}=b\}\mathbb{E}\left[\frac{1\{\bar{Z}_{i}=a\}N}{P(\bar{Z}_{i}=a|T_{i})\hat{N}_{a}}\cdot\frac{1\{\bar{Z}_{j}=b\}N}{P(\bar{Z}_{j}=b|T_{j})\hat{N}_{b}}|T_{i},T_{j}\right]$
and also that
$P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})=$ $\sum_{a,b=0}^{1}1\{Y_{ki}=a,Y_{kj}=b\}P(\bar{Z}_{i}=a,\bar{Z}_{j}=b|T_{i},T_{j})$
5 Theory
In this section, we present consistency and coverage results for the general class of Hájek-normalized estimators described in Section 4. Specifically, we show that our point estimate $\hat{\tau}$ is consistent for a lower bound on the estimand while equaling a Hájek-weighted contrast; we show that the variance estimators $V_{k}$ concentrates and is either unbiased or conservative for the variance that is required for our interval estimates to cover; and we show that our intervals have the desired asymptotic coverage, albeit with a technical modification to cover degenerate settings in which normality does not hold.
To describe our requirements on the experiment design, for $i\in[N]$ let $\Gamma_{i}\subseteq[N]$ denote the smallest subset of units whose treatment assignment is sufficient to determine the values of $T_{i},Z_{i}$, and $1\{X\in R_{i}\}$ for all $X\in\{0,1\}^{N}$. We will call $\Gamma_{i}$ the design neighborhood of unit $i$. We stress that it places no restrictions on the outcome model or interference.
Assumption 1 describes the experiment design that we study. It assumes a sequence of experiments in which the design neighborhoods are bounded in size and overlap, Bernoulli randomization with treatment probabilities bounded away from 0 and 1, and places a positivity condition on $P(\bar{Z}_{i}|T_{i})$ under the hypothesis that $\phi_{i}^{*}=1$. While many experiment designs are possible, we study independent Bernoulli randomization as a starting point, leading to the simplest proofs of consistency and asymptotic normality.
Assumption 1.
The treatment assignments $\{X_{i}\}_{i=1}^{N}$ are indpendent Bernoulli random variables, which satisfy for fixed constants $d_{\max}$ and $p_{\min}>0$ as $N\rightarrow\infty$ that
$|\Gamma_{i}|$ $\leq d_{\max}$ $\forall\ i\in[N]$ (40)
$\sum_{j=1}^{N}1\{x_{i}\in\Gamma_{j}\}$ $\leq d_{\max}$ $\forall\ i\in[N]$ (41)
$p_{\min}\leq P(X_{i}=1)$ $\leq 1-p_{\min}$ $\forall\ i\in[N]$ (42)
and for all $i:\phi_{i}^{*}=1$, it holds for $\bar{Z}$ given by (39) that
[…]
The support condition (43) requires $T_{i}$, $Z_{i}$, and $R_{i}$ to be chosen so that the entries of $X$ needed to determine $Z_{i}$ and membership in $R_{i}$ do not fully overlap with those needed to determine $T_{i}$. For example, if $T_{i}=X_{\eta_{i}}$, then $Z_{i}$ and $R_{i}$ should depend at least partially on units outside of $\eta_{i}$.
Theorem 4 is a consistency result. It states that $\hat{\Phi}^{\textup{Haj}}_{k}$ converges to $\Phi$, and that $\hat{\tau}$ is asymptotically a lower bound for $\tau$, and equals a Hájek weighted contrast between units, given by (46). Table 2 gives examples of (46) for the estimands considered in the paper.
Theorem 4.
Let Assumption 1 hold. Then
$\hat{\Phi}_{k}^{\textup{Haj}}=\Phi+O_{P}(N^{1/2}),\qquad k=1,2$ (44)
and for $\hat{\tau}=N-\min\left\{\max\limits_{\phi\in\{0,1\}}\hat{\Phi}_{1}^{\textup{Haj}}(\phi)\ ,\ \max\limits_{\phi\in\{0,1\}}\hat{\Phi}_{2}^{\textup{Haj}}(\phi)\right\}$ it holds that
$\hat{\tau}\leq\tau+O_{P}(N^{1/2}).$ (45)
For arbitrary experiment designs (i.e., no longer requiring Assumption 1), it holds that
$\hat{\tau}=\left|\frac{1\{X\in R_{i},Z_{i}=1\}N}{P(X\in R_{i},Z_{i}=1|T_{i})\hat{N}_{1}}Y_{i}-\sum_{i=1}^{N}\frac{1\{X\in R_{i},Z_{i}=0\}N}{P(X\in R_{i},Z_{i}=0|T_{i})\hat{N}_{0}}Y_{i}\right|,$ (46)
where $\hat{N}_{a}=\displaystyle{\sum_{i=1}^{N}}\dfrac{1\{X\in R_{i},Z_{i}=a\}}{P(X\in R_{i},Z_{i}=a|T_{i})}$.
Theorem 5 pertains to variance estimation. It states that $V_{k}$ converges to an upper bound on the variance of $\hat{\Phi}_{k}^{\textup{Haj}}$, and gives conditions under which the bound will be tight.
Theorem 5.
Let Assumption 1 hold. Then for any $\delta>0$ it holds that
$V_{k}=\mathbb{E}[V_{k}]+O_{P}(N^{1/2+\delta}),\qquad k=1,2.$ (47)
Additionally, it holds that
$\mathbb{E}V_{k}\geq{\operatorname{Var}\,}\hat{\Phi}_{k}^{\textup{Haj}},$ (48)
with equality if the exposure mappings $\{T_{i}\}_{i=1}^{N}$ are constant and
$P(\bar{Z}_{i}=a,\bar{Z}_{j}=b)>0$ for all $a,b\in\{0,1\}$ and $i,j\in[N]^{2}$.
Theorem 6 gives coverage properties for the proposed confidence interval given by (30). It states that if the variance estimators $V_{k}$ are lower bounded by a vanishing fraction of $N$, then the intervals will have at least the desired asymptotic coverage level.
Theorem 6.
Let Assumption 1 hold. Let $\widetilde{V}_{k}(\phi)$ denote the thresholded variance estimate given by
$\widetilde{V}_{k}(\phi)=\max\left\{V_{k}(\phi)\ ,\ \frac{1.01\ N^{2/3+0.01}}{z_{1-\frac{\alpha}{2}}^{2}\cdot\alpha/2}\right\}$
and let $CI$ denote
$CI=N-\min\left\{\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{1}^{\textup{Haj}}(\phi)+z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{1}(\phi)}\,,\,\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{2}^{\textup{Haj}}(\phi)+z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{2}(\phi)}\right\}$ (49)
Then $\tau\geq CI$ with probability converging to at least $1-\alpha$.
All theorems are proven in the appendix.
6 Examples
6.1 Simulation Study
Table 3 shows the results of a simulation study in which the outcomes depended on both direct and indirect treatment. For all estimands except $\tau^{\textup{nonneighbors}}$, outcomes were generated according to
$Y_{i}=1\{\textup{if }u_{i}\leq\alpha_{0i}+\alpha_{1i}\cdot X_{i}+\alpha_{2i}\cdot W_{i}\}$ (50)
where the parameters $(u_{i},\alpha_{0i},\alpha_{1i},\alpha_{2i})$ were randomly generated, with $u_{i}\sim\textup{Unif }[0,1]$, and $\alpha_{0i},\alpha_{1i},\alpha_{2i}$ nonnegative. For simulations involving $\tau^{\textup{nonneighbors}}$, the outcome model was changed to
$Y_{i}=1\{\textup{if }u_{i}\leq\alpha_{0i}+\alpha_{1i}\cdot X_{i}+\alpha_{2i}\cdot\bar{W}_{i}\}$ (51)
where $\bar{W}_{i}$ denotes the thresholded number of treated units in $\eta_{i}\cup\eta_{i}^{(2)}$ excluding unit $i$, so that neighbors in $\eta_{i}$ and $\eta_{i}^{(2)}$ were equally important in determining the outcome of unit $i$. Unit $i$ was considered to be affected by treatment if $Y_{i}$ varied as a function of $X$ under the generated parameter values, and was considered to be affected by treatment of others if $Y_{i}$ varied as a function of $X_{-i}$ for any value of $X_{i}$, and so forth.
Hájek normalization was used on all estimators. Estimation of $\tau^{\textup{basic}}$ involved both $W$ and $X$, following (14) with Hájek normalization. For comparison of coverage rates and interval widths, the treatment effect
$EATE=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}[Y_{i}|X\in R_{i},Z_{i}=1]-\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}[Y_{i}|X\in R_{i},Z_{i}=0],$
was also estimated, for $\{R_{i}\}$ and $\{Z_{i}\}$ given by Table 1 for each estimand. For example, when estimating $\tau^{\textup{indirect}}$, we compared to the indirect effect given by
$\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}[Y_{i}|W_{i}=1]-\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}[Y_{i}|W_{i}=0].$
For estimation and inference of these treatment effects, Hájek normalization and conservative variance estimates were used following Gao and Ding, (2023).
The results of the simulations are the following. For the estimands $\tau^{\textup{basic}}$, $\tau^{\textup{tr}}$, and $\tau^{\textup{control}}$, point estimates were unbiased (as might be expected given that the unit effects were monotone functions of the contrast variables), and the confidence intervals exhibited approximate calibration of coverage. For these estimands, the exposure mapping $T_{i}$ is constant, allowing for unbiased variance estimation as discussed in Theorem 5. For the remaining estimands, estimation was conservatively biased, and interval coverage rates exceeded the desired level. Table 3 also shows that the interval widths for our estimands were competitive, and often smaller than, those of comparable EATE treatment effects. Coverage of the EATE interval estimates was conservative, always exceeding 99%.
Actual Bias of RMSE of Coverage Avg CI Avg Width
Estimand N Value$/N$ Estimator Estimator of 95% CI Width of EATE CI
$\tau^{\textup{basic}}$ 250 0.40 -0.00 0.09 0.95 0.18 0.27
500 0.36 -0.00 0.06 0.95 0.12 0.22
$\tau^{\textup{tr}}$ 250 0.21 -0.00 0.11 0.94 0.20 0.29
500 0.16 -0.00 0.06 0.95 0.12 0.23
$\tau^{\textup{control}}$ 250 0.21 -0.00 0.10 0.93 0.20 0.28
500 0.18 -0.00 0.06 0.94 0.12 0.23
$\tau^{\textup{indirect}}$ 250 0.36 0.14 0.16 0.99 0.17 0.21
500 0.30 0.12 0.13 1.0 0.10 0.16
$\tau^{\textup{nonneighbor}}$ 250 0.36 0.15 0.18 1.0 0.27 0.21
500 0.35 0.14 0.15 1.0 0.24 0.17
Table 3: Results of Simulation Study. Bold values indicate approximate 95% coverage, or that CI widths (measured as distance from point estimate to interval boundary) were smaller than those of comparable treatment effect.
6.2 Social Networks and the Decision to Insure
We analyze an experiment that was originally described in Cai et al., (2015), quoting first a summary of the experiment from Choi, (2021):
In an experiment described in Cai et al., (2015), rural farmers in China were randomly assigned to information sessions where they would be given the opportunity to purchase weather insurance. The sessions were randomized to give either high or low levels of information about the insurance product. First round sessions were held three days before second round sessions, so that first round attendees would have opportunity for informal conversations with their second round friends, in which they might share their opinions about the insurance product. Social network information was elicited, with the farmers instructed to list 5 close friends with whom they specifically discussed rice production or financial issues.
The goal of the experiment was to broadly demonstrate the importance of information sharing, by measuring its effects in a randomized setting. One of the conclusions of Cai et al., (2015) was that the decision to purchase insurance was affected not only by a farmer’s own treatment assignment, but also by that of their friends; specifically, farmers assigned to a second round low-information session were more likely to purchase insurance if more of their listed friends in the first round were assigned to a high-information session.
In the original analysis of the experiment, the unit outcomes were assumed to depend only on the information received by themselves and their listed friends. To justify this assumption, a subjective argument was made that the three day gap between information sessions was too short for information to travel more than one hop on the observed social network.
In our analysis, we forgo such assumptions and estimate $\tau^{\textup{control}}$, the number of second round units whose outcome under direct control (i.e., being assigned to a low information session) would be affected by the assignment of first round farmers into high and low information sessions. To construct our estimators, we require knowledge of the randomization design, which is contained in Gao and Ding, (2023).
Our point estimate equals 23%, which may be interpreted as meaning that at least 23% of second round farmers, if assigned to a low information session, would be affected by information given to the first round farmers (1-sided 95% CI: at least 9%). This point estimate of 23% is equal to the Hájek weighted comparison of second round units who received low information directly but had many first round friends with high information, versus those in the second round who received low information directly and had few or no first round friends with high information:
$\hat{\tau}=\frac{1}{N}\sum_{i=1}^{N}\left(\frac{1\{X_{i}=0,W_{i}=1\}N}{P(X_{i}=0,W_{i}=1)\hat{N}_{1}}Y_{i}-\frac{1\{X_{i}=0,W_{i}=0\}N}{P(X_{i}=0,W_{i}=0)\hat{N}_{0}}Y_{i}\right)$
where $i\in[N]$ enumerates the second round units, $X_{i}=0$ if unit $i$ was assigned to a low information session, $W_{i}=1$ if all of unit $i$’s first round friends received high information, $Y_{i}$ denotes unit $i$’s decision of whether or not to purchase insurance, and $\hat{N}_{a}=\sum_{i=1}^{N}1\{X_{i}=0,W_{i}=a\}/P(X_{i}=0,W_{i}=a)$ for $a\in\{0,1\}$.
For comparison, we consider an EATE-type treatment effect that considers the relative effects of receiving $(X_{i},W_{i})$ equal to $(0,1)$ versus $(0,0)$:
treatment $\textup{treatment effect}=\frac{1}{N}\sum_{i=1}^{N}\left(\mathbb{E}[Y_{i}|X_{i}=0,W_{i}=1]-\mathbb{E}[Y_{i}|X_{i}=0,W_{i}=0]\right),$
where the expectation is taken over the randomization of treatment under the experiment design. For this target parameter, the estimation method of Gao and Ding, (2023) gives a Hájek-normalized point estimate of 23%, and 95% CI of $[2\%,45\%]$. This confidence interval requires an assumption of no interference between villages. Such an assumption might be debatable, as farmers listed cross-village friendships. In contrast, no assumptions on interference are required for our estimand. Thus for the purposes of demonstrating the presence of social influence, our estimand may be an appropriate target parameter, and has tighter, less questionable CIs when compared to an analogous treatment effect.
Appendix
Appendix A Proofs
The organization of this section is as follows. In Section A.1, we present helper lemmas that will be used in our proofs. Section A.2 proves Theorems 1 - 6. Finally, in Section A.3 we prove Lemmas 1 - 9.
A.1 Auxilliary Results
The Azuma-Hoeffding (or McDiarmid’s) inequality (Boucheron et al., , 2013, Thm 6.2) states that given independent variables $U=(U_{1},\ldots,U_{N})$, and a function $f$ satisfying the bounded difference property
$|f(U)-f(U^{\prime})|\leq c_{i}\text{ if }U_{j}=U_{j}^{\prime}\text{ for all }j\neq i,\qquad i\in[N],$ (52)
it holds that
$\mathbb{P}\left(|f(U)-\mathbb{E}f|>\epsilon\right)\leq 2\exp\left(-2\epsilon^{2}/\sum_{i}c_{i}^{2}\right).$ (53)
Lemma 1 is used to prove Theorem 4. It applies McDiarmid’s inequality to show concentration of the terms appearing in $\hat{\Phi}_{k}^{\textup{Haj}}$ for $k=1,2$.
Lemma 1.
Let Assumption 1 hold. For $i\in[N]$ and $\bar{Z}_{i}$ given by (39), let $u_{i}$ denote the vector
$u_{i}=\left(\frac{1\{\bar{Z}_{i}=1\}}{P(\bar{Z}_{i}=1|T_{i})}\ ,\ \frac{1\{\bar{Z}_{i}=1\}}{P(\bar{Z}_{i}=1|T_{i})}Y_{i}\phi_{i}^{*}\ ,\ \frac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=0|T_{i})}\ ,\ \frac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=0|T_{i})}(1-Y_{i})\phi_{i}^{*}\right)$
and let $\bar{u}=\frac{1}{N}\sum_{i=1}^{N}u_{i}$. Then $\|\bar{u}-\mathbb{E}\bar{u}\|_{\infty}=O_{P}(N^{-1/2})$.
In our proof of Theorem 5, we will use $C_{ij}(T_{i},T_{j})$ to abbreviate the covariance estimates appearing in $V_{k}$. It is given by
$C_{ij}(T_{i},T_{j})=\mathbb{E}[v_{i}v_{j}|T_{i},T_{j}]-\mathbb{E}[v_{i}|T_{i}]\mathbb{E}[v_{j}|T_{j}]\frac{P(T_{i})P(T_{j})}{P(T_{i},T_{j})}$ (54)
Lemmas 2 and 3 are used to prove (47) of Theorem 5. Lemma 2 bounds the values of $\{C_{ij}\}$.
Lemma 2.
Let Assumption 1 hold and let $C_{ij}$ be given by (54). Let $\mathcal{I}_{0}$ and $\mathcal{I}_{1}$ denote the sets
$\mathcal{I}_{0}$ =$=\{i,j:\phi_{i}^{*}\phi_{j}^{*}=1,|\Gamma_{i}\cap\Gamma_{j}|=0\}$
$\mathcal{I}_{1}$ =$=\{i,j:\phi_{i}^{*}\phi_{j}^{*}=1,|\Gamma_{i}\cap\Gamma_{j}|\neq 0\}$
It holds w.p. 1 that
$\max_{i,j\in\mathcal{I}_{0}}C_{ij}(T_{i},T_{j})\leq O\left(\frac{\log N}{N}\right)$ (55)
$\max_{i,j\in\mathcal{I}_{1}}C_{ij}(T_{i},T_{j})\leq 1+O\left(\frac{\log N}{N}\right)$ (56)
To bound the size of the IPW weights appearing in the variance estimates, Lemma 3 states that certain events, which depend on only a finite number of treatments, cannot have probability that is infinitesimally small yet positive.
Lemma 3.
Let Assumption 1 hold, and let $\mathcal{T}_{i}$ denote the support of $T_{i}$. Then there exists a fixed constant $\rho_{\min}>0$ such that
$P(T_{i}=t_{i},T_{j}=t_{j})$$\in\{0\}\cap[\rho_{\min},1]\qquad\forall\ i,j\in[N]^{2}$ (57)
$P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})$$\in\{0\}\cap[\rho_{\min},1]\qquad\forall\ i,j:\phi_{i}^{*}\phi_{j}^{*}=1$ (58)
Lemmas 4 and 5 are used to prove (48) of Theorem 5.
Lemma 4.
Let $C_{ij}$ be given by (54). If $\phi_{i}^{*}\phi_{j}^{*}=1$ and $P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})=0$, then $C_{ij}(T_{i},T_{j})\leq 0$
Lemma 5.
It holds that $\mathbb{E}[C_{ij}(T_{i},T_{j})]\geq{\operatorname{Cov}}(v_{i},v_{j})$.
Lemmas 6 and 7 bound terms appearing in our variance estimators, and will be used to prove Lemma 2.
Lemma 6.
Let Assumption 1 hold, and let $v_{i}$ for $i\in[N]$ denote
$v_{i}=\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{\bar{Z}_{i}=Y_{ki}\}}{P(\bar{Z}_{i}=Y_{ki}|T_{i})},$
where $\hat{N}_{a}=\sum_{j=1}^{N}1\{\bar{Z}_{i}=a\}/P(\bar{Z}_{i}=a|T_{i})$ for $a\in\{0,1\}$. Then
$\mathbb{E}\left[v_{i}|T_{i}\right]$ $=\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|T_{i},\bar{Z}_{i}=Y_{ki}\right].$ (59)
If $\phi_{i}^{*}\phi_{j}^{*}=1$ and $|\Gamma_{i}\cap\Gamma_{j}|=0$, then
$\mathbb{E}\left[v_{i}v_{j}|T_{i},T_{j}\right]$ $=\mathbb{E}\left[\frac{N^{2}}{\hat{N}_{Y_{ki}}\hat{N}_{Y_{kj}}}|T_{i},T_{j},\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\right].$ (60)
If $\phi_{i}^{*}\phi_{j}^{*}=1$ and $|\Gamma_{i}\cap\Gamma_{j}|\neq 0$, then
$\mathbb{E}\left[v_{i},v_{j}|T_{i}T_{j}\right]$ $\leq\frac{1}{p_{\min}^{2}}\mathbb{E}\left[\frac{N^{2}}{\hat{N}_{Y_{ki}}\hat{N}_{Y_{kj}}}|T_{i},T_{j},\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\right].$ (61)
Lemma 7.
Let Assumption 1 hold. Then
$\max_{i:\phi_{i}^{*}=1}\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|T_{i},\bar{Z}_{i}=Y_{ki}\right]$ $=1+O\left(\frac{\log N}{N}\right)$ (62)
$\max_{i,j:\phi_{i}^{*}\phi_{j}^{*}=1}\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|T_{i},\bar{Z}_{i}=Y_{ki},T_{j},\bar{Z}_{j}=Y_{kj}\right]$ $=1+O\left(\frac{\log N}{N}\right)$ (63)
$\max_{i,j:\phi_{i}^{*}\phi_{j}^{*}=1}{\operatorname{Var}\,}\left[\frac{N}{\hat{N}_{Y_{ki}}}|T_{i},\bar{Z}_{i}=Y_{ki},T_{j},\bar{Z}_{j}=Y_{kj}\right]$ $=O\left(\frac{1}{N}\right)$ (64)
Lemma 8 is used to prove Theorem 6. It gives a coverage result for $\hat{\Phi}_{k}$.
Lemma 8.
Let Assumption 1 hold, and let $\widetilde{V}_{k}$ be defined as in Theorem 6. It holds for $k=1,2$ that
$P\left(\Phi-\hat{\Phi}_{k}\geq z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{k}}\right)\leq\frac{\alpha}{2}+o(1)$
Lemmas 9 and 10 are used to prove Lemma 8. Lemma 9 gives a central limit theorem for a linearized version of $\Phi_{k}$. Lemma 10 is a corollary of Fatou’s lemma, and will be used to show that $V_{k}$ is conservative for the variance of the linearized version of $\Phi_{k}$. It appears as Exercise 3.2.4 in (Durrett, , 2019, 5th edition).
Lemma 9.
Let Assumption 1 hold. For $k\in[2]$, let $\mu_{1}$ and $\mu_{0}$ denote
$\mu_{1}$ $=\sum_{i:\phi_{i}^{*}=1}\mathbb{E}[\tilde{f}_{ki}(T_{i})],$ $\mu_{0}$ $=\sum_{i:\phi_{i}^{*}=1}\mathbb{E}[1-\tilde{f}_{ki}(T_{i})].$
where $\tilde{f}_{ki}(T_{i})$ for $i:\phi_{i}^{*}=1$ denotes
$\tilde{f}_{ki}(T_{i})=\begin{cases}\tilde{f}_{i}(T_{i})&\textup{ if $k=1$}\\ 1-\tilde{f}_{i}(T_{i})&\textup{ if $k=2$}\end{cases}$ (65)
Let $h_{i}$ for $i\in[N]$ denote
$h_{i}=\frac{1\{\bar{Z}_{i}=1\}}{P(\bar{Z}_{i}=1|T_{i})}Y_{ki}\phi_{i}^{*}-\mu_{1}+\frac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=0|T_{i})}(1-Y_{ki})\phi_{i}^{*}-\mu_{0}$
Let $\sigma^{2}={\operatorname{Var}\,}\sum_{i}h_{i}$, and suppose that $\sigma^{2}\geq N^{2/3+\delta}$ for any $\delta>0$. Then $\sum_{i}h_{i}/\sigma$ converges to a standard normal.
Lemma 10.
(Durrett, , 2019, Exercise 3.2.4) Let $X_{n}$ denote a sequence of random variables that converges in distribution to $X_{\infty}$, and let $g$ denote a non-negative continuous function. Then $\liminf\limits_{n\rightarrow\infty}\mathbb{E}g(X_{n})\geq\mathbb{E}g(X_{\infty})$
Theorem 7 is taken from Ross, (2011), and gives a bounded dependence central limit theorem. It will be used to prove the asymptotic normality result of Lemma 9.
Theorem 7.
(Ross, , 2011, Thm 3.6) Let $U_{1},\ldots,U_{n}$ be random variables such that $\mathbb{E}[U_{i}^{4}]\leq\infty$, $\mathbb{E}U_{i}=0$, with variance denoted by $\sigma^{2}={\operatorname{Var}\,}\sum_{i=1}^{N}U_{i}$, with dependency neighborhoods $N_{1},\ldots,N_{n}$, which are subsets of $[n]$ (which need not be disjoint), where $U_{j}$ and $U_{i}$ are independent if $U_{j}\notin N_{i}$. Let $D=\max_{i}|N_{i}|$. Let $W=\sum_{i=1}^{N}U_{i}/\sigma$. Then for $Z$ denoting a $N(0,1)$ random variable, and $d_{W}$ denoting Wasserstein metric, it holds that
$d_{W}(W,Z)\leq\frac{D^{3}}{\sigma^{3}}\sum_{i=1}^{n}\mathbb{E}\left[|U_{i}|^{3}\right]+\frac{\sqrt{28}D^{3/2}}{\sqrt{\pi}\cdot\sigma^{2}}\sqrt{\sum_{i=1}^{n}\mathbb{E}\left[X_{i}^{4}\right]}$
A.2 Proof of Theorems
A.2.1 Proof of Theorems 1, 2, and 3
Proof of Theorem 1.
If $\hat{\Phi}_{1}$ and $\hat{\Phi}_{2}$ are within $o_{P}(N)$ of $\Phi$, then $\min\left(\hat{\Phi}_{1}\,,\,\hat{\Phi}_{2}\right)$ is within $o_{P}(N)$ as well. As $\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{k}(\phi)$ upper bounds $\hat{\Phi}_{k}$, it follows that
$\min\left(\max_{\phi\in\{0,1\}}\hat{\Phi}_{1}(\phi)\,,\,\max_{\phi\in\{0,1\}}\hat{\Phi}_{2}(\phi)\right)\geq\Phi+o_{P}(N),$
which implies that $\hat{\tau}\leq\tau+o_{P}(N)$. ∎
Proof of Theorem 2.
It can be seen that
$\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{1}(\phi)$ $=\max_{\phi\in\{0,1\}^{N}}\sum_{i:\phi_{i}=1}\left(\frac{X_{i}Y_{i}}{P(X_{i}=1)}+\frac{(1-X_{i})(1-Y_{i})}{P(X_{i}=0)}\right)\phi_{i}$
$=\sum_{i=1}^{N}\left(\frac{X_{i}Y_{i}}{P(X_{i}=1)}+\frac{(1-X_{i})(1-Y_{i})}{P(X_{i}=0)}\right)$
$=\sum_{i=1}^{N}\left(\frac{X_{i}Y_{i}}{P(X_{i}=1)}-\frac{(1-X_{i})Y_{i}}{P(X_{i}=0)}+\frac{1-X_{i}}{P(X_{i}=0)}\right)$
$=\sum_{i=1}^{N}\left(\frac{X_{i}}{P(X_{i}=1)}Y_{i}-\frac{1-X_{i}}{P(X_{i}=0)}Y_{i}\right)+o_{P}(N)$
Where the final equality holds by the assumptions of the theorem. By similar arguments, it holds that
$\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{2}(\phi)=\sum_{i=1}^{N}\left(\frac{1-X_{i}}{P(X_{i}=0)}Y_{i}-\frac{X_{i}}{P(X_{i}=1)}Y_{i}\right)+o_{P}(N)$
Substituting into $\hat{\tau}=N-\min\left(\max\limits_{\phi\in\{0,1\}}\hat{\Phi}_{1}(\phi)\ ,\ \max\limits_{\phi\in\{0,1\}}\hat{\Phi}_{2}(\phi)\right)$ yields
$\hat{\tau}=\left|\sum_{i=1}^{N}\frac{X_{i}}{P(X_{i}=1)}Y_{i}-\sum_{i=1}^{N}\frac{1-X_{i}}{P(X_{i}=0)}Y_{i}\right|,$
proving the result. ∎
Proof of Theorem 3.
Let $\tilde{f}_{ki}$ be given by (65). Since $\phi_{i}^{*}=1$ and $\bar{Z}_{i}=Y_{ki}$ (which implies $X\in R_{i}$), it holds that $Y_{ki}=\tilde{f}_{ki}(T_{i})$, and hence that
$\mathbb{E}[v_{i}|T_{i}]$ $=\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{\bar{Z}_{i}=Y_{ki}\}}{P(\bar{Z}_{i}=Y_{ki}|T_{i})}|T_{i}\right]$
$=P(X\in R_{i}|T_{i})\cdot\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{\bar{Z}_{i}=Y_{ki}\}}{P(\bar{Z}_{i}=Y_{ki}|T_{i})}|T_{i},X\in R_{i}\right]$ (66)
$=P(X\in R_{i}|T_{i})\cdot\sum_{a=0,1}1\{Y_{ki}=a\}\mathbb{E}\left[\frac{N}{\hat{N}_{a}}\cdot\frac{1\{\bar{Z}_{i}=a\}}{P(\bar{Z}_{i}=a|T_{i})}|T_{i},X\in R_{i}\right]$ (67)
$=\sum_{a=0,1}1\{Y_{ki}=a\}\cdot P(X\in R_{i}|T_{i})\mathbb{E}\left[\frac{N}{\hat{N}_{a}}\cdot\frac{1\{\bar{Z}_{i}=a\}}{P(\bar{Z}_{i}=a|T_{i})}|T_{i},X\in R_{i}\right]$ (68)
$=\sum_{a=0,1}1\{Y_{ki}=a\}\mathbb{E}\left[\frac{N}{\hat{N}_{a}}\cdot\frac{1\{\bar{Z}_{i}=a\}}{P(\bar{Z}_{i}=a|T_{i})}|T_{i}\right]$ (69)
where (66) holds since $01\{\bar{Z}_{i}=Y_{ki}\}=$ 0 if $X\notin R_{i}$; (67) holds since the $Y_{ki}$ term in the conditional expectation equals $\tilde{f}_{ki}(T_{i})$, and hence may be treated as a constant equal to the observed value of $Y_{ki}$; (68) is a simple rearrangement of terms; and (69) uses that $01\{\bar{Z}_{i}=a\}=$ 0 for $a\in\{0,1\}$ if $X\notin R_{i}$.
By similar reasoning, since $\phi_{i}^{*}\phi_{j}^{*}=1$ and $11\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}=$ 1 (which implies that $X\in R_{i}\cap R_{j}$), it holds that $Y_{ki}=\tilde{f}_{ki}(T_{i})$ and $Y_{kj}=\tilde{f}_{kj}(T_{j})$, and hence that
$\mathbb{E}[v_{i}v_{i}|T_{i},T_{j}]$ $=\mathbb{E}\left[\frac{1\{\bar{Z}_{i}=Y_{ki}\}N}{P(\bar{Z}_{i}=Y_{ki}|T_{i})\hat{N}_{Y_{ki}}}\cdot\frac{1\{\bar{Z}_{i}=Y_{ki}\}N}{P(\bar{Z}_{i}=Y_{ki}|T_{i})\hat{N}_{Y_{ki}}}|T_{i},T_{j}\right]$
$=P(X\in R_{i}\cap R_{j})\mathbb{E}\left[\frac{1\{\bar{Z}_{i}=Y_{ki}\}N}{P(\bar{Z}_{i}=Y_{ki}|T_{i})\hat{N}_{Y_{ki}}}\cdot\frac{1\{\bar{Z}_{i}=Y_{ki}\}N}{P(\bar{Z}_{i}=Y_{ki}|T_{i})\hat{N}_{Y_{ki}}}|T_{i},T_{j},X\in R_{i}\cap R_{j}\right]$
$=P(X\in R_{i}\cap R_{j})\sum_{a=0}^{1}\sum_{b=0}^{1}1\{Y_{ki}=a,Y_{kj}=b\}\cdot$
$\hskip 28.45274pt\mathbb{E}\left[\frac{1\{\bar{Z}_{i}=a\}N}{P(\bar{Z}_{i}=a|T_{i})\hat{N}_{a}}\cdot\frac{1\{\bar{Z}_{i}=b\}N}{P(\bar{Z}_{i}=b|T_{i})\hat{N}_{b}}|T_{i},T_{j},X\in R_{i}\cap R_{j}\right]$
$=\sum_{a=0}^{1}\sum_{b=0}^{1}1\{Y_{ki}=a,Y_{kj}=b\}\cdot$
$\hskip 28.45274ptP(X\in R_{i}\cap R_{j})\mathbb{E}\left[\frac{1\{\bar{Z}_{i}=a\}N}{P(\bar{Z}_{i}=a|T_{i})\hat{N}_{a}}\cdot\frac{1\{\bar{Z}_{i}=b\}N}{P(\bar{Z}_{i}=b|T_{i})\hat{N}_{b}}|T_{i},T_{j},X\in R_{i}\cap R_{j}\right]$
$=\sum_{a=0}^{1}\sum_{b=0}^{1}1\{Y_{ki}=a,Y_{kj}=b\}\mathbb{E}\left[\frac{1\{\bar{Z}_{i}=a\}N}{P(\bar{Z}_{i}=a|T_{i})\hat{N}_{a}}\cdot\frac{1\{\bar{Z}_{i}=b\}N}{P(\bar{Z}_{i}=b|T_{i})\hat{N}_{b}}|T_{i},T_{j}\right]$
and
$P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})$ $=P(X\in R_{i}\cap R_{j}|T_{i},T_{j})P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j},X\in R_{j}\cap R_{j})$
$=P(X\in R_{i}\cap R_{j}|T_{i},T_{j})\cdot\sum_{a=0}^{1}\sum_{b=0}^{1}1\{Y_{ki}=a,Y_{kj}=b\}\cdot$
$\hskip 28.45274ptP(\bar{Z}_{i}=a,\bar{Z}_{j}=b|T_{i},T_{j},X\in R_{j}\cap R_{j})$
$=\sum_{a=0}^{1}\sum_{b=0}^{1}1\{Y_{ki}=a,Y_{kj}=b\}\cdot$
$\hskip 28.45274ptP(X\in R_{i}\cap R_{j}|T_{i},T_{j})P(\bar{Z}_{i}=a,\bar{Z}_{j}=b|T_{i},T_{j},X\in R_{j}\cap R_{j})$
$=\sum_{a=0}^{1}\sum_{b=0}^{1}1\{Y_{ki}=a,Y_{kj}=b\}\cdot P(\bar{Z}_{i}=a,\bar{Z}_{j}=b|T_{i},T_{j})$
A.2.2 Proof of Theorem 4
We prove (44), (45), and (46) separately, as follows:
Proof of Theorem 4, Eq. (44).
Proof of (44) follows by Taylor approximation and is more or less standard. Given $k\in[2]$, let $a,b\in\mathbb{R}^{2}$ be given by
[…]
To evaluate $\mathbb{E}a$, we observe that for $a\in\{0,1\}$, $\phi_{i}^{*}=1$, and $\tilde{f}_{ki}(T_{i})$ given by (65), it holds that
$\mathbb{E}\left[\frac{1\{\bar{Z}_{i}=a\}}{P(\bar{Z}_{i}=a|T_{i})}Y_{ki}\right]$ $=\mathbb{E}\left[\mathbb{E}\left\{\frac{1\{\bar{Z}_{i}=a\}}{P(\bar{Z}_{i}=a|T_{i})}Y_{ki}|T_{i}\right\}\right]$
$=\mathbb{E}\left[\mathbb{E}\left\{Y_{ki}|T_{i},\bar{Z}_{i}=a\right\}\right]$
$=\mathbb{E}\left[\mathbb{E}\left\{\tilde{f}_{ki}(T_{i})|T_{i},\bar{Z}_{i}=a\right\}\right]$ (77)
$=\mathbb{E}\left\{\mathbb{E}\left[\tilde{f}_{ki}(T_{i})|T_{i}\right]\right\}$
$=\mathbb{E}[\tilde{f}_{ki}(T_{i})]$
where (77) hold since $Y_{ki}=\tilde{f}_{ki}(T_{i})$ if $\phi_{i}^{*}=1$ and $X\in R_{i}$, which is implied by $\bar{Z}_{i}\in\{0,1\}$. It follows that $\mathbb{E}a$ and $\mathbb{E}b$ equal
$\mathbb{E}a$ $=\left[\begin{array}[]{l}\dfrac{1}{N}\displaystyle\sum_{i:\phi_{i}^{*}=1}\mathbb{E}[\tilde{f}_{ki}(T_{i})]\\ \dfrac{1}{N}\displaystyle\sum_{i:\phi_{i}^{*}=1}\mathbb{E}[1-\tilde{f}_{ki}(T_{i})]\end{array}\right],$ $\mathbb{E}b=\left[\begin{array}[]{l}1\\ 1\end{array}\right]$ (82)
and that $\hat{\Phi}_{k}^{\textup{Haj}}$ and $\Phi$ are given by
$\hat{\Phi}_{k}^{\textup{Haj}}$ $=N\left(\frac{a_{1}}{b_{1}}+\frac{a_{2}}{b_{2}}\right),$ $\Phi$ $=N\left(\frac{\mathbb{E}a_{1}}{\mathbb{E}b_{1}}+\frac{\mathbb{E}a_{2}}{\mathbb{E}b_{2}}\right)$
and the estimation error may be written as
$\frac{1}{N}\left(\hat{\Phi}_{k}^{\textup{Haj}}-\Phi\right)$ $=\frac{\epsilon_{1}}{b_{1}}+\frac{\epsilon_{2}}{b_{2}}$
where $\epsilon\in\mathbb{R}^{2}$ denotes
$\epsilon=\left[\begin{array}[]{l}a_{1}-b_{1}\dfrac{\mathbb{E}a_{1}}{\mathbb{E}b_{1}}\\ \\ a_{2}-b_{2}\dfrac{\mathbb{E}a_{2}}{\mathbb{E}b_{2}}\end{array}\right]$ (86)
which can be seen to satisfy $\mathbb{E}\epsilon=0$.
To show (44), we define the mapping $h$ by
$h(\epsilon,b)=\frac{\epsilon_{1}}{b_{1}}+\frac{\epsilon_{2}}{b_{2}}$
so that $N^{-1}(\hat{\Phi}_{k}^{\textup{Haj}}-\Phi)=h(\epsilon,b)$. Letting $\xi=(\epsilon,b)$, by mean value theorem it holds for some $\tilde{\xi}$ in the convex hull of $\xi$ and $\mathbb{E}\xi$ that
$h(\xi)-h(\mathbb{E}\xi)$ $=\nabla h(\tilde{\xi})^{T}(\xi-\mathbb{E}\xi)$
$=\nabla h(\mathbb{E}\xi)^{T}(\xi-\mathbb{E}\xi)+\left(\nabla h(\tilde{\xi})-\nabla h(\mathbb{E}\xi)\right)^{T}(\xi-\mathbb{E}\xi)$ (87)
where the derivative $\nabla h(\xi)$ is given by
$\nabla h(\xi)=\left[\dfrac{1}{b_{1}}\ \ \ \dfrac{1}{b_{2}}\ \ -\frac{\epsilon_{1}}{b_{1}^{2}}\ \ -\frac{\epsilon_{2}}{b_{2}^{2}}\right]$ (88)
As $\xi$ is an affine transformation of $\bar{u}$ defined in Lemma 1, it follows that
$\|\xi-\mathbb{E}\xi\|_{\infty}=O_{P}(N^{-1/2})$ (89)
from which it follows that
$\nabla h(\mathbb{E}\xi)^{T}(\xi-\mathbb{E}\xi)=O_{P}(N^{-1/2})$ (90)
where we have used $\nabla h(\mathbb{E}\xi)=\left(1,1,0,0\right)$.
As $\tilde{\xi}$ is in the convex hull of $\xi$ and $\mathbb{E}\xi$, (89) implies that $\|\tilde{\xi}-\mathbb{E}\xi\|=O_{P}(N^{-1/2})$. By continuity of $\nabla h$ at $\mathbb{E}\xi$, it follows that $\|\nabla h(\tilde{\xi})-\nabla h(\mathbb{E}\xi)\|=O_{P}(N^{-1/2})$ as well, and hence that
$\left(\nabla h(\tilde{\xi})-\nabla h(\mathbb{E}\xi)\right)^{T}(\xi-\mathbb{E}\xi)=O_{P}(N^{-1})$ (91)
Substituting (90) and (91) into (87) proves that $N^{-1}(\hat{\Phi}_{k}-\Phi)=O_{P}(N^{-1/2})$, which is equivalent to the claim of (44) made by the lemma. ∎
Proof of Theorem 4, Eq. (45).
By (44), it holds that $\hat{\Phi}_{1}^{\textup{Haj}}$ and $\hat{\Phi}_{2}^{\textup{Haj}}$ are within $O_{P}(N^{1/2})$ of $\Phi$, and hence $\min\left(\hat{\Phi}_{1}^{\textup{Haj}}\,,\,\hat{\Phi}_{2}^{\textup{Haj}}\right)$ is within $O_{P}(N^{1/2})$ as well. As $\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{k}(\phi)$ upper bounds $\hat{\Phi}_{k}$, it follows that
$\min\left(\max_{\phi\in\{0,1\}}\hat{\Phi}_{1}^{\textup{Haj}}(\phi)\,,\,\max_{\phi\in\{0,1\}}\hat{\Phi}_{2}^{\textup{Haj}}(\phi)\right)\geq\Phi+O_{P}(N^{1/2}),$
which implies that $\hat{\tau}\leq\tau+O_{P}(N^{1/2})$, proving (45). ∎
Proof of Theorem 4, Eq. (46).
It holds that
$\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{k}^{\textup{Haj}}(\phi)$ $=\sum_{i=1}^{N}\frac{N}{\hat{N}_{1}}\cdot\frac{1\{\bar{Z}_{i}=1\}}{P(\bar{Z}_{i}=1|T_{i})}\cdot Y_{i}+\sum_{i=1}^{N}\frac{N}{\hat{N}_{0}}\cdot\frac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=1|T_{i})}\cdot(1-Y_{i})$
$=\sum_{i=1}^{N}\frac{N}{\hat{N}_{1}}\cdot\frac{1\{\bar{Z}_{i}=1\}}{P(\bar{Z}_{i}=1|T_{i})}\cdot Y_{i}-\sum_{i=1}^{N}\frac{N}{\hat{N}_{0}}\cdot\frac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=1|T_{i})}\cdot Y_{i}+N$ (92)
=$=\begin{cases}N+\Delta^{{\textup{Haj}}}&\textup{ if }k=1\\ N-\Delta^{{\textup{Haj}}}&\textup{ if }k=2\end{cases}$ (93)
where (92) holds since the definition of $\hat{N}_{0}$ implies that
$\sum_{i=1}^{N}\frac{N}{\hat{N}_{0}}\cdot\frac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=0|T_{i})}=1,$
and $\Delta^{{\textup{Haj}}}$ denotes
$\Delta^{{\textup{Haj}}}=\sum_{i=1}^{N}\frac{N}{\hat{N}_{1}}\frac{1\{\bar{Z}_{i}=1\}}{P(\bar{Z}_{i}=1|T_{i})}Y_{i}-\sum_{i=1}^{N}\frac{N}{\hat{N}_{0}}\frac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=0|T_{i})}Y_{i},$
for $\hat{N}_{a}$ defined as in the statement of the theorem.
To show (46), we observe that
$\hat{\tau}$ $=N-\min\left(\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{1}^{\textup{Haj}}(\phi)\ ,\ \max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{2}^{\textup{Haj}}(\phi)\right)$
$=\max\left(N-\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{1}^{\textup{Haj}}(\phi)\ ,\ N-\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{2}^{\textup{Haj}}(\phi)\right)$
$=\max\left(-\Delta^{{\textup{Haj}}},\Delta^{{\textup{Haj}}}\right)$
$=\left|\Delta^{{\textup{Haj}}}\right|$
where the inequality follows by substitution of (93). ∎
A.2.3 Proof of Theorem 5
Proof of Theorem 5, Eq. (47).
For $C_{ij}$ given by (54), let $M\in\mathbb{R}^{N\times N}$ denote the matrix given by
$M_{ij}=\begin{cases}C_{ij}\cdot\dfrac{1\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}}{P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})}&\textup{ if $\phi_{i}^{*}\phi_{j}^{*}=1$}\\ 0&\textup{ else }\end{cases}$
so that $V_{k}=\sum_{ij}M_{ij}$.
By Lemma 2 and (58) of Lemma 3, $C>0$ exists such that $|M_{ij}|$ may be bounded by
$|M_{ij}|\leq\begin{cases}\frac{C\log N}{N\rho_{\min}}&i,j\in\mathcal{I}_{0}\\ \frac{C}{\rho_{\min}}&i,j\in\mathcal{I}_{1}.\end{cases}$ (94)
By Assumption 1, it holds that
$\max_{i\in[N]}|\{j:(i,j)\in\mathcal{I}_{1}\}|\leq d_{\max}.$ (95)
Combining (94) and (95) implies that each row or column of $M$ has entries summing to at most $C(\log N)/\rho_{\min}+Cd_{\max}/\rho_{\min}$.
Changing $X$ in a single entry changes at most $d_{\max}$ rows and columns of $M$. As a result, for a fixed constant $C_{2}>0$ the bounded differences condition holds for $N\geq 2$ and all $X,X^{\prime}$ differing in a single entry that
$|V_{k}(X)-V_{k}(X^{\prime})|\leq C_{2}\log N$
and it holds by McDiarmid’s inequality that
$P(|V_{k}-\mathbb{E}V_{k}|\geq\epsilon)\leq 2\exp\left(\frac{-2\epsilon^{2}}{C_{2}N(\log N)^{2}}\right)$
which implies that $|V-\mathbb{E}V_{k}|=O_{P}(N^{1/2+\delta})$ for any $\delta>0$. ∎
Proof of Theorem 5, Eq. (48).
Recall that $C_{ij}$ is given by (54). To show that $V_{k}$ is not anti-conservative, we observe that
$\mathbb{E}V_{k}$ $=\sum_{i,j:\phi_{i}^{*}\phi_{j}^{*}=1}\ \mathbb{E}\left[\mathbb{E}\left[C_{ij}(T_{i},T_{j})\cdot\frac{1\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}}{P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})}\ \bigg{|}T_{i},T_{j}\right]\right]$
$=\sum_{i,j:\phi_{i}^{*}\phi_{j}^{*}=1}\ \mathbb{E}\left[C_{ij}(T_{i},T_{j})\cdot\mathbb{E}\left[\frac{1\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}}{P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})}\ \bigg{|}T_{i},T_{j}\right]\right]$
$=\sum_{i,j:\phi_{i}^{*}\phi_{j}^{*}=1}\ \sum_{t_{i},t_{j}}P(T_{i}=t_{i},T_{j}=t_{j})\cdot C_{ij}(t_{i},t_{j})\cdot$
$\hskip 28.45274pt1\{P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i}=t_{i},T_{j}=t_{j})\neq 0\}$
$\geq\sum_{i,j:\phi_{i}^{*}\phi_{j}^{*}=1}\ \sum_{t_{i},t_{j}}P(T_{i}=t_{i},T_{j}=t_{j})C_{ij}(t_{i},t_{j})$ (96)
$\geq\sum_{i,j:\phi_{i}^{*}\phi_{j}^{*}=1}\ {\operatorname{Cov}}(v_{j},v_{j})$ (97)
where (96) holds by Lemma 4, and (97) holds by Lemma 5. It can be seen that (97) equals the claim of (48) made by Theorem 5.
To verify the conditions for unbiasedness given by Theorem 5, we observe that if $T_{i}$ is the constant mapping for all units, then $V_{k}$ simplifies to
$V_{k}=\sum_{i:\phi^{*}=1}\sum_{j:\phi^{*}=1}{{\operatorname{Cov}}}\left(v_{i},v_{j}\right)\cdot\frac{1\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}}{P\left(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\right)}$
As this is an importance sampled estimator of covariance terms appearing in ${\operatorname{Var}\,}\Phi_{k}^{{\textup{Haj}}}$, it is unbiased provided that the sampling probabilities $P\left(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\right)$ are non-zero for each term.
A.2.4 Proof of Theorem 6
Proof of Theorem 6.
It holds that
$CI-\tau$ $=\Phi-\min\left\{\max_{\phi\in\{0,1\}}\hat{\Phi}_{1}(\phi)+z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{1}(\phi)}\ ,\ \max_{\phi\in\{0,1\}}\hat{\Phi}_{2}(\phi)+z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{2}(\phi)}\right\}$
$\leq\Phi-\min\left\{\hat{\Phi}_{1}+z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{1}}\ ,\ \hat{\Phi}_{2}+z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{2}}\right\}$
$=\max\left\{\Phi-\hat{\Phi}_{1}-z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{1}}\ ,\ \Phi-\hat{\Phi}_{2}-z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{2}}\right\}$ (98)
Let $\mathcal{E}_{k}$ denote the event
$\mathcal{E}_{k}=\left\{\Phi-\hat{\Phi}_{k}-z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{k}}\geq 0\right\}.$
By (98), it follows that loss of coverage occurs if either $\mathcal{E}_{1}$ or $\mathcal{E}_{2}$ occurs. By Lemma 8, the probability of $\mathcal{E}_{k}$ is bounded by $\alpha/2+o(1)$, and hence by union bound it holds that
$P(\mathcal{E}_{1}\cup\mathcal{E}_{2})\leq\alpha+o(1)$ (99)
Combining (98) and (99) yields that
$P(CI-\tau>0)\leq\alpha+o(1)$ (100)
proving that $CI\leq\tau$ with probability converging to at least $1-\alpha$. ∎
A.3 Proof of Auxilliary Lemmas
Proof of Lemma 1.
For each $i\in[N]$, it can be seen that $\|u_{i}\|\leq 1/p_{\min}$ and that $u_{i}$ depends on $X$ only through $\bar{Z}_{i}$ and $T_{i}$, and hence only through $X_{\Gamma_{i}}$. As a result, changing $X$ in a single entry changes at most $d_{\max}$ entries of $u_{1},\ldots,u_{N}$, and hence changes $\bar{u}$ by at most $C=d_{\max}/p_{\min}$. Applying McDiarmid’s inequality to the $\ell$th entry of $\bar{u}$ for yields that
$P(|\bar{u}_{\ell}-\mathbb{E}\bar{u}_{\ell}|\geq\epsilon)\leq\exp\left(-2\epsilon^{2}/NC\right)$
and hence $\bar{u}_{\ell}-\mathbb{E}\bar{u}_{\ell}=O_{P}(N^{-1/2})$, and $\|\bar{u}-\mathbb{E}\bar{u}\|_{\infty}=O_{P}(N^{-1/2})$ as well. ∎
Proof of Lemma 2.
For $i,j$ satisfying $\phi_{i}^{*}\phi_{j}^{*}=1$, let $F_{\{i\}}$ and $F_{\{i,j\}}$ denote sigma-fields such that
$\mathbb{E}[\ \cdot\ |F_{\{i\}}]$ $=\mathbb{E}[\ \cdot\ |T_{i},\bar{Z}_{i}=Y_{ki}]$
$\mathbb{E}[\ \cdot\ |F_{\{i,j\}}]$ $=\mathbb{E}[\ \cdot\ |T_{i},T_{j},\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}]$
Recall $C_{ij}$ is given by (54). To show (55), we observe that for $i,j\in\mathcal{I}_{0}$ that
$C_{ij}(T_{i},T_{j})$ $=\mathbb{E}[v_{i}v_{j}|T_{i},T_{j}]-\mathbb{E}[v_{i}|T_{i}]\mathbb{E}[v_{j}|T_{j}]\frac{P(T_{i})P(T_{j})}{P(T_{i},T_{j})}$
$=\mathbb{E}\left[\frac{N^{2}}{\hat{N}_{Y_{ki}}\hat{N}_{Y_{kj}}}|F_{\{i,j\}}\right]-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i\}}\right]\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{j\}}\right]$ (101)
$=\mathbb{E}\left[\frac{N^{2}}{\hat{N}_{Y_{ki}}\hat{N}_{Y_{kj}}}|F_{\{i,j\}}\right]-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i,j\}}\right]\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{i,j\}}\right]$
$\hskip 28.45274pt+\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i\}}\right]\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{j\}}\right]-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i,j\}}\right]\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{i,j\}}\right]$
$={\operatorname{Cov}}\left(\frac{N}{\hat{N}_{Y_{ki}}},\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{i,j\}}\right)+\left(\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i\}}\right]-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i,j\}}\right]\right)\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{j\}}\right]$
$\hskip 28.45274pt+\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i,j\}}\right]\left(\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{j\}}\right]-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{i,j\}}\right]\right)$
$\leq\sqrt{{\operatorname{Var}\,}\left(\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i,j\}}\right){\operatorname{Var}\,}\left(\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{i,j\}}\right)}$
$\hskip 28.45274pt+\left(\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i\}}\right]-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i,j\}}\right]\right)\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{j\}}\right]$
$\hskip 28.45274pt+\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i,j\}}\right]\left(\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{j\}}\right]-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{i,j\}}\right]\right)$
$\leq O\left(\frac{1}{N}\right)+O\left(\frac{\log N}{N}\right)$ (102)
where (101) uses Lemma 6, and (102) uses Lemma 7.
To show (56), we observe for $i,j\in\mathcal{I}_{1}$ that
$C_{ij}(T_{i},T_{j})$ $=\mathbb{E}[v_{i}v_{j}|T_{i},T_{j}]-\mathbb{E}[v_{i}|T_{i}]\mathbb{E}[v_{j}|T_{j}]\frac{P(T_{i})P(T_{j})}{P(T_{i},T_{j})}$
$\leq\frac{1}{p_{\min}^{2}}\mathbb{E}\left[\frac{N^{2}}{\hat{N}_{Y_{ki}}\hat{N}_{Y_{kj}}}|F_{\{i,j\}}\right]-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i\}}\right]\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{j\}}\right]\cdot\frac{1}{\rho_{\min}}$ (103)
$\leq\frac{1}{p_{\min}^{2}}\left(\mathbb{E}\left[\frac{N^{2}}{(\hat{N}_{Y_{ki}})^{2}}|F_{\{i,j\}}\right]\cdot\mathbb{E}\left[\frac{N^{2}}{(\hat{N}_{Y_{kj}})^{2}}|F_{\{i,j\}}\right]\right)^{1/2}+$
$\hskip 28.45274pt\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i\}}\right]\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{j\}}\right]\cdot\frac{1}{\rho_{\min}}$
$=O\left(1\right)+O\left(\frac{\log N}{N}\right)$ (104)
where (103) uses Lemma 6 and (57), and (104) uses Lemma 7 and also because
$\mathbb{E}\left[\frac{N^{2}}{(\hat{N}_{Y_{ki}})^{2}}|F_{\{i,j\}}\right]$ $={\operatorname{Var}\,}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i,j\}}\right]+\left(\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i,j\}}\right]\right)^{2}$
$=O(1/N)+O(1)$
Proof of Lemma 3.
Given $i,j\in[N]$, let $\mathcal{X}_{\Gamma_{i}\cup\Gamma_{j}}=\{0,1\}^{|\Gamma_{i}\cup\Gamma_{j}|}$ denote the space of possible treatment assignments for $X_{\Gamma_{i}\cup\Gamma_{j}}$. By Assumption 1, the unit treatments are independently assigned, with $p_{\min}\leq P(X_{i}=1)\leq 1-p_{\min}$. Since $|\Gamma_{i}|\leq d_{\max}$, it follows for $\rho_{\min}=(p_{\min})^{2d_{\max}}$ that
$P(X_{\Gamma_{i}\cup\Gamma_{j}}=x)\geq\rho_{\min}\qquad\textup{ for all $x\in\{0,1\}^{|\Gamma_{i}\cup\Gamma_{j}|}$}$ (105)
As $T_{i}$ and $T_{j}$ depends on $X$ only through $X_{\Gamma_{i}}$ and $X_{\Gamma_{j}}$, we may write
$P(T_{i}=t_{i},T_{j}=t_{j})$ $=P(X_{\Gamma_{i}\cup\Gamma_{j}}\in\mathcal{S})$
for some (possibly empty) set $\mathcal{S}\subseteq\mathcal{X}_{\Gamma_{i}\cup\Gamma_{j}}$. It follows that $P(T_{i}=t_{i},T_{j}=t_{j})$ equals zero if $\mathcal{S}$ is empty, and otherwise lower bounded by $\rho_{\min}$ by (105). This establishes
[…]
which equals (57).
To establish (58), we observe that similar to $T_{i}$, if $\phi_{i}^{*}=1$ then the occurrence of the event $\{\bar{Z}_{i}=Y_{ki}\}$ depends only on $X$ through $X_{\Gamma_{i}}$. It follows that
$P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj},T_{i},T_{j})=P(X_{\Gamma_{i}\cup\Gamma_{j}}\in\mathcal{S})$
for some (possibly empty) set $\mathcal{S}\subseteq\mathcal{X}_{\Gamma_{i}\cup\Gamma_{j}}$. Applying (105) yields
[…]
Combining (106) with Bayes’ rule, which implies
$P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})\geq P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj},T_{i},T_{j})$
with equality when $P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj},T_{i},T_{j})=0$, establishes that
[…]
proving the claim. ∎
Proof of Lemma 4.
Recall that $v_{i}$ and $C_{ij}$ are respectively given by (37) and (54). If $\phi_{i}^{*}\phi_{j}^{*}=1$ and $P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})=0$, then
$\mathbb{E}[v_{i}v_{j}|T_{i},T_{j}]$ $=\mathbb{E}\left[\frac{1\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}}{P(\bar{Z}_{i}=Y_{ki}|T_{i})P(\bar{Z}_{j}=Y_{kj}|T_{j})}\cdot\frac{N^{2}}{\hat{N}_{Y_{ki}}\hat{N}_{Y_{kj}}}|T_{i},T_{j}\right]$
$=0$ (107)
where (107) uses $P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})=0$. It follows that
$C_{ij}(T_{i},T_{j})$ $=\mathbb{E}\left[v_{i}v_{j}|T_{i},T_{j}\right]-\mathbb{E}[v_{i}|T_{i}]\cdot\mathbb{E}[v_{j}|T_{j}]\cdot\frac{P(T_{i})P(T_{j})}{P(T_{i},T_{j})}$
$=0-\mathbb{E}[v_{i}|T_{i}]\cdot\mathbb{E}[v_{j}|T_{j}]\cdot\frac{P(T_{i})P(T_{j})}{P(T_{i},T_{j})}$ (108)
$\leq 0$ (109)
where (108) holds by (107), and (109) holds because $v_{i}\geq 0$ always, which implies that $\mathbb{E}[v_{i}|T_{i}]\geq 0$. ∎
Proof of Lemma 5.
Let $\mathcal{T}_{i}$ denote the support of $T_{i}(X)$.
$\sum_{t_{i}\in\mathcal{T}_{i}}\sum_{t_{j}\in\mathcal{T}_{j}}P(T_{i}=t_{i},T_{j}=t_{j})C_{ij}(t_{i},t_{j})$
$=\sum_{t_{i}\in\mathcal{T}_{i}}\sum_{t_{j}\in\mathcal{T}_{j}}P(T_{i}=t_{i},T_{j}=t_{j})\bigg{\{}\mathbb{E}\left[v_{i}v_{j}|T_{i}=t_{i},T_{j}=t_{j}\right]-$
$\hskip 28.45274pt\mathbb{E}[v_{i}|T_{i}=t_{i}]\mathbb{E}[v_{j}|T_{j}=t_{j}]\cdot\frac{P(T_{i}=t_{i})P(T_{j}=t_{j})}{P(T_{i}=t_{i},T_{j}=t_{j})}\bigg{\}}$ (110)
$=\mathbb{E}[v_{i}v_{j}]-\sum_{t_{i}\in\mathcal{T}_{i},t_{j}\in\mathcal{T}_{j}}1\{P(T_{i}=t_{i},T_{j}=t_{j})\neq 0\}\cdot$
$\hskip 28.45274pt\mathbb{E}[v_{i}|T_{i}=t_{j}]\mathbb{E}[v_{j}|T_{j}=t_{j}]P(T_{i}=t_{i})P(T_{j}=t_{j})$ (111)
$\geq\mathbb{E}[v_{i}v_{j}]-\sum_{t_{i}\in\mathcal{T}_{i},t_{j}\in\mathcal{T}_{j}}\mathbb{E}[v_{i}|T_{i}=t_{i}]\mathbb{E}[v_{j}|T_{j}=t_{j}]P(T_{i}=t_{i})P(T_{j}=t_{j})$ (112)
$=\mathbb{E}[v_{i}v_{j}]-\mathbb{E}[v_{i}]\cdot\mathbb{E}[v_{j}]={\operatorname{Cov}}(v_{i},v_{j})$
where (110) holds by definition of $C_{ij}$ in (54), and (111) holds because
$\sum_{t_{i}\in\mathcal{T}_{i}}\sum_{t_{j}\in\mathcal{T}_{j}}P(T_{i}=t_{i},T_{j}=t_{j})\mathbb{E}\left[v_{i}v_{j}|T_{i}=t_{i},T_{j}=t_{j}\right]=\mathbb{E}[v_{i}v_{j}],$
and (112) holds because
$\mathbb{E}[v_{i}|T_{i}=t_{i}]\mathbb{E}[v_{j}|T_{j}=t_{j}]P(T_{i}=t_{i})P(T_{j}=t_{j})\geq 0$
since $v_{i}\geq 0$ always. ∎
Proof of Lemma 6.
Recall that $v_{i}$ is given by (37). To show (59), we observe that
$\mathbb{E}\left[v_{i}|T_{i}\right]$ $=\mathbb{E}\left[\frac{1\{\bar{Z}_{i}=Y_{ki}\}N}{P(\bar{Z}_{i}=Y_{ki}|T_{i})\hat{N}_{Y_{ki}}}|T_{i}\right]$
$=P(\bar{Z}_{i}=Y_{ki}|T_{i})\mathbb{E}\left[\frac{N}{P(\bar{Z}_{i}=Y_{ki}|T_{i})\hat{N}_{Y_{ki}}}|T_{i},\bar{Z}_{i}=Y_{ki}\right]$
$=\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|T_{i},\bar{Z}_{i}=Y_{ki}\right].$
To show (60), we observe that if $\phi_{i}^{*}\phi_{j}^{*}=1$ and $|\Gamma_{i}\cap\Gamma_{j}|=0$ then
$\mathbb{E}\left[v_{i}v_{j}|T_{i}T_{j}\right]$ $=\mathbb{E}\left[\frac{1\{\bar{Z}_{i}=Y_{ki}\}\cdot 1\{\bar{Z}_{j}=Y_{kj}\}\cdot N^{2}}{P(\bar{Z}_{i}=Y_{ki}|T_{i})P(\bar{Z}_{j}=Y_{kj}|T_{j})\hat{N}_{Y_{ki}}\hat{N}_{Y_{kj}}}|T_{i},T_{j}\right]$
$=P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})\cdot$
$\hskip 28.45274pt\mathbb{E}\left[\frac{N^{2}}{P(\bar{Z}_{i}=Y_{ki}|T_{i})P(\bar{Z}_{j}=Y_{kj}|T_{j})\hat{N}_{Y_{ki}}\hat{N}_{Y_{kj}}}|T_{i},T_{j},\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\right]$
$=\mathbb{E}\left[\frac{N^{2}}{\hat{N}_{Y_{ki}}\hat{N}_{Y_{kj}}}|T_{i},T_{j},\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\right]$
where we have used that the events $\{\bar{Z}_{i}=Y_{ki}\}$ and $\{\bar{Z}_{j}=Y_{kj}\}$ are independent conditioned on $T_{i}$ and $T_{j}$ when $|\Gamma_{i}\cap\Gamma_{i}|=0$.
To show (61), we observe that if $\phi_{i}^{*}\phi_{j}^{*}=1$ and $|\Gamma_{i}\cap\Gamma_{j}|\neq 0$, then
$\mathbb{E}\left[v_{i}v_{j}|T_{i}T_{j}\right]$ $=\mathbb{E}\left[\frac{1\{\bar{Z}_{i}=Y_{ki}\}\cdot 1\{\bar{Z}_{j}=Y_{kj}\}\cdot N^{2}}{P(\bar{Z}_{i}=Y_{ki}|T_{i})P(\bar{Z}_{j}=Y_{kj}|T_{j})\hat{N}_{Y_{ki}}\hat{N}_{Y_{kj}}}|T_{i},T_{j}\right]$
$=P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})\cdot$
$\hskip 28.45274pt\mathbb{E}\left[\frac{N^{2}}{P(\bar{Z}_{i}=Y_{ki}|T_{i})P(\bar{Z}_{j}=Y_{kj}|T_{j})\hat{N}_{Y_{ki}}\hat{N}_{Y_{kj}}}|T_{i},T_{j},\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\right]$
$\leq\frac{1}{p_{\min}^{2}}\mathbb{E}\left[\frac{N^{2}}{\hat{N}_{Y_{ki}}\hat{N}_{Y_{kj}}}|T_{i},T_{j},\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\right]$
Proof of Lemma 7.
Given a set $S\subseteq[N]$, let $\Gamma_{S}$ denote the set
$\Gamma_{S}=\bigcup_{i\in S}\Gamma_{i}$
and let $Q$ and $\bar{S}$ denote the sets
$Q$ =$=\{j:|\Gamma_{j}\cap\Gamma_{S}|=0\},$ $\bar{S}$ $=[N]/Q$
and given $Y_{ki},Q$, and $\bar{S}$, let $\hat{N}_{Q}$ and $\hat{N}_{\bar{S}}$ denote
$\hat{N}_{Q}$ $=\sum_{i\in Q}\frac{1\{\bar{Z}_{j}=Y_{ki}\}}{P\left(\bar{Z}_{j}=Y_{ki}|T_{j}\right)}$ $\hat{N}_{\bar{S}}$ $=\sum_{i\in\bar{S}}\frac{1\{\bar{Z}_{j}=Y_{ki}\}}{P\left(\bar{Z}_{j}=Y_{ki}|T_{j}\right)}$
so that $\hat{N}_{Y_{ki}}$ satisfies
$\hat{N}_{Y_{ki}}=\hat{N}_{Q}+\hat{N}_{\bar{S}}.$
We observe that $\hat{N}_{Q}$ is independent of $X_{\Gamma_{S}}$, and hence that
$\mathbb{E}\left[\hat{N}_{Q}\ \Big{|}\ X_{\Gamma_{S}}\right]=N_{Q}$
where $N_{Q}=N-|\bar{S}|$.
Under Assumption 1, it can be seen that changing $X$ in a single entry changes at $\hat{N}_{Q}$ by at most $C=d_{\max}/p_{\min}$. As a result, by McDiarmid’s inequality it holds that
$P\left(\left|\hat{N}_{Q}-N_{Q}\right|\geq\epsilon\ \Big{|}\ X_{\Gamma_{S}}\right)\leq\exp\left(-\frac{\epsilon^{2}}{NC^{2}}\right)$
Thus for $\epsilon_{B}=\sqrt{3C^{2}N\cdot\log N}$ and the indicator
$B=1\left\{\left|\hat{N}_{Q}-N_{Q}\right|\geq\epsilon_{B}\right\},$
it holds that
$P(B=1\ |\ X_{\Gamma_{S}})\leq 1/N^{3}.$ (113)
For $i,j:\phi_{j}^{*}\phi_{i}^{*}=1$, we let $F_{\{i\}}$ and $F_{\{i,j\}}$ denote sigma-fields such that
$\mathbb{E}[\ \cdot\ |F_{\{i\}}]$ $=\mathbb{E}[\ \cdot\ |T_{i},\bar{Z}_{i}=Y_{ki}]$
$\mathbb{E}[\ \cdot\ |F_{\{i,j\}}]$ $=\mathbb{E}[\ \cdot\ |T_{i},T_{j},\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}]$
as previously used in the proof of Lemma 2. Given $S$ equal to either $\{i\}$ or $\{i,j\}$, we can induce $F_{S}$, $Q$, and $\bar{S}$, and observe that
$\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{S}\right]$ $=\mathbb{E}\left[1+\frac{N-\hat{N}_{Y_{ki}}}{N}+\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S}\right]$ (114)
$\leq 1+\mathbb{E}\left[\frac{|\bar{S}|-\hat{N}_{\bar{S}}}{N}|F_{S}\right]+\mathbb{E}\left[\frac{N_{Q}-\hat{N}_{Q}}{N}|F_{S}\right]+\mathbb{E}\left[\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S}\right]$
$\leq 1+\frac{|\bar{S}|}{Np_{\min}}+0+\mathbb{E}\left[\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S},B=0\right]+P(B=1|F_{S})\cdot\mathbb{E}\left[\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S},B=1\right]$ (115)
$\leq 1+\frac{|\bar{S}|}{Np_{\min}}+\frac{(|\bar{S}|+\epsilon_{B})^{2}}{N(N_{Q}-\epsilon_{B})p_{\min}}+\frac{1}{N^{3}}\cdot\left(\frac{N}{p_{\min}}\right)^{2}\cdot\frac{1}{N}$ (116)
$=O\left(\frac{1}{N}\right)+O\left(\frac{\log N}{N}\right)+O\left(\frac{1}{N^{2}}\right)$ (117)
where (114) holds by the identity $\frac{a}{b}=1+\frac{a-b}{a}+\frac{(a-b)^{2}}{ab}$; (115) holds because conditioned on $F_{S}$, it holds that $1\leq\hat{N}_{\bar{S}}\leq\frac{|\bar{S}|}{p_{\min}}$; (116) holds by definition of the event $\{B=0\}$ and because conditioned on $F_{S}$, it holds that $1\leq\hat{N}_{Y_{ki}}\leq N/p_{\min}$; and (117) holds because $|\bar{S}|\leq 2d_{\max}^{2}$, $\epsilon_{B}=\sqrt{3C^{2}N\log N}$, $N_{Q}\geq N-2d_{\max}$, and (113). Letting $S=\{i\}$ implies (62), while letting $S=\{i,j\}$ implies (63).
To show (64), we let $S=\{i,j\}$ (inducing $F_{S}$, $Q$, and $\bar{S}$), and observe that
${\operatorname{Var}\,}\left(\frac{N}{\hat{N}_{Y_{ki}}}|F_{S}\right)$ $={\operatorname{Var}\,}\left[1+\frac{|\bar{S}|-\hat{N}_{\bar{S}}}{N}+\frac{N_{Q}-\hat{N}_{Q}}{N}+\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S}\right]$
$=5{\operatorname{Var}\,}\left[\frac{|\bar{S}|-\hat{N}_{\bar{S}}}{N}|F_{S}\right]+5{\operatorname{Var}\,}\left[\frac{N_{Q}-\hat{N}_{Q}}{N}|F_{S}\right]+5{\operatorname{Var}\,}\left[\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S}\right]$ (118)
$\leq\left(\frac{|\bar{S}|}{Np_{\min}}\right)^{2}+\frac{Nd_{\max}}{p_{min}^{2}N^{2}}+\mathbb{E}\left[{\operatorname{Var}\,}\left(\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S},B\right)|F_{S}\right]+$
$\hskip 28.45274pt{\operatorname{Var}\,}\left(\mathbb{E}\left[\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S},B\right]|F_{S}\right)$
$\leq\left(\frac{|\bar{S}|}{Np_{\min}}\right)^{2}+\frac{Nd_{\max}}{p_{min}^{2}N^{2}}+\left(\frac{(|\bar{S}|+\epsilon_{B})^{2}}{N(N_{Q}-\epsilon_{B})}\right)^{2}+\left(\frac{N^{2}}{p_{\min}^{2}N}\right)^{2}\cdot\frac{1}{N^{3}}$
$=O\left(\frac{1}{N^{2}}\right)+O\left(\frac{1}{N}\right)+O\left(\frac{\log N}{N}\right)+O\left(\frac{1}{N}\right)$
where (118) follows by the identity
${\operatorname{Var}\,}\sum_{i=1}^{3}a_{i}$ $=\sum_{i=1}^{3}\sum_{j=1}^{3}{\operatorname{Cov}}(a_{i},a_{j})$
$\leq\sum_{i=1}^{3}{\operatorname{Var}\,}a_{i}+\sum_{i=1}^{3}\sum_{j:j\neq i}({\operatorname{Var}\,}a_{i}+{\operatorname{Var}\,}a_{j})$
and the remaining steps follow by similar reasoning as (115)-(117). ∎
Proof of Lemma 8.
Let $h_{i}$ be defined as in Lemma 9, and let $\sigma^{2}={\operatorname{Var}\,}\sum_{i}h_{i}$. Divide the sequence of experiments into two subsequences, depending on whether $\sigma^{2}\geq N^{2/3+2\epsilon}$.
For the subsequence that satisfies $\sigma^{2}\geq N^{2/3+2\epsilon}$, it holds by linearization arguments made in the proof of Theorem 4 that
$\frac{\hat{\Phi}_{k}^{{\textup{Haj}}}-\Phi}{\sigma}$ $=\frac{\nabla h(\mathbb{E}\xi)^{T}(\xi-\mathbb{E}\xi)}{\sigma}+O_{P}\left(\frac{1}{\sigma}\right)$ (119)
$=\frac{\sum_{i=1}^{N}h_{i}}{\sigma}+O_{P}\left(\frac{1}{\sigma}\right)$ (120)
where (119) holds by (87) and (91), while (120) holds by (76)-(86) and (88). By Lemma 9 and (120) it follows that
$\frac{\hat{\Phi}_{k}^{{\textup{Haj}}}-\Phi}{\sigma}\rightarrow N(0,1)$
from which it follows by Lemma 10 that
$\liminf\limits_{N\rightarrow\infty}{\operatorname{Var}\,}\left[\frac{\hat{\Phi}_{k}^{\textup{Haj}}-\Phi}{\sigma}\right]\geq 1$ (121)
Since by Theorem 5 it holds that
$\frac{V_{k}}{\sigma^{2}}$ $=\frac{\mathbb{E}V_{k}}{\sigma^{2}}+O_{P}\left(\frac{N^{1/2+\delta}}{\sigma^{2}}\right)$
$\geq{\operatorname{Var}\,}\left(\frac{\hat{\Phi}_{k}^{\textup{Haj}}}{\sigma}\right)+O_{P}\left(\frac{N^{1/2+\delta}}{\sigma^{2}}\right)$ (122)
combining (121) and (122) yields
$\liminf\limits_{N\rightarrow\infty}\frac{V_{k}}{\sigma^{2}}\geq 1$
which implies that
$P\left(\Phi-\hat{\Phi}_{k}\geq z_{1-\frac{\alpha}{2}}\sqrt{V_{k}}\right)\leq\frac{\alpha}{2}+o(1)$ (123)
for the subsequence of experiments that satisfies $\sigma^{2}\geq N^{2/3+2\epsilon}$.
For the other subsequence, Chebychev inequality implies that
$P\left(\left|\sum_{i}h_{i}\right|\geq\frac{N^{1/3+\epsilon}}{\sqrt{\alpha/2}}\right)\leq\alpha/2$
where we have used that $\sigma^{2}<N^{2/3+2\epsilon}$. In turn, this implies for any $\epsilon^{\prime}>0$ that that
$P\left(\hat{\Phi}_{k}-\Phi\geq(1+\epsilon^{\prime})\frac{N^{1/3+\epsilon}}{\sqrt{\alpha/2}}\right)\leq 1-\alpha+o(1)$ (124)
where we have used that $\hat{\Phi}_{k}-\Phi=\sum_{i}h_{i}+O_{P}(1)$, which holds by (119). In turn, we may choose $\epsilon^{\prime}$ and $\epsilon$ such that (124) is equivalent to
$P\left(\hat{\Phi}_{k}-\Phi\geq z_{1-\frac{\alpha}{2}}\sqrt{\Delta}\right)\leq\alpha/2+o(1)$ (125)
where $\Delta=(1.01\ N^{2/3+0.01})/(z_{1-\frac{\alpha}{2}}^{2}\cdot\alpha/2)$. Combining (123) and (125) yields that for all experiments, it holds that
$P\left(\hat{\Phi}_{k}-\Phi\geq z_{1-\frac{\alpha}{2}}\sqrt{\max\left(V_{k},\Delta\right)}\right)\leq\alpha/2+o(1),$
As $\widetilde{V}_{k}$ can be seen to equal $\max\left(V_{k},\Delta\right)$, this proves the lemma.
Proof of Lemma 9.
It can be seen that $|h_{i}|\leq\frac{1}{p_{\min}}$, and that the dependency neighborhoods of $h_{1},\ldots,h_{N}$ have maximum size bounded by $d_{\max}^{2}$, since each $h_{i}$ depends on $X$ only through $X_{\Gamma_{i}}$. As a result, applying Theorem 7 yields
$d_{W}\left(\frac{\sum_{i}h_{i}}{\sigma}\ ,\ N(0,1)\right)$ $\leq\frac{d_{\max}^{6}N}{\sigma^{3}p_{\min}}+\frac{\sqrt{28}\,d_{\max}^{3}\,\sqrt{N}}{\sqrt{\pi}p_{\min}^{2}\sigma^{2}}$
$=o(1)$ (126)
where (126) uses $\sigma^{2}\geq N^{2/3+\delta}$ for any $\delta>0$. ∎