### v4

# Agnostic Characterization of Interference in Randomized Experiments

Abstract
We give an approach for characterizing interference by lower bounding the number of units whose outcome depends on selected groups of treated individuals, such as depending on the treatment of others, or others who are at least a certain distance away. The approach is applicable to randomized experiments with binary-valued outcomes. Asymptotically conservative point estimates and one-sided confidence intervals may be constructed with no assumptions beyond the known randomization design, allowing the approach to be used when interference is poorly understood, or when an observed network might only be a crude proxy for the underlying social mechanisms. Point estimates are equal to Hájek-weighted comparisons of units with differing levels of treatment exposure. Empirically, we find that the width of our interval estimates is competitive with (and often smaller than) those of the EATE, an assumption-lean treatment effect, suggesting that the proposed estimands may be intrinsically easier to estimate than treatment effects.
1 Introduction
There is growing interest in the analysis of randomized experiments where participants may affect each other (“interference between units”), for example by transmission of disease, sharing of information, peer influence, or economic competition. Such mechanisms often play fundamental roles in our understanding of social behaviors, and their empirical characterization may be of scientific interest.
We believe that a better understanding of interference may not only be of scientific interest, but also of immense practical importance. One of the most well-known findings in the history of public health–that cholera could be transmitted through shared water utilities–was a particularly impactful characterization of interference (Freedman, , 2010, Ch. 3). Conversely, List, (2022) and List, (2023) discuss fields experiments at Uber in which the results were misinterpreted because interference was ignored, leading to decisions that harmed the company.111(List, , 2022, pp. 93-94) reports that giving a rider a promotional coupon increased their Uber usage in a field study, but the possibility of negative effects on other riders (by increasing their waiting time for rides) was ignored. The promotion was subsequently expanded to cover an entire city, leading to long wait times followed by substantial reduction in Uber usage. Similarly, List, (2023) reports that allowing a driver to receive tips increased their productivity in a field experiment, but the possibility of negative effects on the productivity of other drivers (who competed for the same pool of riders) was again ignored. Tipping was subsequently allowed for all drivers, but no gains in driver productivity were realized. The 2007 subprime mortgage crisis and subsequent “Great Recession” have been attributed in part to questionable modeling of causal dependencies between mortgage defaults, which led to catastrophic underestimation of credit risk (Salmon, , 2012). [^2] Interference between mortgages arises if housing prices can affect–and are affected by–the rate of mortgage default.
In this work, we give a general approach for characterizing whether interference-causing mechanisms are widely prevalent in an experimental setting. The approach is applicable to experiments with binary-valued outcomes. It can be used to explore questions such as
  1. 1.
    How many units are affected by (i.e., their outcome depends on) the treatment of others? of distant others?
  2. 2.
    How many units, when receiving a certain level of direct treatment, are affected by the treatment of others? For example, if a vaccine grants immunity then the disease status of a unit receiving the placebo (but not necessarily when receiving the vaccine), might depend on the vaccination status of others.
For these types of questions, our approach gives conservative point estimates and one-sided confidence intervals, both of which lower bound the true value. Under reasonable experiment designs, our point estimates will converge to a lower bound on the true value, while the one-sided interval will cover with probability converging to at least the desired level. These consistency and coverage properties are agnostic, in that they require no assumptions or restrictions on the nature of the interference, and rely only on the known experiment design. As a result, our estimates will remain valid lower bounds (but may become less tight) if an observed social network is only a crude proxy for the actual social mechanisms.
We will demonstrate our approach using an experiment of Cai et al., (2015), which studied whether crop insurance decisions of farmers could be affected by information received by their peers. Such experiments, in which the primary goal is to demonstrate the presence of interference, may be considered part of the collection of “shoe leather” efforts that are necessary to deeply understand a problem setting (Freedman, , 2010).333“Historically, ‘shoe-leather epidemiology’ is epitomized by intensive, door-to-door canvassing that wears out investigators’ shoes” (Freedman, , 2010, p. xiii). Without such efforts, treatment effects and policy recommendations may lead to questionable conclusions, such as in the abovementioned Uber examples where interference was mischaracterized. On the other hand, a more informed understanding of interference may lead to better assumptions and more credible policy recommendations. For the experiment of Cai et al., (2015), our results strengthen their findings, with confidence intervals that are both tighter and require fewer assumptions than those of the expected average treatment effect (EATE), an assumption-lean treatment effect whose study was motivated by interference (Sävje et al., , 2021).
2 Related Work
Our work is closely related to hypothesis testing in the presence of interference, which has been studied in Aronow, (2012); Athey et al., (2018); Pouget-Abadie et al., (2019); Basse et al., (2019); Han et al., (2022); Puelz et al., (2022); Hoshino and Yanagi, (2023), for testing of null effects and hypotheses regarding interference structures, such as the null of no interference between units. In these settings, p-values generally cannot be inverted into confidence intervals for a treatment effect unless additional assumptions are added, and more generally the connection between testing and estimation under interference is not fully understood. Our work complements this literature, as we invert tests of interference specifications to construct one-sided intervals without additional assumptions – however not for a treatment effect, but rather for the number of units who violate the specification.
In doing so, we improve on prior work in this area (Choi, , 2023, 2024) in multiple respects. Most importantly, our confidence intervals are much tigher than those of previous approaches, due to a combination of improved computation, better estimator design, and a novel approach for variance estimation. Our target estimand is a proper parameter, whereas in prior approaches it was defined to be a random quantity whose value depended on the treatment assignment. Finally, our approach generalizes to answer a wide variety of questions regarding the characterization of interference.
The estimation of treatment effects in the presence of interference is an active area of research (Rosenbaum, , 2007; Sävje et al., , 2021; Li et al., , 2019; Hu et al., , 2021; Li and Wager, , 2022; Leung, , 2022; Ogburn et al., , 2022; Harshaw et al., , 2022; Zhao and Ding, , 2022; Park and Kang, , 2023; Gao and Ding, , 2023). In many of these works, the target parameter describes the effect of a marginal departure from the experiment design. For such estimands, consistent estimates can generally be constructed in assumption-lean regimes, although additional assumptions are needed for coverage of interval estimates. Harshaw et al., (2021) focuses on the problem of variance estimation. For settings with interference, Kandiros et al., (2024) proposes a novel, more-tailored approach to experiment design. Estimation of total effects (as opposed to marginal effects) assuming low order interactions between unit treatments is considered in Cortez et al., (2022) under staggered rollout, and in Cortez-Rodriguez et al., (2023) for known exposure neighborhoods. Farias et al., (2022), Dhaouadi et al., (2023), and Johari et al., (2024) consider a particular form of interference (“Markovian”) arising in e-commerce settings. Banerjee et al., (2024) gives an empirical example of information sharing in which interference may be more subtle and complex than is captured by models of local exposure such as infection.
3 Methodology (basic case)
In this section, we present a basic example, in which the estimand is the number of units who are affected by (i.e., their outcome depends on) the assignment of treatment. This quantity, which we denote by $\tau^{\textup{basic}}$, includes units who are affected by their own treatment as well as those who are affected by the treatment of others.
Under certain experiment designs, $\tau^{\textup{basic}}$ may be used to characterize interference. In bipartite network designs, some units are not directly treated, and $\tau^{\textup{basic}}$ may be used to characterize whether these units are affected by the treatment of others (Zigler and Papadogeorgou, , 2020). In experiments involving the assignment of units to groups, or the formation of network ties, $\tau^{\textup{basic}}$ characterizes whether units are affected by their randomly assigned neighbors or group members (Basse et al., , 2024; Li et al., , 2019).
The remaining organization of the paper is as follows. Sections 3.1 through 3.7 respectively present the generative model, the estimand $\tau^{\textup{basic}}$, point estimation, usage of network information, Hájek normalization, inference, and computation. In Section 4, the approach is then generalized to cover a wider range of target parameters, such as the questions presented in the introduction. Section 5 considers asymptotic behavior, and Section 6 gives a simulation study and data example to illustrate usage. All propositions and theorems are proven in the supplemental materials.
3.1 Preliminaries and Oveview
We consider an experiment on $N$ units, with $X=(X_{1},\ldots,X_{N})$ denoting the binary-valued treatment of each unit, and $Y=(Y_{1},\ldots,Y_{N})$ denoting their binary outcomes. We allow for arbitrary interference, so that the outcome $Y_{i}$ of unit $i$ may potentially depend on all $N$ treatments,
$Y_{i}=f_{i}(X_{1},\ldots,X_{N}),\qquad i\in[N]$ (1)
where the potential outcome mappings $\{f_{i}\}_{i=1}^{N}$ may be arbitrary and unknown.
Given an observed network, we will use ${\mathcal{N}}_{i}$ to denote the set consisting of unit $i$ and $i$’s “close” associates, and ${\mathcal{N}}_{i}^{(2)}$ to denote “non-close” associates of unit $i$, as defined by the researcher. For example, ${\mathcal{N}}_{i}$ might denote unit $i$ and its direct neighbors on the network, while ${\mathcal{N}}_{i}^{(2)}$ might denote units who are not in $\mathcal{N}_{i}$ but have overlapping sets of neighbors with unit $i$. We will let $W_{i}$ denote the thresholded number of treated close associates for unit $i$,
[…]
and $W_{i}^{(2)}$ will denote the thresholded number of treated units in $\mathcal{N}_{i}^{(2)}$,
[…]
for some choice of thresholds $t_{i}$ and $t_{i}^{(2)}$.
Given integer $n>0$, we use $[n]$ to denote the set $\{1,\ldots,n\}$. Given a set $S\subseteq[N]$, we let $|S|$ denote its cardinality, and let $X_{S}$ denote the subvector of $X$ corresponding to the members of $S$.
Overview of Approach
To estimate lower bounds on our target estimand $\tau^{\textup{basic}}$, we will take the following approach:
  1. 1.
    Define the estimand $\tau^{\textup{basic}}$ in terms of an particular subset ${\mathcal{I}}\subseteq[N]$, whose identity is unknown
  2. 2.
    Propose idealized estimators $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ which will have good statistical properties, such as consistency and asymptotic normality, but require knowledge of $\mathcal{I}$}
  3. 3.
    Show that $\Delta$, the difference in average outcomes between treated and control (which can be computed without knowledge of $\mathcal{I}$}) converges to a lower bound for $\max(\hat{\tau}_{1},\hat{\tau}_{2})$, so that if $\tau_{1}$ and $\tau_{2}$ are both consistent for $\tau^{\textup{basic}}$, then $\Delta$ is an asymptotic lower bound.
  4. 4.
    Lower bound the confidence intervals induced by $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ and their variance estimates, by solving an integer program over the unknown subset $\mathcal{I}$}.
3.2 Estimand
Let $\tau^{\textup{basic}}$ denote the number of units who are affected by treatment, including both their own treatment or the treatment of others. We may write $\tau^{\textup{basic}}$ as the number of units whose outcome $f_{i}(X)$ is not constant over the possible treatment assignments $X\in\{0,1\}^{N}$.
$\tau^{\textup{basic}}=\sum_{i=1}^{N}1\{f_{i}(X)\textup{ is not constant in }X\}$
Define ${\mathcal{I}}\subset[N]$ as the subset of units who are unaffected by treatment and have constant outcome mappings,
${\mathcal{I}}=\{i:f_{i}(X)\textup{ is constant in }X\}$
so that $\tau^{\textup{basic}}=N-|{\mathcal{I}}|$.
3.3 Point estimation
Let $H_{i}$ denote the indicator of whether unit $i$’s treatment and outcome have the same binary value,
$H_{i}=1\{(X_{i},Y_{i})=(1,1)\textup{ or }(0,0)\},$ (4)
and let $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ denote sampling-based estimators of $\tau^{\textup{basic}}=N-|{\mathcal{I}}|$, in which the unknown cardinality of $\mathcal{I}$} is unbiasedly estimated by a probability-weighted (i.e., Horvitz-Thompson) sample:
$\hat{\tau}_{1}=N-\sum_{i\in{\mathcal{I}}}\frac{1\{H_{i}=1\}}{P(H_{i}=1)}\qquad\textup{ and }\qquad\hat{\tau}_{2}=N-\sum_{i\in{\mathcal{I}}}\frac{1\{H_{i}=0\}}{P(H_{i}=0)}$ (5)
Because $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ involve only units in $\mathcal{I}$} whose outcomes are unaffected by treatment, they often will exhibit simple statistical behavior, even if strong interference exists between units who are not in $\mathcal{I}$}. For example, if treatment is assigned by independent Bernoulli randomization, then $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ are sums of independent variables. [^4] If $i\in{\mathcal{I}}$, then $Y_{i}$ equals a constant $c_{i}$ for all $X\in\{0,1\}$. Thus it holds for example that $P(H_{i}=1,H_{j}=1)=P(X_{i}=c_{i},X_{j}=c_{j})=P(X_{i}=c_{i})P(X_{j}=c_{j})=P(H_{i}=1)P(H_{j}=1)$, if $i,j\in{\mathcal{I}}$ and $X_{i}\perp\!\!\!\!\perp X_{j}$. Similarly, if the dependencies between the unit treatments are bounded, then $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ are sums of variables whose dependencies will be similarly bounded. For this reason, under a variety of designs we may expect the values of $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$, while unknown due to $\mathcal{I}$} being unknown, to concentrate at their expectation (which equals $\tau^{\textup{basic}}$) and to be asymptotically normal.
Our motivation for $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ is the following: under mild conditions on the experiment design, the maximum of $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ is asymptotically lower bounded by the magnitude of the propensity-weighted difference in outcomes between treated and control, given by
$\Delta=\sum_{i=1}^{N}\left(\frac{X_{i}}{P(X_{i}=1)}-\frac{1-X_{i}}{P(X_{i}=0)}\right)Y_{i},$
as stated by Proposition 1 below:
Proposition 1.
Let the total weights of the treated and control converge to their expectations, so that
$\sum_{i=1}^{N}\frac{X_{i}}{P(X_{i}=1)}=N+O_{P}(N^{1/2})\quad\textup{and}\quad\sum_{i=1}^{N}\frac{1-X_{i}}{P(X_{i}=0)}=N+O_{P}(N^{1/2})$ (6)
Then it holds that
$\left|\sum_{i=1}^{N}\left(\frac{X_{i}}{P(X_{i}=1)}-\frac{1-X_{i}}{P(X_{i}=0)}\right)Y_{i}\right|\leq\max(\hat{\tau}_{1},\hat{\tau}_{2})+O_{P}(N^{1/2})$ (7)
As a corollary, it immediately follows that if $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ are consistent for $\tau^{\textup{basic}}$, then $\max(\hat{\tau}_{1},\hat{\tau}_{2})$ will be consistent as well, and $|\Delta|$ will be an asymptotic lower bound for $\tau^{\textup{basic}}$.
Corollary 1.
Let the conditions of Theorem 1 hold. If $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ are within $O_{P}(N^{1/2})$ of $\tau^{\textup{basic}}$, then
$\left|\sum_{i=1}^{N}\left(\frac{X_{i}}{P(X_{i}=1)}-\frac{1-X_{i}}{P(X_{i}=0)}\right)Y_{i}\right|\leq\tau^{\textup{basic}}+O_{P}(N^{1/2})$ (8)
We note that $\Delta$, which involves units outside of $\mathcal{I}$}, need not concentrate in order for (7) or (8) to hold. In contrast to $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$, which depend only on units in $\mathcal{I}$}, it may well be that $\Delta$ becomes less stable as the number of participants grows, but nonetheless is asymptotically a valid lower bound on the estimand.
Discussion
To give intuition for Theorem 1, consider an experiment in which $100$ units are randomly divided into equal-sized groups for treatment and control, with average outcomes of $0.4$ (20/50) for the treatment group and $0.8$ (40/50) for the control. How large can $|{\mathcal{I}}|$ be? Units in $\mathcal{I}$} have constant outcomes and are equally likely to be assigned to treatment or control. It follows that up to sampling variation, such units should be evenly distributed between treatment and control for each outcome level. Since there are 20 treated units and 40 control units with outcome level 1, it trivially holds that at most all 20 of the treated units with outcome level 1 could belong to $\mathcal{I}$}, and thus at most roughly 20 out of the 40 control units with outcome level 1 could belong to $\mathcal{I}$} as well. (Otherwise, units in $\mathcal{I}$} with outcome level 1 would not be evenly distributed between treatment and control.) Likewise, since only 10 control units have outcome level 0, at most all 10 could belong to $\mathcal{I}$}, implying that at most roughly 10 out of the 30 treated units with outcome level 0 could belong to $\mathcal{I}$} as well. This yields an upper bound of $(20+10)\times 2=60$ units who are unaffected by treatment, or equivalently at least 40 units who are affected by treatment. This lower bound of 40 equals the difference in outcomes of 0.4 between treatment and control, and also equals the maximum possible value of $\hat{\tau}_{1}$ over all hypotheses for the unknown $\mathcal{I}$}. The limiting quantities that were used to construct this bound, namely the 20 treated units with outcome equal to 1, and the 10 control units with outcome equal to 0, correspond to the units $\{i:H_{i}=1\}$.
When might this lower bound on $\tau$ be tight? Can $\tau$ be upper bounded? To build intuition, we continue the example by presenting a generative model under which the lower bound is tight. In this model, the contrast variable $X$ appearing in (7) is a perfect predictor of the outcome for units not in $\mathcal{I}$. We then describe two models under which the lower bound is loose, and which suggest that non-vacuous upper bounds on $\tau$ cannot be constructed without additional assumptions.
Example 1.
(SUTVA with nonnegative effects) Of the 100 units in the experiment, suppose that units 1-60 are unaffected by treatment, and units 61-100 are affected by their own treatment, according to
$f_{i}(X)=\begin{cases}1&1\leq i\leq 40\\ 0&41\leq i\leq 60\\ X_{i}&61\leq i\leq 100\end{cases}$
Under this model, $40$ units are affected by treatment, and the expected outcomes for treated and control can be seen to be $0.8$ and $0.4$. Hence the diff-in-means will equal the fraction of units affected by treatment (up to sampling variation).
Example 2.
(SUTVA with cancellation of effects) Suppose that units 1-20 are unaffected by treatment, while units 21-40 are affected negatively by their own treatment, and units 41-100 are affected positively,
$f_{i}(X)=\begin{cases}1&1\leq i\leq 20\\ 1-X_{i}&21\leq i\leq 40\\ X_{i}&41\leq i\leq 100\end{cases}$
While the expected outcomes match the previous example, equaling $0.8$ and $0.4$ for treated and control, the diff-in-means is a loose lower bound on the fraction of units affected by treatment, which is 0.8.
Example 3.
(Interference) Let the outcome models be given by
[…]
where $(\alpha_{i1},\alpha_{i0})$ match the potential outcomes of the first example, equaling $(1,1)$ for units 1-40, (0,0) for units 41-60, and $(1,0)$ for units 61-100, while $(\beta_{i1},\beta_{i0})$ takes values $(1,0)$ and $(0,1)$ in some arbitrary proportion. If half of the units are randomly assigned to treatment, then $\sum_{i}X_{i}$ is even and the outcomes will be indistinguishable from those of the first example; however, unlike the first example, all 100 units are affected by treatment. In this example $\Delta$ does not concentrate at its expectation, but is nonetheless a valid lower bound for $\tau^{\textup{basic}}$.
3.4 Using Network Information
Given an elicited social network, we may consider contrasts other than $\Delta$, based on a new variable $Z=(Z_{1},\ldots,Z_{N})$ which encodes the direct and indirect treatment of each unit. We will consider $Z\in\{1,0,-1\}^{N}$, where $1$ and 0 respectively denote generalized versions of treatment and control, and 1-1 denotes that the unit will be excluded from comparison.
For example, to use $W$, the thresholded number of treated neighbors given by (2), we can define $Z$ according to
$Z_{i}=\begin{cases}1&\textup{ if $X_{i}=W_{i}=1$}\\ 0&\textup{ if $X_{i}=W_{i}=0$}\\ -1&\textup{ if $X_{i}\neq W_{i}$}\end{cases}$ (9)
Under $Z$ so defined, unit $i$ receives “effective treatment” if $X_{i}=1$ and $W_{i}=1$; receives “effective control” if $X_{i}=0$ and $W_{i}=0$; and otherwise is excluded from comparison.
Given $Z$, we can redefine $H=(H_{1},\ldots,H_{N})$ to use $Z$ instead of $X$,
[…]
which induces new versions of $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ following (5) as before. Proposition 1 can then be adapted to show that the difference in outcomes between effective treatment and effective control, given by
$\Delta^{Z}=\sum_{i=1}^{N}\left(\frac{1\{Z_{i}=1\}}{P(Z_{i}=1)}-\frac{1\{Z_{i}=0\}}{P(Z_{i}=0)}\right)Y_{i}$
is an asymptotic lower bound on the maximum of $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$.
Proposition 2.
Let $Z$ be given by (9) and let the total weights of effective treatment and control converge to their expectations,
$\sum_{i=1}^{N}\frac{1\{Z_{i}=1\}}{P(Z_{i}=1)}=N+O_{P}(N^{1/2})\quad\textup{ and }\quad\sum_{i=1}^{N}\frac{1\{Z_{i}=0\}}{P(Z_{i}=0)}=N+O_{P}(N^{1/2})$
and let $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ be given by (5) for $H$ given by (10). Then it holds that
$\left|\sum_{i=1}^{N}\left(\frac{1\{Z_{i}=1\}}{P(Z_{i}=1)}-\frac{1\{Z_{i}=0\}}{P(Z_{i}=0)}\right)Y_{i}\right|\leq\max(\hat{\tau}_{1},\hat{\tau}_{2})+O_{P}(N^{1/2})$ (11)
If $Z$ is a better predictor of the unit outcomes than $X$, then $\Delta^{Z}$ may be larger than the comparison of direct treatment and direct control given by $\Delta$. This results in a larger and tighter lower bound on $\tau^{\textup{basic}}$, assuming consistency of $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$.
3.5 Hájek Normalization
To reduce the variance of our IPW-based estimators, we may use Hájek normalization (Aronow and Samii, , 2017; Gao and Ding, , 2023). As a preliminary step, let $Y_{ki}$ denote transformed outcomes given by
$Y_{ki}=\begin{cases}Y_{i}&\textup{ if }k=1\\ 1-Y_{i}&\textup{ if }k=2,\end{cases}\qquad i\in[N],\ k=1,2,$ (12)
This allows $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ using $H$ given by (10) to be written with a single formula
$\hat{\tau}_{k}=N-\sum_{i\in{\mathcal{I}}}\frac{1\{Z_{i}=Y_{ki}\}}{P(Z_{i}=Y_{ki})},\qquad k=1,2,$ (13)
which will simplify our discussion of Hájek normalization for $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$.
To construct a Hájek-normalized version of $\hat{\tau}_{k}$ as given by (13), we first observe that it equals
$\hat{\tau}_{k}=N-\sum_{i\in{\mathcal{I}}}\left(\frac{1\{Z_{i}=1,Y_{ki}=1\}}{P(Z_{i}=1)}+\frac{1\{Z_{i}=0,Y_{ki}=0\}}{P(Z_{i}=0)}\right)$ (14)
where we have divided the event $\{Z_{i}=Y_{ki}\}$ into sub-events $\{Z_{i}=Y_{ki}=1\}$ and $\{Z_{i}=Y_{ki}=0\}$, and replaced the propensity score $P(Z_{i}=Y_{ki})$ by
$P(Z_{i}=Y_{ki})=Y_{ki}P(Z_{i}=1)+(1-Y_{ki})P(Z_{i}=0),\qquad\forall\ i\in{\mathcal{I}},$
which can be done because $Y_{ki}$ is constant in $X$ for $i\in{\mathcal{I}}$. Eq. (14) suggests a Hájek normalized version given by
$\hat{\tau}_{k}^{{\textup{Haj}}}=N-\sum_{i\in{\mathcal{I}}}\left[\frac{N}{\hat{N}_{1}}\frac{1\{Z_{i}=1\}}{P(Z_{i}=1)}\cdot Y_{ki}+\frac{N}{\hat{N}_{0}}\frac{1\{Z_{i}=0\}}{P(Z_{i}=0)}\cdot(1-Y_{ki})\right],$ (15)
where $\hat{N}_{0}$ and $\hat{N}_{1}$ are estimates of $N$ given by
$\hat{N}_{0}$ $=\sum_{j=1}^{N}\frac{1\{Z_{j}=0\}}{P(Z_{j}=0)},$ $\hat{N}_{1}$ $=\sum_{j=1}^{N}\frac{1\{Z_{j}=1\}}{P(Z_{j}=1)}.$
In Theorem 1, we show that $\max(\hat{\tau}_{1}^{\textup{Haj}},\hat{\tau}_{2}^{\textup{Haj}})$ is lower bounded by a Hájek-weighted comparison of units,
$\left|\sum_{i=1}^{N}\left(\frac{N}{\hat{N}_{1}}\frac{Z_{i}}{P(Z_{i}=1)}-\frac{N}{\hat{N}_{0}}\frac{(1-Z_{i})}{P(Z_{i}=0)}\right)Y_{i}\right|\leq\max(\hat{\tau}_{1}^{\textup{Haj}}\,,\,\hat{\tau}_{2}^{\textup{Haj}})$
noting that the inequality holds non-asymptotically, in contrast to (7) and (11).
3.6 Inference and Variance Estimation
If $\hat{\tau}_{1}^{\textup{Haj}}$ and $\hat{\tau}_{2}^{\textup{Haj}}$ are asymptotically normal, with consistent variance estimators denoted by $\hat{V}_{1}$ and $\hat{V}_{2}$, then by combining 1-sided confidence intervals it holds with probability converging to at least $1-\alpha$ that
$\tau^{\textup{basic}}$ $\geq\max\left\{\hat{\tau}_{1}^{\textup{Haj}}-z_{1-\frac{\alpha}{2}}\sqrt{\hat{V}_{1}},\,\hat{\tau}_{2}^{\textup{Haj}}-z_{1-\frac{\alpha}{2}}\sqrt{\hat{V}_{2}}\right\}.$ (16)
As the right hand side of (16) requires knowledge of $\mathcal{I}$}, it cannot be computed.
To construct a computable one-sided confidence interval for $\tau^{\textup{basic}}$, we will lower bound (16) by minimizing over all possible hypotheses for the unknown $\mathcal{I}$}. Doing so results in the confidence statement that with probability at least $1-\alpha$,
$\tau^{\textup{basic}}\geq\max\left\{\min_{\phi\in\{0,1\}^{N}}\hat{\tau}_{1}^{\textup{Haj}}(\phi)-z_{1-\frac{\alpha}{2}}\sqrt{\hat{V}_{1}(\phi)},\,\min_{\phi\in\{0,1\}^{N}}\hat{\tau}_{2}^{\textup{Haj}}(\phi)-z_{1-\frac{\alpha}{2}}\sqrt{\hat{V}_{2}(\phi)}\,\right\},$ (17)
where $\hat{\tau}_{k}^{\textup{Haj}}(\phi)$ and $\hat{V}_{k}(\phi)$ denote $\hat{\tau}_{k}^{\textup{Haj}}$ and $\hat{V}_{k}$ evaluated under the hypothesis that ${\mathcal{I}}=\{i:\phi_{i}=1\}$ for $\phi\in\{0,1\}^{N}$.
To complete our discussion of interval estimation, we next give an expression for $\hat{V}_{k}$, and then discuss computation.
Variance Estimation
The expression for $\hat{\tau}_{k}^{\textup{Haj}}$ given by (15) is equivalent to
$\hat{\tau}_{k}^{{\textup{Haj}}}=N-\sum_{i\in{\mathcal{I}}}\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{Z_{i}=Y_{ki}\}}{P(Z_{i}=Y_{ki})}$
It follows that the variance of $\hat{\tau}_{k}^{\textup{Haj}}$ can be written as a sum of covariance terms,
${\operatorname{Var}\,}\hat{\tau}_{k}^{\textup{Haj}}=\sum_{i\in{\mathcal{I}}}\sum_{j\in{\mathcal{I}}}{\operatorname{Cov}}\left(\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{Z_{i}=Y_{ki}\}}{P(Z_{i}=Y_{ki})}\ ,\ \frac{N}{\hat{N}_{Y_{kj}}}\cdot\frac{1\{Z_{j}=Y_{kj}\}}{P(Z_{j}=Y_{kj})}\right)$ (18)
To construct the one-sided interval (17), we propose $\hat{V}_{k}$ given by
$\hat{V}_{k}=\sum_{i\in{\mathcal{I}}}\sum_{j\in{\mathcal{I}}}{\operatorname{Cov}}\left(\frac{N}{\hat{N}_{Y_{ki}}}\frac{1\{Z_{i}=Y_{ki}\}}{P(Z_{i}=Y_{ki})}\ ,\ \frac{N}{\hat{N}_{Y_{kj}}}\frac{1\{Z_{j}=Y_{kj}\}}{P(Z_{j}=Y_{kj})}\right)\cdot\frac{1\{Z_{i}=Y_{ki},Z_{j}=Y_{kj}\}}{P(Z_{i}=Y_{ki},Z_{j}=Y_{kj})}$ (19)
Comparing (18) and (19), it can be seen that $\hat{V}_{k}$ is a Horvitz-Thompson estimator which samples the covariance terms appearing in $\operatorname{Var}\hat{\tau}_{k}^{\textup{Haj}}$. Hence it is unbiased provided that the sampling probabilities $P(Z_{i}=Y_{ki},Z_{j}=Y_{kj})$ are non-zero for all $i,j\in{\mathcal{I}}$. Should this condition not hold due to dependencies between the entries of $Z$, our variance estimators will be conservatively biased, as will be established in Theorem 2.
We use the estimated variance $\hat{V}_{k}$, rather than the actual variance given by (18), in the 1-sided interval given by (17). This is because (17) involves a minimization over all possible hypotheses for the unknown set $\mathcal{I}$}, and $\hat{V}_{k}$ can be less sensitive than the actual variance to this optimization, resulting in a less conservative interval. This is particularly the case when the contrast $\Delta^{Z}$ is large, as then either $\hat{V}_{1}$ or $\hat{V}_{2}$ involves only a small number of sampled covariance terms.
To evaluate $\hat{V}_{k}$ given by (19), we may use the following expressions for the covariance and propensity values which hold for all $i,j\in{\mathcal{I}}$,
${\operatorname{Cov}}\left(\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{Z_{i}=Y_{ki}\}}{P(Z_{i}=Y_{ki})}\ ,\ \frac{N}{\hat{N}_{Y_{kj}}}\cdot\frac{1\{Z_{j}=Y_{kj}\}}{P(Z_{j}=Y_{kj})}\right)=\\ \sum_{a=0}^{1}\sum_{b=0}^{1}1\{Y_{ki}=a,Y_{kj}=b\}\ {\operatorname{Cov}}\left(\frac{N}{\hat{N}_{a}}\cdot\frac{1\{Z_{i}=a\}}{P(Z_{i}=a)}\ ,\ \frac{N}{\hat{N}_{b}}\cdot\frac{1\{Z_{j}=b\}}{P(Z_{j}=b)}\right)$ (20)
and
$P(Z_{i}=Y_{ki},Z_{j}=Y_{kj})=\sum_{a=0}^{1}\sum_{b=0}^{1}1\{Y_{ki}=a,Y_{kj}=b\}\ P(Z_{i}=a,Z_{j}=b).$ (21)
The expressions (20) and (21) hold because $Y_{ki}$ and $Y_{kj}$ are constants when $i,j\in{\mathcal{I}}$.
3.7 Computation
To construct the one-sided interval (17), we must solve the optimization problem
$\min_{\phi\in\{0,1\}^{N}}\hat{\tau}_{k}^{\textup{Haj}}(\phi)-z_{1-\frac{\alpha}{2}}\sqrt{\hat{V}_{k}(\phi)},$ (22)
for $k=1,2$. We can write this problem as
$\min_{\phi\in\{0,1\}^{N}}N-b^{T}\phi-z_{1-\frac{\alpha}{2}}\sqrt{\phi^{T}Q\phi}$ (23)
where $\hat{\tau}_{k}^{\textup{Haj}}(\phi)=N-b^{T}\phi$ and $\hat{V}_{k}(\phi)=\phi^{T}Q\phi$ for $b\in\mathbb{R}^{N}$ and $Q\in\mathbb{R}^{N\times N}$ given by
$b_{i}$ $=\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{Z_{i}=Y_{ki}\}}{Y_{ki}\cdot P(Z_{i}=1)+(1-Y_{ki})\cdot P(Z_{i}=0)}$ $i\in[N]$ (24)
$Q_{ij}$ $=C_{ij}\cdot\frac{1\{Z_{i}=Y_{ki},Z_{j}=Y_{kj}\}}{P_{ij}}$ $i,j\in[N]^{2}.$ (25)
in which $C_{ij}$ and $P_{ij}$ appearing in (25) are given by the right hand sides of (20) and (21) respectively
To solve (23), we let $q_{i}$ denote the $i$th row of $Q$, let $B=\sum_{ij}|Q_{ij}|$, and reformulate the problem as
$\max_{\phi,u}$ $\ b^{T}\phi+z_{1-\frac{\alpha}{2}}\sqrt{\sum_{i=1}^{N}u_{i}}$
subject to $\ u_{i}\leq q_{i}^{T}\phi+B(1-\phi_{i})$ $\forall\ i\in[N]$
$\ u_{i}\geq q_{i}^{T}\phi-B(1-\phi_{i})$ $\forall\ i\in[N]$
$\ u_{i}\leq B\phi_{i}$ $\forall\ i\in[N]$
$\ u_{i}\geq-B\phi_{i}$ $\forall\ i\in[N]$
$\ \phi\in\{0,1\}^{N}$
$\ u\in\mathbb{R}^{N}$
It can be seen that $u_{i}=0$ if $\phi_{i}=0$, and $u_{i}=q_{i}^{T}\phi$ if $\phi_{i}=1$, enforcing that $\sum_{i}u_{i}=\phi^{T}Q\phi$. We solve this binary program using Gurobi, a commercial optimization solver. For the instances that we considered, involving 1000 or more units, Gurobi was generally able to solve the problem to acceptable tolerances. If stopped early, Gurobi returns an upper bound, retaining the coverage properties of the actual solution.
4 Generalization to other estimands
We now generalize the approach of Section 3 to cover estimands that more directly characterize interference, such as those described in the introduction. In Section 4.1 we introduce the estimands $\tau^{\textup{indirect}}$, $\tau^{\textup{nonneighbors}}$, $\tau^{\textup{control}}$, and $\tau^{\textup{tr}}$, defining $\mathcal{I}$} and Horvitz-Thompson estimators for each one. In Section 4.2 we then give a general formulation that covers each of these examples, and define the Hájek-normalized estimators $\hat{\tau}_{1}^{\textup{Haj}}$ and $\hat{\tau}_{2}^{\textup{Haj}}$ and the Hájek-normalized contrast $\Delta^{\textup{Haj}}$. Section 4.3 then presents variance estimators and inference.
4.1 Other Estimands
The number of units affected by the treatment of others ($\tau^{\textup{indirect}}$)
Let ${\mathcal{I}}\subseteq[N]$ identify the units who are not affected by the treatment of others
[…]
and let $\tau^{\textup{indirect}}=N-|{\mathcal{I}}|$, the number of units affected by the treatment of others.
To estimate $\tau^{\textup{indirect}}$, we may use estimators $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ given by
$\hat{\tau}_{k}$ $=N-\sum_{i\in{\mathcal{I}}}\frac{1\{W_{i}=Y_{ki}\}}{{P}(W_{i}=Y_{ki}|X_{i})}.$
This estimator differs from (13) in that we condition on $X_{i}$ in the probability weight $P(W_{i}=Y_{ki}|X_{i})$. This is done because the outcome $Y_{ki}$ is no longer constant for $i\in{\mathcal{I}}$, but instead is a function of $X_{i}$. As a result, the probability $P(W_{i}=Y_{ki})$ is unknown if $f_{i}$ is unknown, even if $i\in{\mathcal{I}}$. On the other hand, the conditional probability $P(W_{i}=Y_{ki}|X_{i})$ may be evaluated without knowledge of $f_{i}$, according to
$P(W_{i}=Y_{ki}|X_{i})=Y_{ki}\cdot P(W_{i}=1|X_{i})+(1-Y_{ki})\cdot P(W_{i}=0|X_{i}),\qquad i\in{\mathcal{I}}$
as will be shown by Proposition 4 in Section 4.2.
Long-range interference ($\tau^{\textup{nonneighbors}}$)
Let ${\mathcal{I}}\subseteq[N]$ identify the units who are affected only by their own treatment and that of their close associates,
[…]
and let $\tau^{\textup{nonneighbors}}$ equal $N-|{\mathcal{I}}|$, the number of units affected by the treatment of others excluding close associates. To estimate $\tau^{\textup{nonneighbors}}$, we may use $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ given by
$\hat{\tau}_{k}$ $=N-\sum_{i\in{\mathcal{I}}}\frac{1\{W_{i}^{(2)}=Y_{ki}\}}{{P}(W_{i}^{(2)}=Y_{ki}|X_{{\mathcal{N}}_{i}})}$
where $W_{i}^{(2)}$ is given by (3), and measures treatment of non-close associates.
The number of units whose outcome under direct control (treatment) is affected by the treatment of others ($\tau^{\textup{control}}$, $\tau^{\textup{tr}}$)
Let ${\mathcal{I}}\subseteq[N]$ identify the units whose outcome, when receiving control themselves, does not depend on the treatment of others
[…]
in which case $\tau^{\textup{control}}=N-|{\mathcal{I}}|$ equals the number of units whose outcome under direct control is affected by the treatment of others. For example, if a vaccine grants immunity, then the disease status of a unit when receiving the placebo (but not necessarily when receiving the vaccine) might depend on the vaccination status of others.
To estimate $\tau^{\textup{control}}$, we consider $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ given by
$\hat{\tau}_{k}$ $=N-\sum_{i\in{\mathcal{I}}}\frac{1\{X_{i}=0,W_{i}=Y_{ki}\}}{{P}(X_{i}=0,W_{i}=Y_{ki})}$
We may similarly define $\tau^{\textup{tr}}$, the number of units whose outcome under direct treatment is affected by the treatment of others, by defining $\mathcal{I}$} to be
[…]
and $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ to be
$\hat{\tau}_{k}$ $=N-\sum_{i\in{\mathcal{I}}}\frac{1\{X_{i}=1,W_{i}=Y_{ki}\}}{{P}(X_{i}=1,W_{i}=Y_{ki})}$
4.2 General Functional Form
We give a general functional form for our estimands and estimators. We may write our estimands as
$\tau=N-|{\mathcal{I}}|,$ (27)
where ${\mathcal{I}}\subseteq[N]$ contains the units whose outcome mapping satisfies an interference specification, which may be written as
${\mathcal{I}}=\{i:\textup{ $\exists\ \tilde{f}_{i}$ such that $f_{i}(X)=\tilde{f}_{i}(T_{i}(X))$ for all $X\in R_{i}$}\}$ (28)
where $T_{i}\equiv T_{i}(X)$ equals a vector of summary statistics describing unit $i$’s exposure to treatment, and $R_{i}$ is a subset of $\{0,1\}^{N}$. For units in $\mathcal{I}$}, the value of $T_{i}$ dictates unit $i$’s outcome whenever $X\in R_{i}$. Both $T_{i}$ and $R_{i}$ are chosen by the researcher to define the specification of interest. For example, letting $T_{i}=X_{i}$ and $R_{i}=\{0,1\}^{N}$ induces $\mathcal{I}$} given by (26).
In addition to $T_{i}$ and $R_{i}$, the researcher also chooses $Z_{i}\equiv Z_{i}(X)\in\{1,0,-1\}$, which maps unit $i$’s treatment exposure to one of 3 groups (i.e., effect treatment, effective control, or not used). Table 1 gives values for $\{T_{i},R_{i},Z_{i}\}_{i=1}^{N}$ corresponding to the estimands considered in this paper.
To simplify notation, let $\bar{Z}_{i}$ for $i\in[N]$ equal $Z_{i}$ if $X\in R_{i}$, and equal -1 otherwise,
$\bar{Z}_{i}=\begin{cases}Z_{i}&\textup{ if $X\in R_{i}$}\\ -1&\textup{ if $X\notin R_{i}$,}\end{cases}$ (29)
allowing the event $\{X\in R_{i},Z_{i}=Y_{ki}\}$ to be written more concisely as $\{\bar{Z}_{i}=Y_{ki}\}$ under binary outcomes.
The Horvitz-Thompson estimators $\hat{\tau}_{k}$ for $k=1,2$ are given by
$\hat{\tau}_{k}=N-\sum_{i\in{\mathcal{I}}}\frac{1\{\bar{Z}_{i}=Y_{ki}\}}{P(\bar{Z}_{i}=Y_{ki}|T_{i})}$ (30)
and Hájek-normalized versions of $\hat{\tau}_{k}$ are given by
$\hat{\tau}_{k}^{\textup{Haj}}=N-\sum_{i\in{\mathcal{I}}}\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{\bar{Z}_{i}=Y_{ki}\}}{P(\bar{Z}_{i}=Y_{ki}|T_{i})},$ (31)
where $\hat{N}_{1}$ and $\hat{N}_{0}$ denote Hájek normalization factors
$\hat{N}_{a}$ $=\sum_{i=1}^{N}\frac{1\{\bar{Z}_{i}=a\}}{P(\bar{Z}_{i}=a|T_{i})},\qquad a\in\{0,1\}.$ (32)
As will be shown in Theorem 1, the Hájek-weighted estimates will be lower bounded by the absolute value of a Hájek-weighted comparison of outcomes given by
$\Delta^{{\textup{Haj}}}=\sum_{i=1}^{N}\left(\frac{N}{\hat{N}_{1}}\frac{1\{\bar{Z}_{i}=1\}}{P(\bar{Z}_{i}=1|T_{i})}-\frac{N}{\hat{N}_{0}}\frac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=0|T_{i})}\right)Y_{i}$ (33)
For the estimands considered in this paper, Table 2 gives expressions for $\Delta^{{\textup{Haj}}}$.
Estimand $T_{i}$ $R_{i}$ $Z_{i}$
$\tau^{\textup{basic}}$ 1 $\{0,1\}^{N}$ $X_{i}$
$\tau^{\textup{basic}}$ using (9) 1 $\{0,1\}^{N}$ $1\{X_{i}=W_{i}=1\}-1\{X_{i}\neq W_{i}\}$
$\tau^{\textup{indirect}}$ $X_{i}$ $\{0,1\}^{N}$ $W_{i}$
$\tau^{\textup{nonneighbors}}$ $X_{\mathcal{N}_{i}}$ $\{0,1\}^{N}$ $W_{i}^{(2)}$
$\tau^{\textup{control}}$ 1 $\{X\in\{0,1\}^{N}:X_{i}=0\}$ $W_{i}$
$\tau^{\textup{tr}}$ 1 $\{X\in\{0,1\}^{N}:X_{i}=1\}$ $W_{i}$
Table 1: Exposure mappings, subsets, and contrast mappings for various estimands. Note that letting $T_{i}=1$ for all $X$ enforces $Y_{i}$ to equal a constant for all $X\in R_{i}$, if $i\in{\mathcal{I}}$.
Estimand Hájek-normalized contrast $\Delta^{\textup{Haj}}$
$\tau^{\textup{basic}}$ ${\sum_{i=1}^{N}\left(\frac{N}{\hat{N}_{1}}\frac{X_{i}}{P(X_{i}=1)}-\frac{N}{\hat{N}_{0}}\frac{1-X_{i}}{P(X_{i}=0)}\right)Y_{i}}$
$\tau^{\textup{basic}}$ (w/network) ${\sum_{i=1}^{N}\left(\frac{N}{\hat{N}_{1}}\frac{1\{X_{i}=W_{i}=1\}}{P(X_{i}=W_{i}=1)}-\frac{N}{\hat{N}_{0}}\frac{1\{X_{i}=W_{i}=0\}}{P(X_{i}=W_{i}=0)}\right)Y_{i}}$
$\tau^{\textup{indirect}}$ ${\sum_{i=1}^{N}\left(\frac{N}{\hat{N}_{1}}\frac{1\{W_{i}=1\}}{P(W_{i}=1|X_{i})}-\frac{N}{\hat{N}_{0}}\frac{1\{W_{i}=0\}}{P(W_{i}=0|X_{i})}\right)Y_{i}}$
$\tau^{\textup{nonneighbors}}$ ${\sum_{i=1}^{N}\left(\frac{N}{\hat{N}_{1}}\frac{1\{W_{i}^{(2)}=1\}}{P(W_{i}^{(2)}=1|X_{{\mathcal{N}}_{i}})}-\frac{N}{\hat{N}_{0}}\frac{1\{W_{i}^{(2)}=0\}}{P(W_{i}^{(2)}=0|X_{{\mathcal{N}}_{i}})}\right)Y_{i}}$
$\tau^{\textup{control}}$ ${\sum_{i=1}^{N}\left(\frac{N}{\hat{N}_{1}}\frac{1\{X_{i}=0,W_{i}=1\}}{P(X_{i}=0,W_{i}=1)}-\frac{N}{\hat{N}_{0}}\frac{1\{X_{i}=0,W_{i}=0\}}{P(X_{i}=0,W_{i}=0)}\right)Y_{i}}$
$\tau^{\textup{tr}}$ ${\sum_{i=1}^{N}\left(\frac{N}{\hat{N}_{1}}\frac{1\{X_{i}=1,W_{i}=1\}}{P(X_{i}=1,W_{i}=1)}-\frac{N}{\hat{N}_{0}}\frac{1\{X_{i}=1,W_{i}=0\}}{P(X_{i}=1,W_{i}=0)}\right)Y_{i}}$
Table 2: Hájek weighted point estimate $\hat{\tau}$, expressed as contrasts given by (33). The terms $\hat{N}_{1}$ and $\hat{N}_{0}$ are defined by (32).
Analogous to the discussion of Section 3, Proposition 3 gives intuition for when our lower bounds may be tight, and suggests guidance for the choice of $\bar{Z}$. It states that if for all $i\notin{\mathcal{I}}$ the contrast variable $\bar{Z}_{i}$ is a perfect predictor of the outcome $Y_{i}$ whenever $\bar{Z}_{i}\in\{0,1\}$, then an IPW weighted contrast using $\bar{Z}_{i}$ will be unbiased for the estimand.
Proposition 3.
Let $\Delta^{\bar{Z}}$ equal
$\Delta^{\bar{Z}}=\sum_{i=1}^{N}\left(\frac{1\{\bar{Z}_{i}=1\}}{P(\bar{Z}_{i}=1|T_{i})}-\frac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=0|T_{i})}\right)Y_{i}$
and for all $i\notin{\mathcal{I}}$ let $Y_{i}=\bar{Z}_{i}$ whenever $\bar{Z}_{i}\neq-1$. Then $\mathbb{E}[\Delta^{\bar{Z}}]=N-|{\mathcal{I}}|.$
4.3 Variance Estimation and Inference for General Case
Let $v_{i}$ denote
$v_{i}=\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{\bar{Z}_{i}=Y_{ki}\}}{P(\bar{Z}_{i}=Y_{ki}|T_{i})}$ (34)
so that $\hat{\tau}_{k}^{\textup{Haj}}=N-\sum_{i\in{\mathcal{I}}}v_{i}$. It follows that the variance of $\hat{\tau}_{k}^{\textup{Haj}}$ can be written as a sum of covariances
${\operatorname{Var}\,}\hat{\tau}_{k}^{\textup{Haj}}=\sum_{i\in{\mathcal{I}}}\sum_{j\in{\mathcal{I}}}{\operatorname{Cov}}(v_{i},v_{j})$
Unlike the special case where $T_{i}$ is constant for all units, generally we cannot compute ${\operatorname{Cov}}(v_{i},v_{j})$ exactly without knowledge of the outcome models $f_{i}$ and $f_{j}$, even if $i,j\in{\mathcal{I}}$. As a result, the covariance terms must be estimated. We thus propose variance estimator $\hat{V}_{k}$ given by
$\hat{V}_{k}=\sum_{i\in{\mathcal{I}}}\sum_{j\in{\mathcal{I}}}\Bigg\{\mathbb{E}\left[v_{i}v_{j}|T_{i},T_{j}\right]-\mathbb{E}\left[v_{i}|T_{i}\right]\cdot\mathbb{E}\left[v_{j}|T_{j}\right]\cdot\\ \frac{P(T_{i})\cdot P(T_{j})}{P(T_{i},T_{j})}\Bigg\}\cdot\frac{1\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}}{P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})}$ (35)
We will show in Theorem 2 that (35) is conservatively biased, with the exception that if $T_{i}$ is a constant for each unit (as is the case for $\tau^{\textup{basic}}$, $\tau^{\textup{control}}$, and $\tau^{\textup{tr}}$), then the estimator simplifies to a weighted sample of covariance terms
$\hat{V}_{k}=\sum_{i\in{\mathcal{I}}}\sum_{j\in{\mathcal{I}}}{\operatorname{Cov}}(v_{i},v_{j})\cdot\frac{1\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}}{P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj})},$
and is unbiased if the sampling probabilities $P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj})$ are non-zero, and conservative otherwise.
If $\hat{\tau}_{k}^{\textup{Haj}}$ for $k=1,2$ are consistent and asymptotically normal, with variance consistently (or conservatively) estimated by $\hat{V}_{k}$, then we can construct an asymptotic $(1-\alpha)$ level one-sided confidence interval for $\tau$, with lower boundary given by
$\max\left\{\min_{\phi\in\{0,1\}^{N}}\hat{\tau}_{1}^{\textup{Haj}}(\phi)-z_{1-\frac{\alpha}{2}}\sqrt{\hat{V}_{1}(\phi)}\,,\,\min_{\phi\in\{0,1\}^{N}}\hat{\tau}_{2}^{\textup{Haj}}(\phi)-z_{1-\frac{\alpha}{2}}\sqrt{\hat{V}_{2}(\phi)}\,\right\},$ (36)
where $\hat{\tau}_{k}^{\textup{Haj}}(\phi)$ and $\hat{V}_{k}(\phi)$ denote $\hat{\tau}_{k}^{\textup{Haj}}$ and $\hat{V}_{k}$, evaluated under the hypothesis that ${\mathcal{I}}=\{i:\phi_{i}=1\}$.
To evaluate the terms $\hat{\tau}_{k}^{\textup{Haj}}(\phi)$ and $\hat{V}_{k}(\phi)$ appearing in (36), we may use Propositions 4 and 5. Proposition 4 shows that the sampling weights $P(\bar{Z}_{i}=Y_{ki}|T_{i})$ appearing in $\hat{\tau}_{k}^{\textup{Haj}}$ given by (31) do not require knowledge of the outcome mappings $\{f_{i}\}_{i=1}^{N}$. Proposition 5 gives computable expressions for the terms appearing in $\hat{V}_{k}$ given by (35).
Proposition 4.
If $i\in{\mathcal{I}}$ and $X\in R_{i}$, then it holds that
$P(\bar{Z}_{i}=Y_{ki}|T_{i})=Y_{ki}\cdot P(\bar{Z}_{i}=1|T_{i})+(1-Y_{ki})\cdot P(\bar{Z}_{i}=0|T_{i}).$
Proposition 5.
If $i\in{\mathcal{I}}$ and $\bar{Z}_{i}=Y_{ki}$, then it holds that
$\mathbb{E}[v_{i}|T_{i}]$ $=\sum_{a=0}^{1}1\{Y_{ki}=a\}\mathbb{E}\left[\frac{1\{\bar{Z}_{i}=a\}N}{P(\bar{Z}_{i}=a|T_{i})\hat{N}_{a}}|T_{i}\right],$
If $i,j\in{\mathcal{I}}$ and $1\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}=1$, then it holds that
$\mathbb{E}[v_{i}v_{j}|T_{i},T_{j}]=$ $\sum_{a,b=0}^{1}1\{Y_{ki}=a,Y_{kj}=b\}\mathbb{E}\left[\frac{1\{\bar{Z}_{i}=a\}N}{P(\bar{Z}_{i}=a|T_{i})\hat{N}_{a}}\cdot\frac{1\{\bar{Z}_{j}=b\}N}{P(\bar{Z}_{j}=b|T_{j})\hat{N}_{b}}|T_{i},T_{j}\right]$
and also that
$P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})=$ $\sum_{a,b=0}^{1}1\{Y_{ki}=a,Y_{kj}=b\}P(\bar{Z}_{i}=a,\bar{Z}_{j}=b|T_{i},T_{j})$
5 Theory
In this section, we give consistency and coverage results for the general class of Hájek-normalized estimators described in Section 4. Specifically, we show that $|\Delta^{\textup{Haj}}|$ asymptotically lower bounds the estimand; we show that the variance estimator $\hat{V}_{k}$ concentrates and is either unbiased or conservative; and we show that our intervals have the desired asymptotic coverage, albeit with a technical modification to cover settings approaching degeneracy, in which case normality may not hold.
To describe our requirements on the experiment design, for $i\in[N]$ let $\Gamma_{i}\subseteq[N]$ denote a subset of units whose treatment assignment is sufficient to determine the values of $T_{i}$ and $\bar{Z}_{i}$ for all $X\in\{0,1\}^{N}$, so that for all $X,X^{\prime}\in\{0,1\}^{N}$ satisfying $X_{\Gamma_{i}}=X_{\Gamma_{i}}^{\prime}$, it holds that
$T_{i}(X)=T_{i}(X^{\prime}),\textup{ and }\bar{Z}_{i}(X)=\bar{Z}_{i}(X^{\prime})$
We will call $\Gamma_{i}$ the design neighborhood of unit $i$. We stress that requirements on $\{\Gamma_{i}\}_{i=1}^{N}$ restrict the experiment designs and estimands which may be explored, but places no assumptions on the outcome models $\{f_{i}\}_{i=1}^{N}$ that describe interference, which are unknown and uncontrolled by the researcher.
Assumption 1 describes a simple experiment design that we study. It assumes a sequence of experiments in which the design neighborhoods are bounded in size and overlap, Bernoulli randomization with treatment probabilities bounded away from 0 and 1, and places a positivity condition on $P(\bar{Z}_{i}|T_{i})$.
Assumption 1.
The treatment assignments $\{X_{i}\}_{i=1}^{N}$ are independent Bernoulli random variables, which satisfy for fixed constants $p>0$ and $d>0$ as $N\rightarrow\infty$ that
$|\Gamma_{i}|$ $\leq d$ $\forall\ i\in[N]$ (37)
$\sum_{j=1}^{N}1\{i\in\Gamma_{j}\}$ $\leq d$ $\forall\ i\in[N]$ (38)
$p\leq P(X_{i}=1)$ $\leq 1-p$ $\forall\ i\in[N]$ (39)
$P(\bar{Z}_{i}=a\,|\,T_{i}=t)$ $\geq p$ $\forall\ i\in[N],a\in\{0,1\},t\in\mathcal{T}_{i}$ (40)
where $\mathcal{T}_{i}$ denotes the support of $T_{i}$.
The positivity condition (40) requires $T_{i}$, $Z_{i}$, and $R_{i}$ to be chosen so that $Z_{i}$ and $R_{i}$ are not deterministic conditional on $T_{i}$. For example, if $T_{i}=X_{\mathcal{N}_{i}}$, then $Z_{i}$ and $R_{i}$ should depend at least partially on units outside of $\mathcal{N}_{i}$.
Theorem 1 is a consistency result. It states that $\hat{\tau}^{\textup{Haj}}_{k}$ converges to $\tau$, and that the absolute value of $\Delta^{\textup{Haj}}$, the Hájek weighted contrast between units given by (33), is asymptotically a lower bound for $\tau$.
Theorem 1.
It holds that
$|\Delta^{\textup{Haj}}|$ $\leq\max(\hat{\tau}_{1}^{\textup{Haj}},\hat{\tau}_{2}^{\textup{Haj}})$ (41)
If Assumption 1 holds, then
$\hat{\tau}_{k}^{\textup{Haj}}=\tau+O_{P}\left(N^{1/2}\right),\qquad k=1,2$ (42)
in which case (41) implies $|\Delta^{\textup{Haj}}|\leq\tau+O_{P}\left(N^{1/2}\right)$
Theorem 2 pertains to variance estimation. It states that $\hat{V}_{k}$ concentrates to its expectation which upper bounds the variance of $\hat{\tau}_{k}^{\textup{Haj}}$, and gives conditions under which the bound will be tight.
Theorem 2.
Let Assumption 1 hold. It holds that
$\hat{V}_{k}=\mathbb{E}[\hat{V}_{k}]+O_{P}\left(N^{1/2}\log N\right),\qquad k=1,2.$ (43)
Additionally, it holds that
$\mathbb{E}\hat{V}_{k}\geq{\operatorname{Var}\,}\hat{\tau}_{k}^{\textup{Haj}},$ (44)
with equality if all of the following conditions hold:
  1. 1.
    The exposure mappings $\{T_{i}\}_{i=1}^{N}$ are constant
  2. 2.
    $P(\bar{Z}_{i}=a)>0$ for all $i\in[N]$ and $a\in\{0,1\}$
  3. 3.
    $P(\bar{Z}_{i}=a,\bar{Z}_{j}=b)>0$ for all $i,j\in[N]$ such that $i\neq j$, and $a,b\in\{0,1\}$.
Theorem 3 gives coverage properties for the proposed confidence interval given by (36). It states that if the variance estimators $\hat{V}_{k}$ are lower bounded by a vanishing fraction of $N$, then the interval will have at least the desired asymptotic coverage level.
Theorem 3.
Let Assumption 1 hold. For any $\epsilon>0$, let $\widetilde{V}_{k}(\phi)$ denote the thresholded variance estimate given by
$\widetilde{V}_{k}(\phi)=\max\left\{\hat{V}_{k}(\phi)\ ,\ \frac{N^{2/3+\epsilon}}{z_{1-\frac{\alpha}{2}}^{2}\cdot\alpha/2}\right\}$ (45)
and let $LB_{1-\alpha}$ denote
$LB_{1-\alpha}=\max\left\{\min_{\phi\in\{0,1\}^{N}}\hat{\tau}_{1}^{\textup{Haj}}(\phi)-z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{1}(\phi)}\,,\,\min_{\phi\in\{0,1\}^{N}}\hat{\tau}_{2}^{\textup{Haj}}(\phi)-z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{2}(\phi)}\right\}$ (46)
Then $\tau\geq LB_{1-\alpha}$ with probability converging to at least $1-\alpha$.
In practice, we expect usage of $\tilde{V}_{k}(\phi)$ rather than $\hat{V}_{k}(\phi)$ to make little or no difference, as the two terms are equal except in cases involving interval half-widths of order $O(N^{1/3})$ or smaller. This is a vanishing fraction of the usual $O(\sqrt{N})$ rate that arises for estimation of population sizes under simple random sampling, and we do not expect the interval widths found by (46), which are widened by the optimization over $\phi$, to be so small in practice.
Additional Experiment Designs
We have considered independent Bernoulli randomization primarily as a starting point for its simplicity. In the supplemental materials, we generalize the theorems to include two-stage experiment designs, in which the size of the design neighborhoods may grow with $N$. The result will allow contrasts (i.e., $\bar{Z}$) involving not only the treatment of each unit and a bounded number of close friends, but also the number of treated units in some broader division such as geographic region, or overlapping collections of neighboring geographic regions, in which the region sizes may grow at rate $O(N^{1-\delta})$ for any $\delta>0$.
6 Examples
6.1 Simulation Study
Table 3 shows the results of a simulation study in which the outcomes depended on both direct and indirect treatment. For all estimands except $\tau^{\textup{nonneighbors}}$, outcomes were generated according to
$Y_{i}=1\{\textup{if }u_{i}\leq\alpha_{0i}+\alpha_{1i}\cdot X_{i}+\alpha_{2i}\cdot W_{i}\}$ (47)
where the scalar parameters $u_{i},\alpha_{0i},\alpha_{1i},\alpha_{2i}$ were randomly generated, with $u_{i}\sim\textup{Unif }[0,1]$, and $\alpha_{0i},\alpha_{1i},\alpha_{2i}$ nonnegative. For simulations involving $\tau^{\textup{nonneighbors}}$, in (47) we change $W_{i}$ to $\bar{W}_{i}$, the thresholded number of treated units in ${\mathcal{N}}_{i}\cup{\mathcal{N}}_{i}^{(2)}$ excluding unit $i$, so that units in ${\mathcal{N}}_{i}$ and ${\mathcal{N}}_{i}^{(2)}$ were equally important in determining the outcome of unit $i$. To compute the true value of the estimand, unit $i$ was considered to be affected by treatment if $Y_{i}$ varied as a function of $X$ under the generated parameter values, and was considered to be affected by treatment of others if $Y_{i}$ varied as a function of $X_{-i}$ for any value of $X_{i}$, and so forth.
Estimation of $\tau^{\textup{basic}}$ used the contrast variable $Z$ given by (9). Hájek normalization was used on all estimators. For comparison, the treatment effect
$EATE=\frac{1}{N}\sum_{i=1}^{N}\Big(\mathbb{E}[Y_{i}|\bar{Z}_{i}=1]-\mathbb{E}[Y_{i}|\bar{Z}_{i}=0]\Big),$
was also estimated, where $\bar{Z}$ followed Table 1 for each estimand. For example, when estimating $\tau^{\textup{indirect}}$, we compared to the indirect effect given by
$\frac{1}{N}\sum_{i=1}^{N}\Big(\mathbb{E}[Y_{i}|W_{i}=1]-\mathbb{E}[Y_{i}|W_{i}=0]\Big).$
For estimation and inference of these treatment effects, Hájek normalization and conservative variance estimates were used following Gao and Ding, (2023). [^5] Variance estimates for EATE followed Eq. (11) (“adjusted HAC covariance estimator”) of Gao and Ding, (2023) (arxiv version v3), using our $\bar{Z}_{i}$ for their treatment condition $T_{i}$. For the simulations, the correctly specified bandwidth $b_{n}=2$ was used.
The results of the simulations are the following. For the estimands $\tau^{\textup{basic}}$, $\tau^{\textup{tr}}$, and $\tau^{\textup{control}}$, point estimates were unbiased (as might be expected given that the unit effects were nonnegative, consistent with the intuition of Example 1), and the confidence intervals exhibited approximate calibration of coverage. For these estimands, the exposure mapping $T_{i}$ is constant, allowing for unbiased variance estimation as discussed in Theorem 2. For the remaining estimands, estimation was conservatively biased, and interval coverage rates were 100%. Table 3 also shows that the interval half-widths for our estimands were competitive, and often smaller than, those of comparable EATE treatment effects. Coverage of the EATE interval estimates was conservative, always exceeding 99%.
Our intuition for the smaller intervals of our approach is the following. Whereas treatment effect estimation uses $\Delta^{\textup{Haj}}$ as a point estimate and therefore must control its variability, we use $\Delta^{\textup{Haj}}$ rather as a lower bound for the point estimators $\hat{\tau}_{1}^{\textup{Haj}}$ and $\hat{\tau}_{2}^{\textup{Haj}}$. Thus, to control anticonservativeness we need to consider the variability of these point estimators, which we expect to be less than that of $\Delta^{\textup{Haj}}$ and also easier to estimate, as only units in $\mathcal{I}$} are considered.
Actual $|\Delta^{\textup{Haj}}|/N$ Coverage CI Half- $\widehat{EATE}$ EATE CI
Estimand N Value/N/N Avg SD of 95% CI width Avg SD Half-width
$\tau^{\textup{basic}}$ 250 0.43 0.43 0.08 0.94 0.16 0.43 0.08 0.26
500 0.37 0.37 0.05 0.96 0.11 0.36 0.05 0.21
$\tau^{\textup{tr}}$ 250 0.22 0.22 0.08 0.95 0.15 0.22 0.08 0.26
500 0.18 0.18 0.05 0.95 0.11 0.18 0.05 0.20
$\tau^{\textup{control}}$ 250 0.23 0.23 0.10 0.95 0.16 0.23 0.10 0.29
500 0.21 0.21 0.06 0.96 0.12 0.21 0.06 0.23
$\tau^{\textup{indirect}}$ 250 0.37 0.22 0.06 1.0000 0.15 0.22 0.06 0.20
500 0.33 0.20 0.04 1.0000 0.10 0.19 0.04 0.16
$\tau^{\textup{nonneighbor}}$ 250 0.37 0.22 0.07 1.0000 0.21 0.22 0.06 0.20
500 0.33 0.20 0.04 1.0000 0.18 0.19 0.04 0.16
Table 3: Results of Simulation Study. Bold values indicate approximate 95% coverage, or that CI half-widths were smaller than those of comparable treatment effect. CI Half-width defined as average distance from point estimate to the interval boundary. For our estimands, $|\Delta^{\textup{Haj}}|/N$ was point estimate. 500 trials per simulation.
6.2 Social Networks and the Decision to Insure
We analyze an experiment that was originally described in Cai et al., (2015), and re-analyzed in Gao and Ding, (2023).666data available at https://www.openicpsr.org/openicpsr/project/113593/version/V1/view In this experiment, rural farmers were randomized to receive high or low information regarding a crop insurance product, with opportunity to purchase the product at the end of the information session. Before attending their assigned sessions, the farmers could communicate informally with others who had attended earlier sessions (sessions were administered in two rounds, with a three day gap between them), allowing for sharing of information through social contacts. Social network information was elicited, with the farmers instructed to list 5 close friends with whom they specifically discussed rice production or financial issues. Farmers assigned to a second round low-information session were more likely to purchase insurance if more of their listed friends in the first round were assigned to a high-information session.
Table 8 in the supplement shows results from the original regression analysis of Cai et al., (2015), which found that for farmers assigned to to low information sessions, each friend assigned to a first round high-information session caused an 8.9% increase in insurance purchasing rates (95% CI: [4.5%, 13.2%]). In their analysis, it was assumed that a unit’s outcome depends only on the information received by themselves and their listed friends. To justify this assumption, a subjective argument was made that the three day gap between information sessions was long enough for information to reach direct friends, but too short for information to diffuse across an entire village.
In our analysis, we make no assumptions on interference and estimate $\tau^{\textup{control}}$, the number of second round units whose outcome under direct control (i.e., assignment to a low info session) would be affected by the assignment of first round farmers into high and low information sessions. Interference is permitted between distant or even disconnected units; for example, a farmer might ask a question during an information session, affecting the information received by (and subsequently shared by) all farmers at the session. To construct our estimators, we require knowledge of the randomization design, which used stratification and is described in Gao and Ding, (2023).
Our asymptotic lower bound equals 23%, which may be interpreted as meaning that at least 23% of the 880 second round farmers, if assigned to a low information session, would be affected by information given to the first round farmers (1-sided 95% CI: at least 9%). This point estimate of 23% is equal to a Hájek weighted comparison of the second round units receiving low information, grouped according to their number of friends receiving high information in the previous round:
$\frac{\Delta^{\textup{Haj}}}{N}=\frac{1}{N}\sum_{i=1}^{N}\left(\frac{1\{X_{i}=0,W_{i}=1\}N}{P(X_{i}=0,W_{i}=1)\hat{N}_{1}}-\frac{1\{X_{i}=0,W_{i}=0\}N}{P(X_{i}=0,W_{i}=0)\hat{N}_{0}}\right)Y_{i}$ (48)
Here $i\in[N]$ enumerates the $N=880$ second round units with at least one first round friend; $X_{i}=0$ indicates that unit $i$ was assigned to a low information session; $W_{i}=1$ indicates that all of unit $i$’s first round friends received high information; $Y_{i}=1$ indicates that unit $i$ purchased insurance; and $\hat{N}_{a}=\sum_{i=1}^{N}1\{X_{i}=0,W_{i}=a\}/P(X_{i}=0,W_{i}=a)$ for $a\in\{0,1\}$. Additional results can be found in Appendix B of the supplemental materials.
Table 7 in the supplement shows results from Gao and Ding, (2023), who estimate the spillover effects of having friends receive high information. They report weakly significant or non-significant spillover effects of the first round assignments on the second round units, assuming conditions on interference which are technical in nature. It may be difficult to judge whether these conditions are suitable for this application. In contrast, no assumptions on interference are required for our estimates. For the purposes of demonstrating the presence of social influence, our estimand may be an appropriate target parameter, and has tighter, less questionable CIs. Thus it may serve as a useful complement to prior studies of this experiment.
Supplemental Materials
The supplement contains discussion of future work; further discussion and results for the simulations and data example; and proofs of all propositions and theorems, with generalization of Theorems 1 - 3 to include two-stage randomization.
Supplementary Material for “Agnostic Characterization of Interference in Randomized Experiments”
Appendix A Directions for Future Work
In principle, our approach can be extended to settings with categorical or continuous outcomes in various ways. The simplest approach is simply to binarize the outcomes, requiring a choice of threshold. Under binarization, our estimates will continue to lower bound the quantity of interest. For example, $\tau^{\textup{basic}}$ after outcome binarization denotes the number of units whose binarized outcome is affected by treatment, which is a subset of the units whose original outcome (i.e., pre-binarization) is affected by treatment.
To avoid thresholding of continuous-valued outcomes, we conjecture that rank-based estimators reminiscent of Rosenbaum, (2007) may also be devised, or kernel-based density estimates of the outcomes conditional on receiving treatment or control. Under the latter approach, the total variation difference between the conditional densities takes the place of the diff-in-means contrast $\Delta$, although we are unsure whether the optimization approach of (46) can be successfully adapted for inference.
We expect that our approach can be adapted to most experiment designs that have been considered for estimation of treatment effects under the assumption of a correctly specified exposure model. This is because consistency and normality theorems that require a correctly specified exposure model (normally a strong assumption) can be applied to $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ without question, as these estimators only consider units in $\mathcal{I}$}, for which the exposure model condition holds by definition.
Experiments involving group randomization or random edge formation may be of particular interest. As previously mentioned, for such experiments $\tau^{\textup{basic}}$ characterizes interference, as it denotes the number of units who are affected by their randomly assigned neighbors or group members. A possible approach might be similar to Li et al., (2019) which studied peer effects, while removing assumptions regarding interference that were required by that work. Let $A\in\{0,1\}^{N\times N}$ denote a random adjacency matrix whose distribution is jointly exchangeable following Bickel and Chen, (2009), so that $A$ and $\Pi A\Pi^{T}$ are equally probable for any permutation matrix $\Pi\in\{0,1\}^{N\times N}$. For group randomization experiments, the adjacency matrix $A$ would represent the randomly assigned groups as a collection of cliques, while edge formation experiments could use various network models, such as latent space or stochastic blockmodel (Bickel and Chen, , 2009; Hoff et al., , 2002). Let $O=(O_{1},\ldots,O_{N})\in\{0,1\}^{N}$ denote binary pre-treatment covariates for the $N$ units, with covariate classes given by $\mathcal{O}_{a}=\{i:O_{i}=a\}$ for $a\in\{0,1\}$. Let the treatment variables $X=(X_{1},\ldots,X_{N})$ denote the thresholded number of neighbors whose observable level is 1, so that
$X_{i}=1\left\{\sum_{j=1}^{N}A_{ij}O_{j}\geq t_{O_{i}}\right\}$
where $t_{0}$ and $t_{1}$ denote thresholds for units in $\mathcal{O}_{0}$ and $\mathcal{O}_{1}$ respectively. We conjecture that conditional on the number of units in $\mathcal{O}_{0}$ and $\mathcal{O}_{1}$ who were treated (i.e., for whom $X_{i}=1$), inference will take simple form in which the conditional distribution of $X$ is given by sampling without replacement within $\mathcal{O}_{k}$ for $k=0,1$. We additionally conjecture that this property will facilitate adaptation of the necessary asymptotic results, given by Theorems 1 - 3, to this setting.
Appendix B Additional Notes on Examples
Simulations
Choi, (2023) gives a proof-of-concept attempt to estimate $\tau^{\textup{indirect}}$, which was tested on the simulations described in Section 6.1. The approach was found to be lacking in power compared to ours, with 1-sided intervals whose boundaries were considerably closer to zero (and included zero considerably more often), as summarized in Table 4.
Data Example
For the estimation of $\tau^{\textup{control}}$ in the data analysis, $W_{i}$ was defined to equal 1 if all of unit $i$’s first round friends were assigned to high info sessions. This choice of threshold for $W_{i}$ was inspired by Choi, (2024), who found an effect threshold at 2 treated (i.e., first round high info) friends. Thresholding $W_{i}$ at 2 treated friends gives similar results for $\tau^{\textup{control}}$ (point estimate: 22%, 1-sided 95% CI: at least 8.8%), but restricts analysis to units with at least 2 first round friends. Thresholds exceeding 2 would remove most of the units from the analysis. These results are summarized in Table 5, which shows estimates of $\tau^{\textup{control}}$ for varying choices of threshold and under Bonferoni correction (which did not cause the intervals to become weakly significant or non-significant).
Choi, (2023) Our approach
NN P(Significant) Avg. $N^{-1}LB_{1-\alpha}$ bias P(Significant) Avg. $N^{-1}LB_{1-\alpha}$ bias
$N=250$ 5% 0.0005 -0.26 80% 0.076 -0.15
$N=500$ 20% 0.0027 -0.23 97% 0.093 -0.13
Table 4: Comparison of Choi, (2023) and our approach for estimating $\tau^{\textup{indirect}}$ on simulated data. P(Significant) denotes the fraction of trials for which the 1-sided interval excluded zero. Avg. $N^{-1}LB_{1-\alpha}$ denotes the average value of the lower boundary (as a fracton of $N$) of the 1-sided interval. Bias reports the average difference (as a fraction of $N$) between the point estimator and $\tau^{\textup{indirect}}$.
Threshold $t_{i}$ Point Estimate 1-sided 95% CI Bonferoni corrected
$\geq 1$ treated friend 8.8% at least 0% at least 0%
$\geq 2$ treated friends 22% at least 8.8% at least 5.7%
All 1st round friends treated 23% at least 9.9% at least 7.0%
Table 5: Estimates of $\tau^{\textup{control}}$ for data example of Cai et al., (2015), under varying choices of threshold $t_{i}$. Bonferoni correction is for 3 tests ($\approx 98.33\%$ confidence level).
Our result is significantly stronger than the proof-of-concept attempt of Choi, (2024) to estimate the number of second round units who were affected by the first round treatments. Their analysis did not produce a point estimate, and gave a much more conservative value of 6 units (or 2% of a matched subset of units) for the boundary of the 1-sided confidence interval, which is smaller than ours by a factor of 13. The approach of Choi, (2023) also performs poorly, producing a conservative point estimate of at least 6 affected units, which was not statistically significant (95% CI: at least zero units).
We may also compare to an EATE-type treatment effect that considers the relative effects of receiving $(X_{i},W_{i})$ equal to $(0,1)$ versus $(0,0)$:
$\textup{treatment effect}=\frac{1}{N}\sum_{i=1}^{N}\left(\mathbb{E}[Y_{i}|X_{i}=0,W_{i}=1]-\mathbb{E}[Y_{i}|X_{i}=0,W_{i}=0]\right),$ (49)
where the expectation is taken over the randomization of treatment under the experiment design. For this target parameter, the estimation method of Gao and Ding, (2023) gives a Hájek-normalized point estimate of 24%, with a 95% CI ranging in width from $1.4\%,43\%]$ to $5.3\%,40\%]$ depending on the choice of kernel matrix $K_{n}$, as summarized in Table 6.
The role of the kernel matrix $K_{n}$ is to choose which pairs of units should be used by the variance estimator of Eq. (11) in Gao and Ding, (2023). Given an observed network, they propose a bandwidth rule to balance between bias (from excluding important pairs of units whose residuals are strongly correlated) and variance (from including too many unit pairs). In this rule, pairs of units are used for variance estimation if the length of the shortest path between them is less than or equal to a parameter $b_{n}$. However, it seems difficult to discern when one choice of $b_{n}$ will be more accurate than another, or if any of the tested choices will necessarily be accurate for the fixed size of the data at hand. As a result, the correctness of our intervals for (49) may be questionable.
Table 7 shows selected results from the analysis of Gao and Ding, (2023). It shows point estimates and estimated standard errors for the effect on insurance takeup of having at least one first round friend assigned to a high information session. Under all choices of $b_{n}$ exceeding zero, the effect is not significant at 95% confidence level. [^7] Gao and Ding, (2023) also report weak significance using uncorrected variance estimates (which may be anticonservative), in Tables S4 and S5 of their paper. Unlike our analysis, Gao and Ding, (2023) threshold the number of treated friends at 1, and includes second round units who were assigned to high info sessions. Changing both of these choices yields the above-mentioned effect of (49).
Table 8 shows selected regression results from the original analysis of Cai et al., (2015), which we include for reference.
Software
We used the commerical solver Gurobi, through MATLAB after installing the optimization toolbox YALMIP (Lofberg, , 2004).
Choice of $K_{n}$ (i.e., criteria for $K_{n}(i,j)=1$ to hold) 95% CI
$b_{n}=2$ (Units $i$ and $j$ have path distance $\leq 2$): $1.7\%\,,43\%]$
$b_{n}=4$ (Units $i$ and $j$ have path distance $\leq 4$): $2.8\%\,,42\%]$
Units $i$ and $j$ in the same village, or have friends
      in the same village: $5.3\%\,,40\%]$
Units $i$ and $j$ in same strata, or have friends
      in the same strata: $1.4\%\,,43\%]$
Table 6: 95% CI for EATE estimate of (49), using Eq. (10) of Gao and Ding, (2023) under varying choices of kernel matrix $K_{n}$. The last choice of $K_{n}$ (same strata/friends in same strata) captures dependencies that arise when treatments are assigned by stratified random sampling without replacement, or equivalently when inference is conditional on the stratified treatment rates.
Unadj Add Sat
Point Estimate: 0.056 0.058 0.055
std. err, $b_{n}=0$: 0.027 0.027 0.027
std. err, $b_{n}=2$: 0.033 0.033 0.033
std. err, $b_{n}=3$: 0.031 0.031 0.031
std. err, $b_{n}=4$: 0.031 0.031 0.031
std. err, $b_{n}=5$: 0.032 0.032 0.032
Table 7: Estimated spillover effects reported by Table S5 of Gao and Ding, (2023), for “two-dimensional exposure mapping” model and “WLS+ SE” standard errors. Regression specifications, where $z_{i}$ and $x_{i}$ respectively denote treatment and covariate variables, are as follows. Unaj: $Y_{i}\sim 1+z_{i}$. Add: $Y_{i}\sim 1+z_{i}+x_{i}$. Sat: $1+z_{i}+(x_{i}-\bar{x})+z_{i}(x_{i}-\bar{x})$.
Regression Variable (2) (4)
Invited to high info session: 0.0298 0.0809
(0.0332) (0.0397)
Fraction of friends invited to 1st round high info session (NET): 0.291*** 0.444***
(0.0820) (0.109)
(NET)*(invited to high info session): -0.329**
Table 8: Regression results reported by Table 2 of Cai et al., (2015) for specifications (2) and (4). For the effect of treating a single friend, Cai et al., (2015) divided the NET coefficient by 5, the number of elicited friends.
Appendix C Asymptotic Results
The organization of this section is as follows. In Section C.1, we present Assumption 2, a more general version of Assumption 1 which allows other experiment designs to be considered. We also present Theorems 4 and 5 which respectively describe conditions under which Bernoulli randomization designs and two-stage randomization designs will satisfy the assumption. In Section C.2, we present Theorems 68, which are generalized versions of Theorems 13, and respectively give results on consistency, variance estimation, and interval coverage under Assumption 2. Proofs of all propositions and theorems are given in Section D.
C.1 Experiment Designs
Assumption 2 describes a class of experiment designs, which includes independent Bernoulli randomization and two stage randomization designs. It assumes that units can be divided into $M$ groups $G_{1},\ldots,G_{M}$, such that the subvectors $X_{G_{1}},\ldots,X_{G_{M}}$ of $X$ are independent random variables.
Assumption 2.
The units may be partitioned into $M$ non-overlapping groups $G_{1},\ldots,G_{M}\subset[N]$ such that the subvectors $X_{G_{1}},\ldots,X_{G_{M}}$ are independent random variables, and the following hold:
  1. 1.
    For fixed constants $C>0$ and $0<\delta\leq 1$ it holds as $N\rightarrow\infty$ that
    $|G_{m}|\leq C\frac{N}{M}\qquad\textup{ and }\qquad|G_{m}|\leq N^{1-\delta}\qquad\forall m\in[M],$ (50)
  2. 2.
    There exists group neighborhoods $\eta_{1},\ldots,\eta_{M}\subset[M]$ satisfying for fixed $d>0$ that
    $\max_{m}|\eta_{m}|$ $\leq d$ (51)
    $\max_{\ell}\sum_{m=1}^{M}1\{\ell\in\eta_{m}\}$ $\leq d$ (52)
    such that the design neighborhoods $\{\Gamma_{i}\}_{i=1}^{N}$ satisfy
    $\Gamma_{i}=\bigcup_{m\in\eta_{g(i)}}G_{m},$ (53)
    where $g(i)$ denotes the group of unit $i$ (that is, $g(i)=m$ if $i\in G_{m}$).
  3. 3.
    For a fixed constant $\rho>0$ it holds for all $i,j\in[N]$ and $a,b\in\{0,1\}$ that
    $P(\bar{Z}_{i}=a|T_{i}=t_{i})$ $\geq\rho$ $\forall t_{i}\in\mathcal{T}_{i}$ (54)
    $\frac{P(T_{i}=t_{i})P(T_{j}=t_{j})}{P(T_{i}=t_{i},T_{j}=t_{j})}$ $\leq\frac{1}{\rho}$ $\forall(t_{i},t_{j})\in\mathcal{T}_{ij}$ (55)
    $P(\bar{Z_{i}}=a,\bar{Z}_{j}=b|T_{i}=t_{i},T_{j}=t_{j})$ $\in\{0\}\cup[\rho,1]$ $\forall(t_{i},t_{j})\in\mathcal{T}_{ij}$ (56)
    where $\mathcal{T}_{i}$ denotes the support of $t_{i}$, and $\mathcal{T}_{ij}$ denotes the support of $(t_{i},t_{j})$
The conditions of Assumption 2 may be interpreted as follows. Condition (50) requires the size of the largest group to be within a constant factor of the average group size, and allows the group sizes to grow sublinearly as $N\rightarrow\infty$. Condition (53) restricts the design neighbornoods $\{\Gamma_{i}\}_{i=1}^{N}$ to equal equal each unit’s own and neighboring groups, where the group neighborhoods $\{\eta_{m}\}_{m=1}^{M}$ are required by (51)-(52) to have bounded size and overlap. Conditions (54)-(56) are support conditions, which bound the magnitude of the inverse probability weights used by $\hat{\tau}_{k}^{\textup{Haj}}$ and $\hat{V}_{k}$.
Example: Bernoulli Design
Theorem 4 states that the setting of Assumption 1, in which the treatments $\{X_{i}\}_{i=1}^{N}$ are independent Bernoulli random variables, can be modeled under Assumption 2 by assigning each unit to its own group, so that $M=N$.
Theorem 4.
Let Assumption 1 hold. Then Assumption 2 holds with $M=N$.
Example: Two-stage Design
Theorem 5 gives conditions under which $|\Gamma_{i}|$ may grow at rate $N^{1-\delta}$ for any $\delta>0$, in which the treatments are randomized by a two-stage design. In this setting, the units are divided into groups, and treatment is randomized first at the group level, and then at the unit level conditional on group level treatment. The contrast variable $\bar{Z}_{i}$ may consider each unit’s direct treatment, the treatment of their close associates $\mathcal{N}_{i}$, and the number of treated units in their own and neighboring groups, where the group sizes may grow with $N$ at any sublinear polynomial rate.
Theorem 5.
Let $g(i)$ denote the group membership of unit $i$, so that $g(i)=m$ if $i\in G_{m}$. Assume the following hold for fixed $C>0$, $0<\delta\leq 1$, $d>0$, and $p>0$ as $N\rightarrow\infty$
  1. A1.
    The groups $G_{1},\ldots,G_{M}$ form a partition of $[N]$ satisfying (50), with group neighborhoods $\{\eta_{m}\}_{m=1}^{M}$ which satisfy the size and overlap conditions (51) and (52).
  2. A2.
    Each treatment subvector $X_{G_{m}}$ is independently generated by sampling $\theta_{m}$ units without replacement from $G_{m}$, where $\theta_{1},\ldots,\theta_{M}$ are independent random variables with support satisfying
    $\operatorname{supp}(\theta_{m})\subseteq\{0\}\cup\{\theta_{m}^{\textup{low}},\ldots,\theta_{m}^{\textup{high}}\}\cup\{|G_{m}|\}$ (57)
    for $\theta_{m}^{\textup{low}}\geq p|G_{m}|$ and $\theta_{m}^{\textup{high}}\leq(1-p)|G_{m}|$, and which additionally satisfy
    $P(\theta_{m}=t)\geq p\qquad\forall t\in\operatorname{supp}(\theta_{m})$ (58)
  3. A3.
    For $i\in[N]$, $\bar{Z}_{i}$ satisfies
    $\bar{Z}_{i}$ $=\zeta_{i}(X_{\mathcal{N}_{i}},\theta_{\eta_{g(i)}}),$ $\forall i\in[N],$ (59)
    for some set of mappings $\{\zeta_{i}\}_{i=1}^{N}$ and unit neighborhoods $\{\mathcal{N}_{i}\}_{i=1}^{N}$. Each neighborhood $\mathcal{N}_{i}$ is a subset of unit $i$’s own and neighboring groups so that
    $\mathcal{N}_{i}\subset\bigcup_{m\in\eta_{g(i)}}G_{m}$ (60)
    holds, and also satisfies $i\in\mathcal{N}_{i}$ and the size and overlap conditions
    $\max_{i}|\mathcal{N}_{i}|\leq d\qquad\textup{ and }\qquad\max_{i}\sum_{j=1}^{N}1\{i\in\mathcal{N}_{j}\}\leq d$ (61)
  4. A4.
    Either $T_{i}=1$ for all $i\in[N]$, or $T_{i}=X_{i}$ for all $i\in[N]$, or $T_{i}=X_{\mathcal{N}_{i}}$ for all $i\in[N]$
  5. A5.
    It holds that
    $P(\bar{Z}_{i}=a\,|\,T_{i}=t)$ $\geq p$ $\forall\ i\in[N],a\in\{0,1\},t\in\mathcal{T}_{i}$ (62)
    where $\mathcal{T}_{i}$ is the support of $T_{i}$.
Then Assumption 2 holds. Additionally, Assumption 2 continues to hold if Conditions A3 and A4 are replaced by Condition A6, given below:
  1. A6.
    For $i\in[N]$, $\bar{Z}_{i}$ and $T_{i}$ satisfy
    $\bar{Z}_{i}$ $=\zeta_{i}(\theta_{\eta_{g(i)}})\qquad\textup{ and }$ $T_{i}$ $=X_{G_{g(i)}},$ $\qquad\forall i\in[N],$ (63)
    for some set of mappings $\{\zeta_{i}\}_{i=1}^{N}$.
The conditions Theorem 5 may be interpreted as follows. The conditions (57) and (58) disallows treatment of a vanishing fraction of units, and requires the group level treatment probabilities to be bounded away from zero. The condition (59) requires $\bar{Z}_{i}$ to depend on each unit’s own treatment and the treatment of a bounded number of close associates, and on the number of treated units in their own and neighboring groups. (60) restricts $\mathcal{N}_{i}$ to come from unit $i$’s own and neighboring groups, with size and overlap conditions given by (61). (62) places a support condition on $\bar{Z}_{i}$. (63), which replaces conditions A3 and A4, allows the estimand to count the number of units who are affected by treatments outside of their own group, whose size may increase with $N$.
C.2 Main Theorems
Theorem 6 is a consistency result. It states that $\hat{\tau}^{\textup{Haj}}_{k}$ converges to $\tau$, and that the absolute value of $\Delta^{\textup{Haj}}$, the Hájek weighted contrast between units given by (33), is asymptotically a lower bound for $\tau$. Table 2 gives examples of (33) for the estimands considered in the paper.
Theorem 6.
It holds that
$|\Delta^{\textup{Haj}}|$ $\leq\max(\hat{\tau}_{1}^{\textup{Haj}},\hat{\tau}_{2}^{\textup{Haj}})$ (64)
If Assumption 2 holds, then
$\hat{\tau}_{k}^{\textup{Haj}}=\tau+O_{P}\left(\frac{N}{\sqrt{M}}\right),\qquad k=1,2$ (65)
in which case (64) implies $|\Delta^{\textup{Haj}}|\leq\tau+O_{P}\left(\frac{N}{\sqrt{M}}\right)$
Theorem 7 pertains to variance estimation. It states that $\hat{V}_{k}$ concentrates to its expectation which upper bounds the variance of $\hat{\tau}_{k}^{\textup{Haj}}$, and gives conditions under which the bound will be tight.
Theorem 7.
Let Assumption 2 hold. It holds that
$\hat{V}_{k}=\mathbb{E}[\hat{V}_{k}]+O_{P}\left(\frac{N^{2}\log M}{M^{3/2}}\right),\qquad k=1,2.$ (66)
Additionally, it holds that
$\mathbb{E}\hat{V}_{k}\geq{\operatorname{Var}\,}\hat{\tau}_{k}^{\textup{Haj}},$ (67)
with equality if all of the following conditions hold:
  1. 1.
    The exposure mappings $\{T_{i}\}_{i=1}^{N}$ are constant
  2. 2.
    $P(\bar{Z}_{i}=a)>0$ for all $i\in[N]$ and $a\in\{0,1\}$
  3. 3.
    $P(\bar{Z}_{i}=a,\bar{Z}_{j}=b)>0$ for all $i,j\in[N]$ such that $i\neq j$, and $a,b\in\{0,1\}$.
Theorem 8 gives coverage properties for the proposed confidence interval given by (36). It states that if the variance estimators $\hat{V}_{k}$ are lower bounded by a vanishing fraction of $\frac{N^{2}}{M}$, then the interval will have at least the desired asymptotic coverage level.
Theorem 8.
Let Assumption 2 hold. For any $\epsilon>0$, let $\widetilde{V}_{k}(\phi)$ denote the thresholded variance estimate given by
$\widetilde{V}_{k}(\phi)=\max\left\{\hat{V}_{k}(\phi)\ ,\ \frac{1}{z_{1-\frac{\alpha}{2}}^{2}\cdot\alpha/2}\cdot\frac{N^{2}}{M^{4/3-\epsilon}}\right\}$ (68)
and let $LB_{1-\alpha}$ denote
$LB_{1-\alpha}=\max\left\{\min_{\phi\in\{0,1\}^{N}}\hat{\tau}_{1}^{\textup{Haj}}(\phi)-z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{1}(\phi)}\,,\,\min_{\phi\in\{0,1\}^{N}}\hat{\tau}_{2}^{\textup{Haj}}(\phi)-z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{2}(\phi)}\right\}$ (69)
Then $\tau\geq LB_{1-\alpha}$ with probability converging to at least $1-\alpha$.
Remark
Theorems 1 - 3 can be seen to follow directly from Theorems 6 - 8, by substituting $M=N$ and applying Theorem 4.
Remark
The rates appearing in Theorems 6 - 8 may be more familiar or intuitive after normalizing the estimand by $N$, so that it equals $1-|{\mathcal{I}}|/N$, the fraction of units not in $\mathcal{I}$}, and then normalizing $\hat{\tau}_{k}^{\textup{Haj}}$ and $\hat{V}_{k}$ accordingly by $N$ and $N^{2}$ accordingly. For this choice of normalization, Theorem 6 implies $O_{P}(1/\sqrt{M})$ estimation error, as might be expected for an experiment whose randomization was constructed from $M$ independent draws. Similarly, Theorem 7 implies $o_{P}(1/M)$ error for the variance estimator after normalization by $N$, with Theorem 8 lower bounding the estimated variance by $o_{P}(1/M)$ as well.
Appendix D Proofs
Propositions 1 - 5 are proven in Section D.1. Theorems 4 and 5 are proven in Section D.2. Proofs of Theorems 6, 7, and 8 (which are generalized versions of Theorems 1, 2, and 3), including auxilliary lemmas, are respectively contained in Sections D.3, D.4, and D.5.
D.1 Proof of Propositions 1 - 5
Proof of Proposition 1.
It can be seen that
$\hat{\tau}_{1}$ $=N-\sum_{i\in{\mathcal{I}}}\left(\frac{X_{i}Y_{i}}{P(X_{i}=1)}+\frac{(1-X_{i})(1-Y_{i})}{P(X_{i}=0)}\right)$ (70)
$\geq N-\sum_{i=1}^{N}\left(\frac{X_{i}Y_{i}}{P(X_{i}=1)}+\frac{(1-X_{i})(1-Y_{i})}{P(X_{i}=0)}\right)$ (71)
$=N-\sum_{i=1}^{N}\left(\frac{X_{i}Y_{i}}{P(X_{i}=1)}-\frac{(1-X_{i})Y_{i}}{P(X_{i}=0)}+\frac{1-X_{i}}{P(X_{i}=0)}\right)$ (72)
$=\sum_{i=1}^{N}\left(\frac{1-X_{i}}{P(X_{i}=0)}-\frac{X_{i}}{P(X_{i}=1)}\right)Y_{i}+O_{P}(N^{1/2})$ (73)
where (70) uses the definition of $\hat{\tau}_{1}$ given in (5) and that
$P(H_{i}=1)=Y_{i}P(X_{i}=1)+(1-Y_{i})P(X_{i}=0)\qquad\forall\ i\in{\mathcal{I}}$
as $Y_{i}$ is constant if $i\in{\mathcal{I}}$; (71) holds as the terms in the summand are nonnegative; (72) is simple rearrangement of terms, and (73) uses (6) assumed by the lemma.
By essentially identical arguments, it holds that
$\hat{\tau}_{2}$ $=N-\sum_{i\in{\mathcal{I}}}\left(\frac{X_{i}(1-Y_{i})}{P(X_{i}=1)}+\frac{(1-X_{i})Y_{i}}{P(X_{i}=0)}\right)$
$\geq N-\sum_{i=1}^{N}\left(\frac{X_{i}(1-Y_{i})}{P(X_{i}=1)}+\frac{(1-X_{i})Y_{i}}{P(X_{i}=0)}\right)$
$\geq N-\sum_{i=1}^{N}\left(-\frac{X_{i}Y_{i}}{P(X_{i}=1)}+\frac{(1-X_{i})Y_{i}}{P(X_{i}=0)}+\frac{X_{i}}{P(X_{i}=1)}\right)$
$=\sum_{i=1}^{N}\left(\frac{X_{i}}{P(X_{i}=1)}-\frac{1-X_{i}}{P(X_{i}=0)}\right)Y_{i}+O_{P}(N^{1/2})$ (74)
Combining (73) and (74) yields
$\max(\hat{\tau}_{1},\hat{\tau}_{2})\geq\left|\sum_{i=1}^{N}\left(\frac{X_{i}}{P(X_{i}=1)}-\frac{1-X_{i}}{P(X_{i}=0)}\right)Y_{i}\right|,$
proving the result. ∎
Proof of Proposition 2.
Substitute $Z_{i}$ for $X_{i}$ in the proof of Theorem 1. ∎
Proof of Proposition 3.
By (28) and the conditions of the proposition, it holds for $a\in\{0,1\}$ that
$1\{\bar{Z}_{i}=a\}Y_{i}=\begin{cases}1\{\bar{Z}_{i}=a\}\tilde{f}_{i}(T_{i})&\textup{ if }i\in{\mathcal{I}}\\ 1\{\bar{Z}_{i}=a\}\bar{Z}_{i}&\textup{ if }i\notin{\mathcal{I}}\end{cases}$ (75)
where we have used that $\bar{Z}_{i}\in\{0,1\}$ implies $X\in R_{i}$ by (29). Plugging (75) into $\Delta^{\bar{Z}}$ yields
$\Delta^{\bar{Z}}$ $=\sum_{i\in{\mathcal{I}}}\left(\frac{1\{\bar{Z}_{i}=1\}}{P(\bar{Z}_{i}=1|T_{i})}-\frac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=0|T_{i})}\right)\tilde{f}_{i}(T_{i})+\sum_{i\notin{\mathcal{I}}}\left(\frac{1\{\bar{Z}_{i}=1\}}{P(\bar{Z}_{i}=1|T_{i})}-\frac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=0|T_{i})}\right)\bar{Z}_{i}$
It can be seen that the first summation (which is over $i\in{\mathcal{I}}$) has expectation zero, while the remaining summation has expectation $|\{i:i\notin{\mathcal{I}}\}|=N-|{\mathcal{I}}|$, proving the proposition. ∎
Proof of Proposition 4.
Let $\tilde{f}_{ki}$ for $i\in{\mathcal{I}}$ denote the mapping
$\tilde{f}_{ki}(T_{i})=\begin{cases}\tilde{f}_{i}(T_{i})&\textup{ if $k=1$}\\ 1-\tilde{f}_{i}(T_{i})&\textup{ if $k=2$}\end{cases}$ (76)
As $Y_{ki}=\tilde{f}_{ki}(T_{i})$ whenever $X\in R_{i}$ and $i\in{\mathcal{I}}$, it follows that
$P(X\in R_{i},Z_{i}=Y_{ki}|T_{i})$ $=P(X\in R_{i},Z_{i}=\tilde{f}_{ki}(T_{i})|T_{i})$ (77)
$=\sum_{a=0}^{1}1\{\tilde{f}_{ki}(T_{i})=a\}\ P(X\in R_{i},Z_{i}=a|T_{i},X\in R_{i})$
$=\tilde{f}_{ki}(T_{i})\cdot P(X\in R_{i},Z_{i}=1|T_{i})+(1-\tilde{f}_{ki}(T_{i}))\cdot P(X\in R_{i},Z_{i}=0|T_{i})$
$=Y_{ki}\cdot P(X\in R_{i},Z_{i}=1|T_{i})+(1-Y_{ki})\cdot P(X\in R_{i},Z_{i}=0|T_{i})$ (78)
where (77) holds because $i\in{\mathcal{I}}$, and (78) holds because $X\in R_{i}$ and $i\in{\mathcal{I}}$, so that we $Y_{ki}=\tilde{f}_{ki}(T_{i})$
Proof of Proposition 5.
Let $\tilde{f}_{ki}$ be given by (76). Since $i\in{\mathcal{I}}$ and $\bar{Z}_{i}=Y_{ki}$ (which implies $X\in R_{i}$), it holds that $Y_{ki}=\tilde{f}_{ki}(T_{i})$, and hence that
$\mathbb{E}[v_{i}|T_{i}]$ $=\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{\bar{Z}_{i}=Y_{ki}\}}{P(\bar{Z}_{i}=Y_{ki}|T_{i})}|T_{i}\right]$
$=P(X\in R_{i}|T_{i})\cdot\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{\bar{Z}_{i}=Y_{ki}\}}{P(\bar{Z}_{i}=Y_{ki}|T_{i})}|T_{i},X\in R_{i}\right]$ (79)
$=P(X\in R_{i}|T_{i})\cdot\sum_{a=0,1}1\{Y_{ki}=a\}\mathbb{E}\left[\frac{N}{\hat{N}_{a}}\cdot\frac{1\{\bar{Z}_{i}=a\}}{P(\bar{Z}_{i}=a|T_{i})}|T_{i},X\in R_{i}\right]$ (80)
$=\sum_{a=0,1}1\{Y_{ki}=a\}\cdot P(X\in R_{i}|T_{i})\mathbb{E}\left[\frac{N}{\hat{N}_{a}}\cdot\frac{1\{\bar{Z}_{i}=a\}}{P(\bar{Z}_{i}=a|T_{i})}|T_{i},X\in R_{i}\right]$ (81)
$=\sum_{a=0,1}1\{Y_{ki}=a\}\mathbb{E}\left[\frac{N}{\hat{N}_{a}}\cdot\frac{1\{\bar{Z}_{i}=a\}}{P(\bar{Z}_{i}=a|T_{i})}|T_{i}\right]$ (82)
where (79) holds since $1\{\bar{Z}_{i}=Y_{ki}\}=0$ if $X\notin R_{i}$; (80) holds since the $Y_{ki}$ term in the conditional expectation equals $\tilde{f}_{ki}(T_{i})$, and hence may be treated as a constant equal to the observed value of $Y_{ki}$; (81) is a simple rearrangement of terms; and (82) uses that $1\{\bar{Z}_{i}=a\}=0$ for $a\in\{0,1\}$ if $X\notin R_{i}$.
By similar reasoning, if $i,j\in{\mathcal{I}}$ and $1\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}=1$ (which implies that $X\in R_{i}\cap R_{j}$), it holds that $Y_{ki}=\tilde{f}_{ki}(T_{i})$ and $Y_{kj}=\tilde{f}_{kj}(T_{j})$, and hence that
$\mathbb{E}[v_{i}v_{i}|T_{i},T_{j}]$ $=\mathbb{E}\left[\frac{1\{\bar{Z}_{i}=Y_{ki}\}N}{P(\bar{Z}_{i}=Y_{ki}|T_{i})\hat{N}_{Y_{ki}}}\cdot\frac{1\{\bar{Z}_{i}=Y_{ki}\}N}{P(\bar{Z}_{i}=Y_{ki}|T_{i})\hat{N}_{Y_{ki}}}|T_{i},T_{j}\right]$
$=P(X\in R_{i}\cap R_{j})\mathbb{E}\left[\frac{1\{\bar{Z}_{i}=Y_{ki}\}N}{P(\bar{Z}_{i}=Y_{ki}|T_{i})\hat{N}_{Y_{ki}}}\cdot\frac{1\{\bar{Z}_{i}=Y_{ki}\}N}{P(\bar{Z}_{i}=Y_{ki}|T_{i})\hat{N}_{Y_{ki}}}|T_{i},T_{j},X\in R_{i}\cap R_{j}\right]$
$=P(X\in R_{i}\cap R_{j})\sum_{a=0}^{1}\sum_{b=0}^{1}1\{Y_{ki}=a,Y_{kj}=b\}\cdot$
$\hskip 28.45274pt\mathbb{E}\left[\frac{1\{\bar{Z}_{i}=a\}N}{P(\bar{Z}_{i}=a|T_{i})\hat{N}_{a}}\cdot\frac{1\{\bar{Z}_{i}=b\}N}{P(\bar{Z}_{i}=b|T_{i})\hat{N}_{b}}|T_{i},T_{j},X\in R_{i}\cap R_{j}\right]$
$=\sum_{a=0}^{1}\sum_{b=0}^{1}1\{Y_{ki}=a,Y_{kj}=b\}\cdot$
$\hskip 28.45274ptP(X\in R_{i}\cap R_{j})\mathbb{E}\left[\frac{1\{\bar{Z}_{i}=a\}N}{P(\bar{Z}_{i}=a|T_{i})\hat{N}_{a}}\cdot\frac{1\{\bar{Z}_{i}=b\}N}{P(\bar{Z}_{i}=b|T_{i})\hat{N}_{b}}|T_{i},T_{j},X\in R_{i}\cap R_{j}\right]$
$=\sum_{a=0}^{1}\sum_{b=0}^{1}1\{Y_{ki}=a,Y_{kj}=b\}\mathbb{E}\left[\frac{1\{\bar{Z}_{i}=a\}N}{P(\bar{Z}_{i}=a|T_{i})\hat{N}_{a}}\cdot\frac{1\{\bar{Z}_{i}=b\}N}{P(\bar{Z}_{i}=b|T_{i})\hat{N}_{b}}|T_{i},T_{j}\right]$
and
$P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})$ $=P(X\in R_{i}\cap R_{j}|T_{i},T_{j})P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j},X\in R_{j}\cap R_{j})$
$=P(X\in R_{i}\cap R_{j}|T_{i},T_{j})\cdot\sum_{a=0}^{1}\sum_{b=0}^{1}1\{Y_{ki}=a,Y_{kj}=b\}\cdot$
$\hskip 28.45274ptP(\bar{Z}_{i}=a,\bar{Z}_{j}=b|T_{i},T_{j},X\in R_{j}\cap R_{j})$
$=\sum_{a=0}^{1}\sum_{b=0}^{1}1\{Y_{ki}=a,Y_{kj}=b\}\cdot$
$\hskip 28.45274ptP(X\in R_{i}\cap R_{j}|T_{i},T_{j})P(\bar{Z}_{i}=a,\bar{Z}_{j}=b|T_{i},T_{j},X\in R_{j}\cap R_{j})$
$=\sum_{a=0}^{1}\sum_{b=0}^{1}1\{Y_{ki}=a,Y_{kj}=b\}\cdot P(\bar{Z}_{i}=a,\bar{Z}_{j}=b|T_{i},T_{j})$
D.2 Proof of Theorems 4 and 5
Proof of Theorem 4.
Given design neighborhoods $\Gamma_{i}$ satisfying (37) and (38), for $i\in[N]$ let $G_{i}=\{i\}$ and $\eta_{i}=\Gamma_{i}$. The conditions (50) - (56) of Assumption 2 can be verified as follows:
  1. 1.
    (50) and (53) follow from the definitions of $\{G_{i}\}$ and $\{\eta_{i}\}$
  2. 2.
    (51) and (52) follow because $\eta_{i}=\Gamma_{i}$, which satisfies (37) and (38).
  3. 3.
    (54) holds because it is assumed as (40)
  4. 4.
    To show (55) and (56), we observe that $|\Gamma_{i}\cup\Gamma_{j}|\leq 2d$, and hence by (39) and independence of the treatment assignments it holds that
    $P(X_{\Gamma_{i}\cup\Gamma_{j}}=x)\geq\min(p,1-p)^{2d}\qquad\forall x\in\{0,1\}^{|\Gamma_{i}\cup\Gamma_{j}|}$
    As a result, for any $(t_{i},t_{j})\in\operatorname{supp}(T_{i},T_{j})$ it holds that
    $P(T_{i}=t_{i},T_{j}=t_{j})\geq\min(p,1-p)^{2d}$ (83)
    where we have used that $(T_{i},T_{j})$ depend on $X$ only through $X_{\Gamma_{i}\cup\Gamma_{j}}$. Similarly, for any $(a,b,t_{i},t_{j})\in\operatorname{supp}(\bar{Z}_{i},\bar{Z}_{j},T_{i},T_{j})$ it also holds that
    $P(\bar{Z}_{i}=a,\bar{Z}_{j}=b,T_{i}=t_{i},T_{j}=t_{j})\geq\min(p,1-p)^{2d},$ (84)
    where we have used that $(\bar{Z}_{i},\bar{Z}_{j},T_{i},T_{j})$ depends on $X$ only through $X_{\Gamma_{i}\cup\Gamma_{j}}$. By (83) and (84), it follows that (55) and (56) hold for $\rho=\min(p,1-p)^{2d}$.
We prove Theorem 5 in two parts. First we prove the result under the original conditions A1 - A5. We then show that the result continues to hold when Conditions A3 and A4 are replaced by Condition A6, as stated by the theorem.
Proof of Theorem 5, using Conditions A1 - A5.
The conditions (50) - (56) of Assumption 2 can be verified as follows:
  1. 1.
    (50), (51), (52), and (54) hold because they are explicitly assumed by the conditions of the theorem.
  2. 2.
    (53) holds because $\bar{Z}_{i}$ and $T_{i}$ depend only on $X_{i}$, $X_{\mathcal{N}_{i}}$, and $\theta_{\eta_{g(i)}}$, all of which depend only on the treatment of unit $i$’s own and neighboring groups.
  3. 3.
    To show (55) and (56), we will show that for $\rho$ given by
    $\rho=p^{2d}\left(\frac{p}{1+(1-p)2d}\right)^{2d}$
    it holds that
    $P(T_{i}=t_{i},T_{j}=t_{j})$ $\geq\rho\qquad\forall(t_{i},t_{j})\in\operatorname{supp}(T_{i},T_{j})$ (85)
    which implies (55), and it also holds that
    $P(\bar{Z}_{i}=a,\bar{Z}_{j}=b,T_{i}=t_{i},T_{j}=t_{j})$ $\geq\rho\qquad\forall(a,b,t_{i},t_{j})\in\operatorname{supp}(\bar{Z}_{i},\bar{Z}_{j},T_{i},T_{j})$ (86)
    which implies (56).
To complete the proof, we now show (86), which also implies (85). Let the sets $S\subset[N]$ and $Q\subset[M]$ be given by
$S=\mathcal{N}_{i}\cup\mathcal{N}_{j}\qquad\textup{and}\qquad Q=\eta_{g(i)}\cup\eta_{g(j)}$
Because $(\bar{Z}_{i},\bar{Z}_{j},T_{i},T_{j})$ depend on $X$ and $\theta$ only through $X_{S}$ and $\theta_{Q}$, to show (86) it suffices to show that
$P(X_{S}=x,\theta_{Q}=\xi)$ $\geq\rho\qquad\forall(x,\xi)\in\operatorname{supp}(X_{S},\theta_{Q})$ (87)
To show (87), let $X_{S,i}$ denote the $i$th entry of $X_{S}$, let $X_{S,1:i-1}$ denote the first (i1)(i-1) entries of $X_{S}$, let $\bar{p}=\min(p,1-p)$, and let $(a_{i},b_{i},k_{i},\tilde{k}_{i})$ denote terms to be defined shortly. It then holds for any $(x,\xi)\in\operatorname{supp}(X_{S},\theta_{Q})$ that
$P(X_{S}=x,\theta_{Q}=\xi)$ $\geq p^{2d}P(X_{S}=x|\theta_{Q}=\xi)$ (88)
$=p^{2d}\prod_{i=1}^{|S|}P(X_{S,i}=x_{i}|\theta_{Q}=\xi,X_{S,1:i-1}=x_{1:i-1})$ (89)
$=p^{2d}\prod_{i=1}^{|S|}\frac{a_{i}-k_{i}}{b_{i}-\tilde{k}_{i}}$ (90)
$\geq p^{2d}\left(\frac{\bar{p}}{1+(1-\bar{p})2d}\right)^{2d}$ (91)
by the following steps:
  1. 1.
    (88) holds because $\theta_{1},\ldots,\theta_{M}$ are indpendent and satisfy (58)
  2. 2.
    (89) writes $P(X_{S}|\theta_{Q})$ as a product of conditional probabilities
  3. 3.
    (90) holds because the entries of $X_{S}$ conditional on $\theta_{Q}$ are generated by sampling without replacement from each group, with $(a_{i},b_{i},k_{i},\tilde{k}_{i})$ defined as follows. Let $\{\ell_{1},\ldots,\ell_{|S|}\}\subset[N]$ enumerate the units in $S$, so that $X_{S}=(X_{\ell_{1}},\ldots,X_{\ell_{|S|}})$ and $X_{S,i}=X_{\ell_{i}}$. Then it holds that
    $P(X_{S,i}=x_{i}|\theta_{Q}=\xi,X_{S,1:i-1}=x_{1:i-1})=\frac{a_{i}-k_{i}}{b_{i}-\tilde{k}_{i}}$
    for $(a_{i},b_{i},k_{i},\tilde{k}_{i})$ given by
    $a_{i}$ $=\begin{cases}\theta_{g(\ell_{i})}&\textup{if }x_{i}=1\\ |G_{g(\ell_{i})}|-\theta_{g}(\ell_{i})&\textup{if }x_{i}=0,\end{cases}$ $b_{i}$ $=|G_{g(\ell_{i})}|$
    $k_{i}$ $=\sum_{j=1}^{i-1}1\{x_{j}=x_{i},g(\ell_{j})=g(\ell_{i})\},$ $\tilde{k}_{i}$ $=\sum_{j=1}^{i-1}1\{g(\ell_{j})=g(\ell_{i})\}$ (92)
    because the conditional probability of unit $\ell_{i}$ receiving treatment is equal to the probability of drawing a colored ball from an urn with $|G_{g(\ell_{i})}|$ balls, of which $\theta_{g(\ell_{i})}$ are colored, conditional on previously drawing without replacement a certain number of colored and non-colored balls corresponding to the treatment assignment of the units in $\{\ell_{1},\ldots,\ell_{i-1}\}$ who are in the same group as $\ell_{i}$.
  4. 4.
    To show (91) holds we observe that
    $\frac{a_{i}-k_{i}}{b_{i}-\tilde{k}_{i}}=1\qquad\textup{ if }\theta_{g(\ell_{i})}=0\textup{ or }\theta_{g(\ell_{i})}=|G_{g(\ell_{i})}|$ (93)
    which holds because the treatment of unit $\ell_{i}$ can only take one value when all units in $\ell_{i}$’s group receive the same treatment, and
    $\frac{a_{i}-k_{i}}{b_{i}-\tilde{k}_{i}}\geq\frac{\bar{p}}{1+(1-\bar{p})d}\qquad\textup{ if }\theta_{g(\ell_{i})}\neq 0\textup{ and }\theta_{g(\ell_{i})}\neq|G_{g(\ell_{i})}|$ (94)
    which holds by the following steps for $\theta_{g(\ell_{i})}\notin\{0,|G_{g(\ell_{i})}|\}$:
    $\frac{a_{i}-k_{i}}{b_{i}-\tilde{k}_{i}}$ $\geq\frac{a_{i}-k_{i}}{b_{i}-k_{i}}$ (95)
    $\geq\frac{a_{i}-k_{i}}{\frac{a_{i}}{\bar{p}}-k_{i}}$ (96)
    $=\frac{\bar{p}(a_{i}-k_{i})}{(a_{i}-k_{i})+(1-\bar{p})k_{i}}$ (97)
    $\geq\frac{\bar{p}}{1+(1-\bar{p})2d}$ (98)
    where (95) because $\tilde{k}_{i}\geq k_{i}$ by their definitions in (92); (96) holds because $b_{i}=|G_{g(\ell_{i})}|\leq a_{i}/\bar{p}$ by (57); (97) follows from algebraic manipulation; and (98) holds because $k_{i}\leq|S|\leq 2d$, and because the expression
    $\frac{\bar{p}(a_{i}-k_{i})}{(a_{i}-k_{i})+(1-\bar{p})2d}$
    is increasing in $(a_{i}-k_{i})$, where $(a_{i}-k_{i})\geq 1$ since $x$ is in the support of $X_{S}$ conditional on $\theta_{Q}$. Plugging into (90) the bounds for $\frac{a_{i}-k_{i}}{b_{i}-\tilde{k}_{i}}$ given by (93) and (94) and bounding $|S|\leq 2d$ establishes (91).
Proof of Theorem 5, using Condition A6.
To verify the conditions (50) - (56) of Assumption 2 when Conditions A3 and A4 are replaced by Condition A6, we observe that
  1. 1.
    As before, (50), (51), (52), and (54) hold because they are explicitly assumed by the conditions of the theorem.
  2. 2.
    (53) holds because $\bar{Z}_{i}$ and $T_{i}$ depend only on $\theta_{\eta_{g(i)}}$, the number of treated units in $i$’s own and neighboring groups.
  3. 3.
    (55) holds because $T_{i}$ and $T_{j}$ are independent for $i\neq j$, under the conditions of the theorem.
  4. 4.
    To show (56), we observe that $(\bar{Z}_{i},\bar{Z}_{j},T_{i},T_{j})$ depend on $X$ only through the group-level variables $\{\theta_{m}:m\in\eta_{g(i)}\cup\eta_{g(j)}\}$. Since $\{\theta_{m}\}_{m=1}^{M}$ are independent variables which satisfy the support condition (58), and $|\eta_{g(i)}\cup\eta_{g(j)}|\leq 2d$ by (51), it follows for any $(a,b,t_{i},t_{j})$ in the support of $(\bar{Z}_{i},\bar{Z}_{j},T_{i},T_{j})$ that
    $P(\bar{Z}_{i}=a,\bar{Z}_{j}=b,T_{i}=t_{i},T_{j}=T_{j})\geq p^{2d}$ (99)
    which implies (56) for $\rho=p^{2d}$.
D.3 Proof of Theorem 6
Preliminaries
The Azuma-Hoeffding (or McDiarmid’s) inequality (Boucheron et al., , 2013, Thm 6.2) states that given independent variables $U=(U_{1},\ldots,U_{N})$, and a function $f$ satisfying the bounded difference property
$|f(U)-f(U^{\prime})|\leq c_{i}\text{ if }U_{j}=U_{j}^{\prime}\text{ for all }j\neq i,\qquad i\in[N],$ (100)
it holds that
$\mathbb{P}\left(|f(U)-\mathbb{E}f|>\epsilon\right)\leq 2\exp\left(-2\epsilon^{2}/\sum_{i}c_{i}^{2}\right).$ (101)
Lemma 1 applies McDiarmid’s inequality to show concentration of the terms appearing in $\hat{\tau}_{k}^{\textup{Haj}}$ for $k=1,2$.
Lemma 1.
Let Assumption 2 hold, and let $\phi^{*}\in\{0,1\}^{N}$ encode $\mathcal{I}$} by
$\phi_{i}^{*}=1\{i\in{\mathcal{I}}\},\qquad i\in[N]$ (102)
For $i\in[N]$ and $\bar{Z}_{i}$ given by (29), let $u_{i}$ denote the vector
$u_{i}=\left(\frac{1\{\bar{Z}_{i}=1\}}{P(\bar{Z}_{i}=1|T_{i})}\ ,\ \frac{1\{\bar{Z}_{i}=1\}}{P(\bar{Z}_{i}=1|T_{i})}Y_{i}\phi_{i}^{*}\ ,\ \frac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=0|T_{i})}\ ,\ \frac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=0|T_{i})}(1-Y_{i})\phi_{i}^{*}\right)$
and let $\bar{u}=\frac{1}{N}\sum_{i=1}^{N}u_{i}$. Then $\|\bar{u}-\mathbb{E}\bar{u}\|_{\infty}=O_{P}\left(\frac{N}{\sqrt{M}}\right)$.
Proof.
Let $U_{m}$ for $m\in[M]$ be given by
$U_{m}=\frac{1}{N}\sum_{i\in G_{m}}u_{i}$
[…]
$P(|\bar{u}_{\ell}-\mathbb{E}\bar{u}_{\ell}|\geq\epsilon)\leq\exp\left(-\frac{2\epsilon^{2}}{M\left(\frac{dC}{M\rho}\right)^{2}}\right)$
and hence $\bar{u}_{\ell}-\mathbb{E}\bar{u}_{\ell}=O_{P}(M^{-1/2})$. Since $\bar{u}$ is 4-dimensional, it follows that $\|\bar{u}-\mathbb{E}\bar{u}\|_{\infty}=O_{P}(M^{-1/2})$ as well. ∎
Proof of Theorem 6
The claims of Theorem 6, namely (64) and (65), are proven separately as follows:
Proof of Theorem 6, Eq. (64).
It holds that
$\hat{\tau}_{k}^{\textup{Haj}}$ $=N-\left(\sum_{i\in{\mathcal{I}}}\frac{N}{\hat{N}_{1}}\cdot\frac{1\{\bar{Z}_{i}=1\}}{P(\bar{Z}_{i}=1|T_{i})}\cdot Y_{ki}+\sum_{i\in{\mathcal{I}}}\frac{N}{\hat{N}_{0}}\cdot\frac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=1|T_{i})}\cdot(1-Y_{ki})\right)$
$\geq N-\left(\sum_{i=1}^{N}\frac{N}{\hat{N}_{1}}\cdot\frac{1\{\bar{Z}_{i}=1\}}{P(\bar{Z}_{i}=1|T_{i})}\cdot Y_{ki}+\sum_{i=1}^{N}\frac{N}{\hat{N}_{0}}\cdot\frac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=1|T_{i})}\cdot(1-Y_{ki})\right)$ (103)
$=N-\left(\sum_{i=1}^{N}\frac{N}{\hat{N}_{1}}\cdot\frac{1\{\bar{Z}_{i}=1\}}{P(\bar{Z}_{i}=1|T_{i})}\cdot Y_{ki}-\sum_{i=1}^{N}\frac{N}{\hat{N}_{0}}\cdot\frac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=1|T_{i})}\cdot Y_{ki}\right)-N$ (104)
$=\begin{cases}-\Delta^{{\textup{Haj}}}&\textup{ if }k=1\\ \Delta^{{\textup{Haj}}}&\textup{ if }k=2\end{cases}$ (105)
where (103) holds because the terms in the summand are non-negative; (104) uses
$\sum_{i=1}^{N}\frac{N}{\hat{N}_{a}}\cdot\frac{1\{\bar{Z}_{i}=a\}}{P(\bar{Z}_{i}=a|T_{i})}=N,\qquad\forall\ a\in\{0,1\},$
and (105) uses the definition of $\Delta^{\textup{Haj}}$ given by (33). It follows that
$\max(\hat{\tau}_{1}^{\textup{Haj}},\hat{\tau}_{2}^{\textup{Haj}})$ $\geq\max(\Delta^{\textup{Haj}},-\Delta^{\textup{Haj}})$
$=|\Delta^{\textup{Haj}}|$
proving (64).
Proof of Theorem 6, Eq. (65).
Proof of (65) follows by Taylor approximation and is more or less standard. Given $k\in[2]$, let $a,b\in\mathbb{R}^{2}$ be given by
$a$ $=\left[\begin{array}[]{l}\dfrac{1}{N}\displaystyle\sum_{i=1}^{N}\dfrac{1\{\bar{Z}_{i}=1\}}{P(\bar{Z}_{i}=1|T_{i})}\cdot Y_{ki}\phi_{i}^{*}\\ \\ \dfrac{1}{N}\displaystyle\sum_{i=1}^{N}\dfrac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=0|T_{i})}\cdot(1-Y_{ki})\phi_{i}^{*}\end{array}\right]$  $\qquad\textup{ and }\qquad b$ $=\left[\begin{array}[]{c}\dfrac{1}{N}\displaystyle\sum_{i=1}^{N}\dfrac{1\{\bar{Z}_{i}=1\}}{P(\bar{Z}_{i}=1|T_{i})}\\ \\ \dfrac{1}{N}\displaystyle\sum_{i=1}^{N}\dfrac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=0|T_{i})}\end{array}\right]$ (112)
To evaluate $\mathbb{E}a$, we observe that for $z\in\{0,1\}$, $i\in{\mathcal{I}}$, and $\tilde{f}_{ki}(T_{i})$ given by (76), it holds that
$\mathbb{E}\left[\frac{1\{\bar{Z}_{i}=z\}}{P(\bar{Z}_{i}=z|T_{i})}Y_{ki}\right]$ $=\mathbb{E}\left[\mathbb{E}\left\{\frac{1\{\bar{Z}_{i}=z\}}{P(\bar{Z}_{i}=z|T_{i})}Y_{ki}|T_{i}\right\}\right]$
$=\mathbb{E}\left[\mathbb{E}\left\{Y_{ki}|T_{i},\bar{Z}_{i}=z\right\}\right]$
$=\mathbb{E}\left[\mathbb{E}\left\{\tilde{f}_{ki}(T_{i})|T_{i},\bar{Z}_{i}=z\right\}\right]$ (113)
$=\mathbb{E}\left\{\mathbb{E}\left[\tilde{f}_{ki}(T_{i})|T_{i}\right]\right\}$
$=\mathbb{E}[\tilde{f}_{ki}(T_{i})]$
where (113) hold since $Y_{ki}=\tilde{f}_{ki}(T_{i})$ if $i\in{\mathcal{I}}$ and $X\in R_{i}$, which is implied by $\bar{Z}_{i}\in\{0,1\}$. It follows that $\mathbb{E}a$ and $\mathbb{E}b$ equal
$\mathbb{E}a$ $=\left[\begin{array}[]{l}\dfrac{1}{N}\displaystyle\sum_{i\in{\mathcal{I}}}\mathbb{E}[\tilde{f}_{ki}(T_{i})]\\ \dfrac{1}{N}\displaystyle\sum_{i\in{\mathcal{I}}}\mathbb{E}[1-\tilde{f}_{ki}(T_{i})]\end{array}\right],$ $\mathbb{E}b=\left[\begin{array}[]{l}1\\ 1\end{array}\right]$ (118)
and that $\hat{\tau}_{k}^{\textup{Haj}}$ and $\tau$ are given by
$\hat{\tau}_{k}^{\textup{Haj}}$ $=N\left(\frac{a_{1}}{b_{1}}+\frac{a_{2}}{b_{2}}\right),$ $\tau$ $=N\left(\frac{\mathbb{E}a_{1}}{\mathbb{E}b_{1}}+\frac{\mathbb{E}a_{2}}{\mathbb{E}b_{2}}\right)$
and the estimation error may be written as
$\frac{1}{N}\left(\hat{\tau}_{k}^{\textup{Haj}}-\tau\right)$ $=\frac{\epsilon_{1}}{b_{1}}+\frac{\epsilon_{2}}{b_{2}}$
where $\epsilon\in\mathbb{R}^{2}$ denotes
$\epsilon=\left[\begin{array}[]{l}a_{1}-b_{1}\dfrac{\mathbb{E}a_{1}}{\mathbb{E}b_{1}}\\ \\ a_{2}-b_{2}\dfrac{\mathbb{E}a_{2}}{\mathbb{E}b_{2}}\end{array}\right]$ (122)
which can be seen to satisfy $\mathbb{E}\epsilon=0$.
To show (65), we define the mapping $h$ by
$h(\epsilon,b)=\frac{\epsilon_{1}}{b_{1}}+\frac{\epsilon_{2}}{b_{2}}$
so that $N^{-1}(\hat{\tau}_{k}^{\textup{Haj}}-\tau)=h(\epsilon,b)$. Letting $\xi=(\epsilon,b)$, by mean value theorem it holds for some $\tilde{\xi}$ in the convex hull of $\xi$ and $\mathbb{E}\xi$ that
$h(\xi)-h(\mathbb{E}\xi)$ $=\nabla h(\tilde{\xi})^{T}(\xi-\mathbb{E}\xi)$
$=\nabla h(\mathbb{E}\xi)^{T}(\xi-\mathbb{E}\xi)+\left(\nabla h(\tilde{\xi})-\nabla h(\mathbb{E}\xi)\right)^{T}(\xi-\mathbb{E}\xi)$ (123)
where the derivative $\nabla h(\xi)$ is given by
$\nabla h(\xi)=\left[\dfrac{1}{b_{1}}\ \ \ \dfrac{1}{b_{2}}\ \ -\frac{\epsilon_{1}}{b_{1}^{2}}\ \ -\frac{\epsilon_{2}}{b_{2}^{2}}\right]$ (124)
As $\xi$ is an affine transformation of $\bar{u}$ defined in Lemma 1, it follows from that Lemma that
$\|\xi-\mathbb{E}\xi\|_{\infty}=O_{P}(M^{-1/2})$ (125)
from which it follows that
$\nabla h(\mathbb{E}\xi)^{T}(\xi-\mathbb{E}\xi)=O_{P}(M^{-1/2})$ (126)
where we have used $\nabla h(\mathbb{E}\xi)=\left(1,1,0,0\right)$.
As $\tilde{\xi}$ is in the convex hull of $\xi$ and $\mathbb{E}\xi$, (125) implies that $\|\tilde{\xi}-\mathbb{E}\xi\|=O_{P}(M^{-1/2})$. By continuity of $\nabla h$ at $\mathbb{E}\xi$, it follows that $\|\nabla h(\tilde{\xi})-\nabla h(\mathbb{E}\xi)\|=O_{P}(M^{-1/2})$ as well, and hence that
$\left(\nabla h(\tilde{\xi})-\nabla h(\mathbb{E}\xi)\right)^{T}(\xi-\mathbb{E}\xi)=O_{P}(M^{-1})$ (127)
Substituting (126) and (127) into (123) proves that $N^{-1}(\hat{\tau}_{k}-\tau)=O_{P}(M^{-1/2})$, which is equivalent to the claim of (65) made by the lemma. ∎
D.4 Proof of Theorem 7
In our proof of Theorem 7, we will use $C_{ij}(T_{i},T_{j})$ to abbreviate the covariance estimates appearing in $\hat{V}_{k}$. It is given by
$C_{ij}(T_{i},T_{j})=\mathbb{E}[v_{i}v_{j}|T_{i},T_{j}]-\mathbb{E}[v_{i}|T_{i}]\mathbb{E}[v_{j}|T_{j}]\frac{P(T_{i})P(T_{j})}{P(T_{i},T_{j})}$ (128)
To prove the theorem, we will use Lemmas 2 - 7, whose proofs will be given following the proof of Theorem 7. Lemma 2 bounds the values of $\{C_{ij}\}$.
Lemma 2.
Let Assumption 2 hold and let $C_{ij}$ be given by (128). Let $\mathcal{J}_{0}$ and $\mathcal{J}_{1}$ denote the sets
$\mathcal{J}_{0}$ $=\{i,j\in{\mathcal{I}}:|\Gamma_{i}\cap\Gamma_{j}|=0\}$
$\mathcal{J}_{1}$ $=\{i,j\in{\mathcal{I}}:|\Gamma_{i}\cap\Gamma_{j}|\neq 0\}$
It holds w.p. 1 that
$\max_{i,j\in\mathcal{J}_{0}}C_{ij}(T_{i},T_{j})$ $\leq O\left(\frac{\log M}{M}\right)$ (129)
$\max_{i,j\in\mathcal{J}_{1}}C_{ij}(T_{i},T_{j})$ $\leq O(1)$ (130)
Lemmas 3 and 4 are used to prove (67) of Theorem 7.
Lemma 3.
Let $C_{ij}$ be given by (128). If $i,j\in{\mathcal{I}}$ and $P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})=0$, then $C_{ij}(T_{i},T_{j})\leq 0$
Lemma 4.
It holds that $\mathbb{E}[C_{ij}(T_{i},T_{j})]\geq{\operatorname{Cov}}(v_{i},v_{j})$.
Lemmas 5 and 6 bound terms appearing in our variance estimators, and will be used to prove Lemma 2.
Lemma 5.
Let Assumption 2 hold, and let $v_{i}$ for $i\in[N]$ denote
$v_{i}=\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{\bar{Z}_{i}=Y_{ki}\}}{P(\bar{Z}_{i}=Y_{ki}|T_{i})},$
where $\hat{N}_{a}=\sum_{j=1}^{N}1\{\bar{Z}_{i}=a\}/P(\bar{Z}_{i}=a|T_{i})$ for $a\in\{0,1\}$. Then
$\mathbb{E}\left[v_{i}|T_{i}\right]$ $=\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|T_{i},\bar{Z}_{i}=Y_{ki}\right].$ (131)
If $i,j\in{\mathcal{I}}$ and $|\Gamma_{i}\cap\Gamma_{j}|=0$, then
$\mathbb{E}\left[v_{i}v_{j}|T_{i},T_{j}\right]$ $=\mathbb{E}\left[\frac{N^{2}}{\hat{N}_{Y_{ki}}\hat{N}_{Y_{kj}}}|T_{i},T_{j},\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\right].$ (132)
If $i,j\in{\mathcal{I}}$ and $|\Gamma_{i}\cap\Gamma_{j}|\neq 0$, then
$\mathbb{E}\left[v_{i},v_{j}|T_{i}T_{j}\right]$ $\leq\frac{1}{\rho^{2}}\mathbb{E}\left[\frac{N^{2}}{\hat{N}_{Y_{ki}}\hat{N}_{Y_{kj}}}|T_{i},T_{j},\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\right].$ (133)
Lemma 6.
Let Assumption 2 hold. Then
$\max_{i\in{\mathcal{I}}}\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|T_{i},\bar{Z}_{i}=Y_{ki}\right]$ $=1+O\left(\frac{\log M}{M}\right)$ (134)
$\max_{i,j\in{\mathcal{I}}}\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|T_{i},\bar{Z}_{i}=Y_{ki},T_{j},\bar{Z}_{j}=Y_{kj}\right]$ $=1+O\left(\frac{\log M}{M}\right)$ (135)
$\max_{i,j\in{\mathcal{I}}}{\operatorname{Var}\,}\left[\frac{N}{\hat{N}_{Y_{ki}}}|T_{i},\bar{Z}_{i}=Y_{ki},T_{j},\bar{Z}_{j}=Y_{kj}\right]$ $=O\left(\frac{1}{M}\right)$ (136)
Lemma 7 bounds the extent to which the design neighborhoods $\{\Gamma_{i}\}_{i=1}^{N}$ may overlap, and the number of design neighborhoods to which a group $G_{m}$ may belong.
Lemma 7.
Let Assumption 2 hold. It holds that
$\max_{i\in[N]}\sum_{j}1\{|\Gamma_{i}\cup\Gamma_{j}|\neq 0\}$ $\leq\frac{d^{2}CN}{M}$ (137)
$\max_{m\in[M]}\sum_{i}1\{G_{m}\subseteq\Gamma_{i}\}$ $\leq\frac{dCN}{M}$ (138)
Proof of Theorem 7
The claims of Theorem 7, namely (66), (67), and the conditions under which $\hat{V}_{k}$ is unbiased, are proven separately as follows.
Proof of Theorem 7, Eq. (66).
For $C_{ij}$ given by (128), let $Q\in\mathbb{R}^{N\times N}$ denote the matrix given by
$Q_{ij}=\begin{cases}C_{ij}\cdot\dfrac{1\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}}{P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})}&\textup{ if $i,j\in{\mathcal{I}}$}\\ 0&\textup{ else }\end{cases}$ (139)
so that $\hat{V}_{k}=\sum_{ij}Q_{ij}$.
To bound $|Q_{ij}|$, we may apply Lemma 2 and (55) to bound $|C_{ij}|$ and then (56) to bound the term $1/P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})$ appearing in (139), yielding for some $C_{1}>0$ that
$|Q_{ij}|\leq\begin{cases}\displaystyle{\frac{C_{1}\log M}{M}}&i,j\in\mathcal{J}_{0}\\ \displaystyle{C_{1}}&i,j\in\mathcal{J}_{1}.\end{cases}$ (140)
By (137) it holds that
$\max_{i\in[N]}|\{j:(i,j)\in\mathcal{J}_{1}\}|\leq\frac{d^{2}CN}{M}.$ (141)
Changing $X_{G_{m}}$ changes at most $\frac{dCN}{M}$ rows and columns of $Q$, by (138). Combining (140) and (141) implies that each row or column of $Q$ has entries summing to at most $C_{2}\frac{N}{M}(\log M+1)$ for some $C_{2}>0$. It thus follows for some $C_{3}>0$ that for all $X,X^{\prime}$ differing only in $X_{G_{m}}$ (for any $m\in[M]$) that
$|\hat{V}_{k}(X)-\hat{V}_{k}(X^{\prime})|\leq C_{3}\frac{N^{2}}{M^{2}}(\log M+1)$
and it follows by McDiarmid’s inequality that
$P(|\hat{V}_{k}-\mathbb{E}\hat{V}_{k}|\geq\epsilon)\leq 2\exp\left(\frac{-2\epsilon^{2}}{M(CC_{1})^{2}N^{4}M^{-4}(\log M+1)^{2}}\right)$
which implies that $|V-\mathbb{E}\hat{V}_{k}|=O_{P}\left(\frac{N^{2}\log M}{M^{3/2}}\right)$. ∎
Proof of Theorem 7, Eq. (67).
Recall that $C_{ij}$ is given by (128). To show that $\hat{V}_{k}$ is not anti-conservative, we observe that
$\mathbb{E}\hat{V}_{k}$ $=\sum_{i,j\in{\mathcal{I}}}\ \mathbb{E}\left[\mathbb{E}\left[C_{ij}(T_{i},T_{j})\cdot\frac{1\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}}{P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})}\ \bigg|T_{i},T_{j}\right]\right]$
$=\sum_{i,j\in{\mathcal{I}}}\ \mathbb{E}\left[C_{ij}(T_{i},T_{j})\cdot\mathbb{E}\left[\frac{1\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}}{P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})}\ \bigg|T_{i},T_{j}\right]\right]$
$=\sum_{i,j\in{\mathcal{I}}}\ \sum_{t_{i},t_{j}}P(T_{i}=t_{i},T_{j}=t_{j})\cdot C_{ij}(t_{i},t_{j})\cdot$
$\hskip 28.45274pt1\{P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i}=t_{i},T_{j}=t_{j})\neq 0\}$
$\geq\sum_{i,j\in{\mathcal{I}}}\ \sum_{t_{i},t_{j}}P(T_{i}=t_{i},T_{j}=t_{j})C_{ij}(t_{i},t_{j})$ (142)
$\geq\sum_{i,j\in{\mathcal{I}}}\ {\operatorname{Cov}}(v_{j},v_{j})$ (143)
where (142) holds by Lemma 3, and (143) holds by Lemma 4. It can be seen that (143) equals the claim of (67) made by Theorem 7. ∎
Proof of Theorem 7, unbiasedness claim.
To verify the conditions for unbiasedness given by Theorem 7, we observe that if $T_{i}$ is the constant mapping for all units, then $\hat{V}_{k}$ simplifies to
$\hat{V}_{k}=\sum_{i,j\in{\mathcal{I}}}{{\operatorname{Cov}}}\left(v_{i},v_{j}\right)\cdot\frac{1\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}}{P\left(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\right)}$
where $Y_{ki}$ is constant for all $i\in[N]$ such that $i\in{\mathcal{I}}$. As this is an importance sampled estimator of covariance terms appearing in ${\operatorname{Var}\,}\tau_{k}^{{\textup{Haj}}}$, it is unbiased provided that the sampling probabilities $P\left(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\right)$ (which equals $P(\bar{Z}_{i}=Y_{ki})$ for $i=j$) are non-zero for each $i,j\in{\mathcal{I}}$.
Proof of Lemmas 2 - 6
Proof of Lemma 2.
For $i,j$ satisfying $i,j\in{\mathcal{I}}$, let $F_{\{i\}}$ and $F_{\{i,j\}}$ denote sigma-fields such that
$\mathbb{E}[\ \cdot\ |F_{\{i\}}]$ $=\mathbb{E}[\ \cdot\ |T_{i},\bar{Z}_{i}=Y_{ki}]$
$\mathbb{E}[\ \cdot\ |F_{\{i,j\}}]$ $=\mathbb{E}[\ \cdot\ |T_{i},T_{j},\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}]$
Recall $C_{ij}$ is given by (128). To show (129), we observe that for $i,j\in\mathcal{J}_{0}$ that
$C_{ij}(T_{i},T_{j})$ $=\mathbb{E}[v_{i}v_{j}|T_{i},T_{j}]-\mathbb{E}[v_{i}|T_{i}]\mathbb{E}[v_{j}|T_{j}]\frac{P(T_{i})P(T_{j})}{P(T_{i},T_{j})}$
$=\mathbb{E}\left[\frac{N^{2}}{\hat{N}_{Y_{ki}}\hat{N}_{Y_{kj}}}|F_{\{i,j\}}\right]-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i\}}\right]\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{j\}}\right]$ (144)
$=\mathbb{E}\left[\frac{N^{2}}{\hat{N}_{Y_{ki}}\hat{N}_{Y_{kj}}}|F_{\{i,j\}}\right]-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i,j\}}\right]\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{i,j\}}\right]$
$\hskip 28.45274pt-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i\}}\right]\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{j\}}\right]+\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i,j\}}\right]\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{i,j\}}\right]$
$={\operatorname{Cov}}\left(\frac{N}{\hat{N}_{Y_{ki}}},\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{i,j\}}\right)-\left(\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i\}}\right]-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i,j\}}\right]\right)\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{j\}}\right]$
$\hskip 28.45274pt-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i,j\}}\right]\left(\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{j\}}\right]-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{i,j\}}\right]\right)$
$\leq\sqrt{{\operatorname{Var}\,}\left(\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i,j\}}\right){\operatorname{Var}\,}\left(\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{i,j\}}\right)}$
$\hskip 28.45274pt-\left(\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i\}}\right]-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i,j\}}\right]\right)\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{j\}}\right]$
$\hskip 28.45274pt-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i,j\}}\right]\left(\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{j\}}\right]-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{i,j\}}\right]\right)$
$\leq O\left(\frac{\log M}{M}\right)$ (145)
where (144) uses Lemma 5, and (145) uses Lemma 6.
To show (130), we observe for $i,j\in\mathcal{J}_{1}$ that
$C_{ij}(T_{i},T_{j})$ $=\mathbb{E}[v_{i}v_{j}|T_{i},T_{j}]-\mathbb{E}[v_{i}|T_{i}]\mathbb{E}[v_{j}|T_{j}]\frac{P(T_{i})P(T_{j})}{P(T_{i},T_{j})}$
$\leq\frac{1}{\rho^{2}}\mathbb{E}\left[\frac{N^{2}}{\hat{N}_{Y_{ki}}\hat{N}_{Y_{kj}}}|F_{\{i,j\}}\right]-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i\}}\right]\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{j\}}\right]\cdot\frac{1}{\rho}$ (146)
$\leq\frac{1}{\rho^{2}}\left(\mathbb{E}\left[\frac{N^{2}}{(\hat{N}_{Y_{ki}})^{2}}|F_{\{i,j\}}\right]\cdot\mathbb{E}\left[\frac{N^{2}}{(\hat{N}_{Y_{kj}})^{2}}|F_{\{i,j\}}\right]\right)^{1/2}+$
$\hskip 28.45274pt\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i\}}\right]\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{j\}}\right]\cdot\frac{1}{\rho}$
$=O\left(1\right)+O\left(\frac{\log M}{M}\right)$ (147)
where (146) uses Lemma 5 and (55), and (147) uses Lemma 6 and the identity
$\mathbb{E}\left[\frac{N^{2}}{(\hat{N}_{Y_{ki}})^{2}}|F_{\{i,j\}}\right]$ $={\operatorname{Var}\,}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i,j\}}\right]+\left(\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i,j\}}\right]\right)^{2}$
$=O\left(\frac{1}{M}\right)+O(1)$
which holds by Lemma 6 as well. ∎
Proof of Lemma 3.
Recall that $v_{i}$ and $C_{ij}$ are respectively given by (34) and (128). If $i,j\in{\mathcal{I}}$ and $P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})=0$, then
$\mathbb{E}[v_{i}v_{j}|T_{i},T_{j}]$ $=\mathbb{E}\left[\frac{1\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}}{P(\bar{Z}_{i}=Y_{ki}|T_{i})P(\bar{Z}_{j}=Y_{kj}|T_{j})}\cdot\frac{N^{2}}{\hat{N}_{Y_{ki}}\hat{N}_{Y_{kj}}}|T_{i},T_{j}\right]$
$=0$ (148)
where (148) uses $P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})=0$. It follows that
$C_{ij}(T_{i},T_{j})$ $=\mathbb{E}\left[v_{i}v_{j}|T_{i},T_{j}\right]-\mathbb{E}[v_{i}|T_{i}]\cdot\mathbb{E}[v_{j}|T_{j}]\cdot\frac{P(T_{i})P(T_{j})}{P(T_{i},T_{j})}$
$=0-\mathbb{E}[v_{i}|T_{i}]\cdot\mathbb{E}[v_{j}|T_{j}]\cdot\frac{P(T_{i})P(T_{j})}{P(T_{i},T_{j})}$ (149)
$\leq 0$ (150)
where (149) holds by (148), and (150) holds because $v_{i}\geq 0$ always, which implies that $\mathbb{E}[v_{i}|T_{i}]\geq 0$. ∎
Proof of Lemma 4.
Let $\mathcal{T}_{i}$ denote the support of $T_{i}(X)$.
$\sum_{t_{i}\in\mathcal{T}_{i}}\sum_{t_{j}\in\mathcal{T}_{j}}P(T_{i}=t_{i},T_{j}=t_{j})C_{ij}(t_{i},t_{j})$
$=\sum_{t_{i}\in\mathcal{T}_{i}}\sum_{t_{j}\in\mathcal{T}_{j}}P(T_{i}=t_{i},T_{j}=t_{j})\bigg\{\mathbb{E}\left[v_{i}v_{j}|T_{i}=t_{i},T_{j}=t_{j}\right]$-
$\hskip 28.45274pt\mathbb{E}[v_{i}|T_{i}=t_{i}]\mathbb{E}[v_{j}|T_{j}=t_{j}]\cdot\frac{P(T_{i}=t_{i})P(T_{j}=t_{j})}{P(T_{i}=t_{i},T_{j}=t_{j})}\bigg\}$ (151)
$=\mathbb{E}[v_{i}v_{j}]-\sum_{t_{i}\in\mathcal{T}_{i},t_{j}\in\mathcal{T}_{j}}1\{P(T_{i}=t_{i},T_{j}=t_{j})\neq 0\}\cdot$
$\hskip 28.45274pt\mathbb{E}[v_{i}|T_{i}=t_{j}]\mathbb{E}[v_{j}|T_{j}=t_{j}]P(T_{i}=t_{i})P(T_{j}=t_{j})$ (152)
$\geq\mathbb{E}[v_{i}v_{j}]-\sum_{t_{i}\in\mathcal{T}_{i},t_{j}\in\mathcal{T}_{j}}\mathbb{E}[v_{i}|T_{i}=t_{i}]\mathbb{E}[v_{j}|T_{j}=t_{j}]P(T_{i}=t_{i})P(T_{j}=t_{j})$ (153)
$=\mathbb{E}[v_{i}v_{j}]-\mathbb{E}[v_{i}]\cdot\mathbb{E}[v_{j}]={\operatorname{Cov}}(v_{i},v_{j})$
where (151) holds by definition of $C_{ij}$ in (128), and (152) holds because
$\sum_{t_{i}\in\mathcal{T}_{i}}\sum_{t_{j}\in\mathcal{T}_{j}}P(T_{i}=t_{i},T_{j}=t_{j})\mathbb{E}\left[v_{i}v_{j}|T_{i}=t_{i},T_{j}=t_{j}\right]=\mathbb{E}[v_{i}v_{j}],$
and (153) holds because
$\mathbb{E}[v_{i}|T_{i}=t_{i}]\mathbb{E}[v_{j}|T_{j}=t_{j}]P(T_{i}=t_{i})P(T_{j}=t_{j})\geq 0$
since $v_{i}\geq 0$ always. ∎
Proof of Lemma 5.
Recall that $v_{i}$ is given by (34). To show (131), we observe that
$\mathbb{E}\left[v_{i}|T_{i}\right]$ $=\mathbb{E}\left[\frac{1\{\bar{Z}_{i}=Y_{ki}\}N}{P(\bar{Z}_{i}=Y_{ki}|T_{i})\hat{N}_{Y_{ki}}}|T_{i}\right]$
$=P(\bar{Z}_{i}=Y_{ki}|T_{i})\mathbb{E}\left[\frac{N}{P(\bar{Z}_{i}=Y_{ki}|T_{i})\hat{N}_{Y_{ki}}}|T_{i},\bar{Z}_{i}=Y_{ki}\right]$
$=\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|T_{i},\bar{Z}_{i}=Y_{ki}\right].$
To show (132), we observe that if $i,j\in{\mathcal{I}}$ and $|\Gamma_{i}\cap\Gamma_{j}|=0$ then
$\mathbb{E}\left[v_{i}v_{j}|T_{i}T_{j}\right]$ $=\mathbb{E}\left[\frac{1\{\bar{Z}_{i}=Y_{ki}\}\cdot 1\{\bar{Z}_{j}=Y_{kj}\}\cdot N^{2}}{P(\bar{Z}_{i}=Y_{ki}|T_{i})P(\bar{Z}_{j}=Y_{kj}|T_{j})\hat{N}_{Y_{ki}}\hat{N}_{Y_{kj}}}|T_{i},T_{j}\right]$
$=P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})\cdot$
$\hskip 28.45274pt\mathbb{E}\left[\frac{N^{2}}{P(\bar{Z}_{i}=Y_{ki}|T_{i})P(\bar{Z}_{j}=Y_{kj}|T_{j})\hat{N}_{Y_{ki}}\hat{N}_{Y_{kj}}}|T_{i},T_{j},\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\right]$
$=\mathbb{E}\left[\frac{N^{2}}{\hat{N}_{Y_{ki}}\hat{N}_{Y_{kj}}}|T_{i},T_{j},\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\right]$
where we have used that the events $\{\bar{Z}_{i}=Y_{ki}\}$ and $\{\bar{Z}_{j}=Y_{kj}\}$ are independent conditioned on $T_{i}$ and $T_{j}$ when $|\Gamma_{i}\cap\Gamma_{i}|=0$.
To show (133), we observe that if $i,j\in{\mathcal{I}}$ and $|\Gamma_{i}\cap\Gamma_{j}|\neq 0$, then
$\mathbb{E}\left[v_{i}v_{j}|T_{i}T_{j}\right]$ $=\mathbb{E}\left[\frac{1\{\bar{Z}_{i}=Y_{ki}\}\cdot 1\{\bar{Z}_{j}=Y_{kj}\}\cdot N^{2}}{P(\bar{Z}_{i}=Y_{ki}|T_{i})P(\bar{Z}_{j}=Y_{kj}|T_{j})\hat{N}_{Y_{ki}}\hat{N}_{Y_{kj}}}|T_{i},T_{j}\right]$
$=P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})\cdot$
$\hskip 28.45274pt\mathbb{E}\left[\frac{N^{2}}{P(\bar{Z}_{i}=Y_{ki}|T_{i})P(\bar{Z}_{j}=Y_{kj}|T_{j})\hat{N}_{Y_{ki}}\hat{N}_{Y_{kj}}}|T_{i},T_{j},\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\right]$
$\leq\frac{1}{\rho^{2}}\mathbb{E}\left[\frac{N^{2}}{\hat{N}_{Y_{ki}}\hat{N}_{Y_{kj}}}|T_{i},T_{j},\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\right]$
where the final inequality uses (54). ∎
Proof of Lemma 6.
Given a set $S\subseteq[N]$, let $\Gamma_{S}$ denote the set
$\Gamma_{S}=\bigcup_{i\in S}\Gamma_{i}$
and let $Q$ and $\bar{S}$ denote the sets
$Q$ $=\{j:|\Gamma_{j}\cap\Gamma_{S}|=0\},$ $\bar{S}$ $=[N]/Q$
and given $Y_{ki},Q$, and $\bar{S}$, let $\hat{N}_{Q}$ and $\hat{N}_{\bar{S}}$ denote
$\hat{N}_{Q}$ $=\sum_{i\in Q}\frac{1\{\bar{Z}_{j}=Y_{ki}\}}{P\left(\bar{Z}_{j}=Y_{ki}|T_{j}\right)}$ $\hat{N}_{\bar{S}}$ $=\sum_{i\in\bar{S}}\frac{1\{\bar{Z}_{j}=Y_{ki}\}}{P\left(\bar{Z}_{j}=Y_{ki}|T_{j}\right)}$ (154)
so that $\hat{N}_{Y_{ki}}$ satisfies
$\hat{N}_{Y_{ki}}=\hat{N}_{Q}+\hat{N}_{\bar{S}}.$
We observe that $\hat{N}_{Q}$ is independent of $X_{\Gamma_{S}}$, and hence that
$\mathbb{E}\left[\hat{N}_{Q}\ \Big|\ X_{\Gamma_{S}}\right]=N_{Q}$
where $N_{Q}=N-|\bar{S}|$.
From (138), it can be seen that changing $X_{G_{m}}$ for any $m\in[M]$ changes at most $\frac{dCN}{M}$ entries in $Q$, and hence changes $\hat{N}_{Q}$ by at most $\frac{dCN}{M\rho}$, where we have used (54). As a result, by McDiarmid’s inequality it holds for $C_{1}=dC/\rho$ that
$P\left(\left|\hat{N}_{Q}-N_{Q}\right|\geq\epsilon\ \Big|\ X_{\Gamma_{S}}\right)\leq\exp\left(-\frac{\epsilon^{2}M}{C_{1}^{2}N^{2}}\right)$ (155)
Thus for $\epsilon_{B}=\frac{C_{1}N}{\sqrt{M}}\sqrt{(2\delta^{-1}+1)\log M}$ and the indicator
$B=1\left\{\left|\hat{N}_{Q}-N_{Q}\right|\geq\epsilon_{B}\right\},$
it follows from (155) that
$P(B=1\ |\ X_{\Gamma_{S}})$ $\leq\frac{1}{M^{\frac{2}{\delta}+1}}$
$\leq\frac{1}{N^{2+\delta}}$ (156)
where the last inequality uses that $|G_{m}|\leq N^{1-\delta}$, which implies $M\geq N^{\delta}$.
For $i,j\in{\mathcal{I}}$, we let $F_{\{i\}}$ and $F_{\{i,j\}}$ denote sigma-fields such that
$\mathbb{E}[\ \cdot\ |F_{\{i\}}]$ $=\mathbb{E}[\ \cdot\ |T_{i},\bar{Z}_{i}=Y_{ki}]$
$\mathbb{E}[\ \cdot\ |F_{\{i,j\}}]$ $=\mathbb{E}[\ \cdot\ |T_{i},T_{j},\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}]$
as previously used in the proof of Lemma 2. Given $S$ equal to either $\{i\}$ or $\{i,j\}$, we can induce $F_{S}$, $Q$, and $\bar{S}$, and observe that
$\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{S}\right]$ $=\mathbb{E}\left[1+\frac{N-\hat{N}_{Y_{ki}}}{N}+\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S}\right]$ (157)
$\leq 1+\left|\mathbb{E}\left[\frac{|\bar{S}|-\hat{N}_{\bar{S}}}{N}|F_{S}\right]\right|+\left|\mathbb{E}\left[\frac{N_{Q}-\hat{N}_{Q}}{N}|F_{S}\right]\right|+$
$\hskip 28.45274pt\left|\mathbb{E}\left[\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S}\right]\right|$ (158)
$\leq 1+\frac{|\bar{S}|}{N\rho}+0+\left|\mathbb{E}\left[\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S},B=0\right]\right|+$
$\hskip 28.45274ptP(B=1|F_{S})\cdot\left|\mathbb{E}\left[\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S},B=1\right]\right|$ (159)
$\leq 1+\frac{|\bar{S}|}{N\rho}+\frac{(|\bar{S}|/\rho+\epsilon_{B})^{2}}{N(N_{Q}-\epsilon_{B})}+\frac{1}{N^{2+\delta}}\cdot\left(\frac{N}{\rho}\right)^{2}\cdot\frac{1}{N}$ (160)
$=1+O\left(\frac{1}{M}\right)+O\left(\frac{\frac{N^{2}}{M^{2}}+\frac{N^{2}}{M}\log M}{N^{2}}\right)+O\left(\frac{1}{N^{1+\delta}}\right)$ (161)
$=1+O\left(\frac{\log M}{M}\right)$ (162)
where (157) holds by the identity $\frac{a}{b}=1+\frac{a-b}{a}+\frac{(a-b)^{2}}{ab}$; (158) decomposes $N-\hat{N}_{Y_{ki}}=|\bar{S}|-\hat{N}_{\bar{S}}+N_{Q}-\hat{N}_{Q}$; (159) uses the bound
$\left|\mathbb{E}\left[\frac{|\bar{S}|-\hat{N}_{|\bar{S}|}}{N}|F_{S}\right]\right|\leq\frac{|\bar{S}|}{N\rho},$
which holds because $0\leq\hat{N}_{\bar{S}}\leq\frac{|\bar{S}|}{\rho}$, and uses
$\mathbb{E}\left[\frac{N_{Q}-\hat{N}_{Q}}{N}|F_{S}\right]=0,$
which holds because $\mathbb{E}[\hat{N}_{Q}|F_{S}]=N_{Q}$; (160) uses $P(B=1|F_{S})\leq 1/N^{2+\delta}$ which holds by (156), and also uses
$\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}\leq\begin{cases}\displaystyle{\frac{(|\bar{S}|/\rho+\epsilon_{B})^{2}}{N(N_{Q}-\epsilon_{B})}}&\textup{ if }B=0\\ \\ \displaystyle{\left(\frac{N}{\rho}\right)^{2}\cdot\frac{1}{N}}&\textup{ if }B=1\end{cases}$ (163)
which will be proven after the main result; (161) holds because $|\bar{S}|\leq\frac{2CN}{M}$ since $|S|\leq 2$, and also by definition of $\epsilon_{B}$, and $N_{Q}\geq N-\frac{2CN}{M}$. Letting $S=\{i\}$ implies (134), while letting $S=\{i,j\}$ implies (135).
To show (136), we let $S=\{i,j\}$ (inducing $F_{S}$, $Q$, and $\bar{S}$), and observe that
${\operatorname{Var}\,}\left(\frac{N}{\hat{N}_{Y_{ki}}}|F_{S}\right)$ $={\operatorname{Var}\,}\left[1+\frac{|\bar{S}|-\hat{N}_{\bar{S}}}{N}+\frac{N_{Q}-\hat{N}_{Q}}{N}+\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S}\right]$
$=5{\operatorname{Var}\,}\left[\frac{|\bar{S}|-\hat{N}_{\bar{S}}}{N}|F_{S}\right]+5{\operatorname{Var}\,}\left[\frac{N_{Q}-\hat{N}_{Q}}{N}|F_{S}\right]+5{\operatorname{Var}\,}\left[\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S}\right]$ (164)
$\leq\left(\frac{|\bar{S}|}{N\rho}\right)^{2}+\frac{Cd}{\rho^{2}M}+\mathbb{E}\left[{\operatorname{Var}\,}\left(\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S},B\right)|F_{S}\right]+$
$\hskip 28.45274pt{\operatorname{Var}\,}\left(\mathbb{E}\left[\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S},B\right]|F_{S}\right)$ (165)
$\leq\left(\frac{|\bar{S}|}{N\rho}\right)^{2}+\frac{Cd}{\rho^{2}M}+\left(\frac{(|\bar{S}|/\rho+\epsilon_{B})^{2}}{N(N_{Q}-\epsilon_{B})}\right)^{2}+\left(\frac{N^{2}}{\rho^{2}N}\right)^{2}\cdot\frac{1}{N^{2+\delta}}$ (166)
$=O\left(\frac{1}{M^{2}}\right)+O\left(\frac{1}{M}\right)+O\left(\left[\frac{\log M}{M}\right]^{2}\right)+O\left(\frac{1}{M}\right)$ (167)
$=O\left(\frac{1}{M}\right)$ (168)
where (164) follows by the identity
${\operatorname{Var}\,}\sum_{i=1}^{3}a_{i}$ $=\sum_{i=1}^{3}\sum_{j=1}^{3}{\operatorname{Cov}}(a_{i},a_{j})$
$\leq\sum_{i=1}^{3}{\operatorname{Var}\,}a_{i}+\sum_{i=1}^{3}\sum_{j:j\neq i}({\operatorname{Var}\,}a_{i}+{\operatorname{Var}\,}a_{j}),$
(165) uses the bounds
${\operatorname{Var}\,}\left[\frac{|\bar{S}|-\hat{N}_{\bar{S}}}{N}|F_{S}\right]$ $\leq\left(\frac{|\bar{S}|}{N\rho}\right)^{2}$ (169)
${\operatorname{Var}\,}\left[\frac{N_{Q}-\hat{N}_{Q}}{N}|F_{S}\right]$ $\leq\frac{CN^{2}d^{2}}{\rho^{2}MN^{2}}=\frac{Cd^{2}}{\rho^{2}M},$ (170)
which will be proven after the main result, and also uses the law of total variance to decompose ${\operatorname{Var}\,}\left[\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}\right]$; (166) uses the bounds
$\mathbb{E}\left[{\operatorname{Var}\,}\left(\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S},B\right)|F_{S}\right]$ $\leq\left(\frac{(|\bar{S}|/\rho+\epsilon_{B})^{2}}{N(N_{Q}-\epsilon_{B})}\right)^{2}+\frac{1}{N^{2+\delta}}\left(\frac{N^{2}}{\rho N}\right)^{2}$ (171)
${\operatorname{Var}\,}\left(\mathbb{E}\left[\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S},B\right]|F_{S}\right)$ $\leq\frac{1}{N^{2+\delta}}\left(\frac{N^{2}}{\rho N}\right)^{2}$ (172)
which will be proven shortly; (167) holds because $|\bar{S}|\leq\frac{2CM}{N}$, by definition of $\epsilon_{B}$, and $N_{Q}\geq N-\frac{2CM}{N}$.
It remains to show (163), (169) – (170), and (171) – (172):
  1. 1.
    To show (163), we let $\textup{card}(\bar{S})$ denote $|\bar{S}|$, and observe that if $B=0$ then it holds that
    $\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}$ $=\frac{(|\bar{S}|-\hat{N}_{\bar{S}}+N_{Q}-\hat{N}_{Q})^{2}}{N(\hat{N}_{\bar{S}}+\hat{N}_{Q})}$
    $\leq\frac{(|\textup{card}(\bar{S})-\hat{N}_{\bar{S}}|+|N_{Q}-\hat{N}_{Q}|)^{2}}{N(\hat{N}_{\bar{S}}+\hat{N}_{Q})}$
    $\leq\frac{(|\bar{S}|/\rho+\epsilon_{B})^{2}}{N(N_{Q}-\epsilon_{B})},$ (173)
    where (173) uses that $|\hat{N}_{Q}-N_{Q}|<\epsilon_{B}$ when $B=0$, and also that $\hat{N}_{\bar{S}}\geq 0$ to lower bound the denominator, and also that $\hat{N}_{\bar{S}}\leq|\bar{S}|/\rho$ which implies $|\textup{card}(\bar{S})-\hat{N}_{\bar{S}}|\leq|\bar{S}|/\rho$.
    On the other hand, if $B=1$, it holds conditional on $F_{S}$
    $\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}$ $\leq\left(\frac{N}{\rho}\right)^{2}\cdot\frac{1}{N}$ (174)
    where the numerator uses $\hat{N}_{Y_{ki}}\leq N/\rho$ and the denominator uses that $\hat{N}_{Y_{ki}}\geq 1$ conditional on $F_{S}$. Combining (173) and (174) proves (163).
  2. 2.
    To show (169) we observe that $0\leq\hat{N}_{\bar{S}}\leq\frac{|\bar{S}|}{\rho}$.
  3. 3.
    To show (170) we observe that $\hat{N}_{Q}/N$ given by (154) is a sum of $\leq N$ terms whoses magnitudes are bounded by $(N\rho)^{-1}$ and whose dependency graph has at most $N\cdot\frac{d^{2}CN}{M}$ edges, where we have used (137).
  4. 4.
    To show (171), we observe that
    $\mathbb{E}\left[{\operatorname{Var}\,}\left(\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S},B\right)|F_{S}\right]$
    $=P(B=0|F_{S}){\operatorname{Var}\,}\left(\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S},B=0\right)+$
    $\hskip 28.45274ptP(B=1|F_{S}){\operatorname{Var}\,}\left(\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S},B=1\right)$
    $\leq\left(\frac{(|\bar{S}|/\rho+\epsilon_{B})^{2}}{N(N_{Q}-\epsilon_{B})}\right)^{2}+\frac{1}{N^{2+\delta}}\left(\frac{N^{2}}{\rho N}\right)^{2}$ (175)
    where (175) holds by (173), (174), and (156).
  5. 5.
    To show (172), we observe that
    ${\operatorname{Var}\,}\left(\mathbb{E}\left[\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S},B\right]|F_{S}\right)$
    $\leq P(B=1|F_{S})P(B=0|F_{S})\cdot\left(2\max\left\{\frac{(|\bar{S}|/\rho+\epsilon_{B})^{2}}{N(N_{Q}-\epsilon_{B})}\,,\,\frac{N}{\rho}\right\}\right)^{2}$
    $\leq\frac{1}{N^{2+\delta}}\left(\frac{N}{\rho}\right)^{2}$
    which uses (173), (174), (156), and that the variance of a+(ab)Ua+(a-b)U, where $U\sim\textup{Bernoulli}(p)$ and $a,b$ are scalars, equals $p(1-p)(a-b)^{2}$
This shows (134), (135), and (136), as required by the Lemma.
Proof of Lemma 7.
To prove (137), we observe that for any $i\in[N]$, it holds that
$\sum_{j=1}^{N}1\{|\Gamma_{i}\cup\Gamma_{j}|\neq 0\}$ $=\sum_{j=1}^{N}1\{|\eta_{g(i)}\cup\eta_{g(j)}|\neq 0\}$ (176)
$=\sum_{j=1}^{N}\sum_{m=1}^{M}1\{g(j)=m\}\cdot 1\{|\eta_{g(i)}\cup\eta_{m}|\neq 0\}$ (177)
$=\sum_{m=1}^{M}1\{|\eta_{g(i)}\cup\eta_{m}|\neq 0\}\sum_{j=1}^{N}1\{g(j)=m\}$ (178)
$\leq\sum_{m=1}^{M}1\{|\eta_{g(i)}\cup\eta_{m}|\neq 0\}\cdot\frac{CN}{M}$ (179)
$\leq d^{2}\frac{CN}{M}$ (180)
where (176) holds by (53), (177) and (178) are algebraic manipulations, (179) holds by (50), and (180) holds because
$\sum_{m=1}^{M}1\{|\eta_{g(i)}\cap\eta_{m}|\neq 0\}\leq d^{2}$ (181)
which holds because a group neighborhood $\eta_{g(i)}$ contains at most $d$ groups by (51), each of which are members of at most $d$ group neighborhoods by (52).
To prove (138), we observe that for any $m\in[M]$, it holds that
$\sum_{i=1}^{N}1\{G_{m}\subseteq\Gamma_{i}\}$ $=\sum_{i=1}^{N}1\{m\in\eta_{g(i)}\}$ (182)
$=\sum_{i=1}^{N}\sum_{\ell=1}^{M}1\{g(i)=\ell\}\cdot 1\{m\in\eta_{\ell}\}$ (183)
$=\sum_{\ell=1}^{M}1\{m\in\eta_{\ell}\}\cdot\sum_{i=1}^{N}1\{g(i)=\ell\}$ (184)
$\leq\sum_{\ell=1}^{N}1\{m\in\eta_{\ell}\}\cdot\frac{CN}{M}$ (185)
$\leq d\frac{CN}{M}$ (186)
where (182) holds by (53), (183) and (184) are algebraic manipulations, (185) holds by (50), and (186) holds by (51). ∎
D.5 Proof of Theorem 8
To prove Theorem 8, we will use Lemmas 8-10, and also Theorem 9 which is a local dependence central limit theorem.
Lemma 8 gives a coverage result for $\hat{\tau}_{k}$.
Lemma 8.
Let Assumption 2 hold, and let $\widetilde{V}_{k}(\phi)$ for $\phi\in\{0,1\}^{N}$ and $k=1,2$ be given by (68). It holds for $k=1,2$ that
$P\left(\hat{\tau}_{k}^{\textup{Haj}}-\tau\geq z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{k}}\right)\leq\frac{\alpha}{2}+o(1)$
Lemmas 9 and 10 are used to prove Lemma 8. Lemma 9 gives a central limit theorem for a linearized version of $\hat{\tau}_{k}^{\textup{Haj}}$. Lemma 10 is a corollary of Fatou’s lemma, and will be used to show that $\hat{V}_{k}$ is conservative for the variance of the linearized version of $\hat{\tau}_{k}^{\textup{Haj}}$. It appears as Exercise 3.2.4 in (Durrett, , 2019, 5th edition).
Lemma 9.
Let Assumption 2 hold. For $k\in[2]$, let $\mu_{1}$ and $\mu_{0}$ denote
$\mu_{1}$ $=\sum_{i\in{\mathcal{I}}}\mathbb{E}[\tilde{f}_{ki}(T_{i})],$ $\mu_{0}$ $=\sum_{i\in{\mathcal{I}}}\mathbb{E}[1-\tilde{f}_{ki}(T_{i})].$
where $\tilde{f}_{ki}$ is given by (76). Let $h_{i}$ for $i\in[N]$ denote
$h_{i}=\frac{1\{\bar{Z}_{i}=1\}}{P(\bar{Z}_{i}=1|T_{i})}Y_{ki}\phi_{i}^{*}-\mu_{1}+\frac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=0|T_{i})}(1-Y_{ki})\phi_{i}^{*}-\mu_{0}$
where $\phi^{*}$ is given by (102). Let $\sigma^{2}={\operatorname{Var}\,}\sum_{i}h_{i}$, and suppose that $\sigma^{2}\geq N^{2}/M^{4/3-\epsilon}$ for any $\epsilon>0$. Then $\sum_{i}h_{i}/\sigma$ converges to a standard normal.
Lemma 10.
(Durrett, , 2019, Exercise 3.2.4) Let $X_{n}$ denote a sequence of random variables that converges in distribution to $X_{\infty}$, and let $g$ denote a non-negative continuous function. Then $\liminf\limits_{n\rightarrow\infty}\mathbb{E}g(X_{n})\geq\mathbb{E}g(X_{\infty})$
Theorem 9 is taken from Ross, (2011), and gives a bounded dependence central limit theorem. It will be used to prove the asymptotic normality result of Lemma 9.
Theorem 9.
(Ross, , 2011, Thm 3.6) Let $U_{1},\ldots,U_{n}$ be random variables such that $\mathbb{E}[U_{i}^{4}]\leq\infty$, $\mathbb{E}U_{i}=0$, with variance denoted by $\sigma^{2}={\operatorname{Var}\,}\sum_{i=1}^{N}U_{i}$, with dependency neighborhoods $N_{1},\ldots,N_{n}$, which are subsets of $[n]$ (which need not be disjoint) such that $U_{j}$ and $U_{i}$ are independent if $U_{j}\notin N_{i}$. Let $D=\max_{i}|N_{i}|$. Let $W=\sum_{i=1}^{N}U_{i}/\sigma$. Then for $Z$ denoting a $N(0,1)$ random variable, and $d_{W}$ denoting Wasserstein metric, it holds that
$d_{W}(W,Z)\leq\frac{D^{2}}{\sigma^{3}}\sum_{i=1}^{n}\mathbb{E}\left[|U_{i}|^{3}\right]+\frac{\sqrt{28}D^{3/2}}{\sqrt{\pi}\cdot\sigma^{2}}\sqrt{\sum_{i=1}^{n}\mathbb{E}\left[U_{i}^{4}\right]}$
Proof of Theorem 8
Proof of Theorem 8.
It holds that
$LB_{1-\alpha}-\tau$ $=\max\left\{\min_{\phi\in\{0,1\}}\hat{\tau}_{1}(\phi)-z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{1}(\phi)}\ ,\ \min_{\phi\in\{0,1\}}\hat{\tau}_{2}(\phi)-z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{2}(\phi)}\right\}-\tau$
$\leq\max\left\{\hat{\tau}_{1}-z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{1}}\ ,\ \hat{\tau}_{2}-z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{2}}\right\}-\tau$
$=\max\left\{\hat{\tau}_{1}-z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{1}}-\tau\ ,\ \hat{\tau}_{2}-z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{2}}-\tau\right\}$ (187)
Let $\mathcal{E}_{k}$ denote the event
$\mathcal{E}_{k}=\left\{\hat{\tau}_{k}-z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{k}}-\tau\geq 0\right\}.$
By (187), it follows that loss of coverage occurs if either $\mathcal{E}_{1}$ or $\mathcal{E}_{2}$ occurs. By Lemma 8, the probability of $\mathcal{E}_{k}$ is bounded by $\alpha/2+o(1)$, and hence by union bound it holds that
$P(\mathcal{E}_{1}\cup\mathcal{E}_{2})\leq\alpha+o(1)$ (188)
Combining (187) and (188) yields that
$P(LB_{1-\alpha}-\tau>0)\leq\alpha+o(1)$ (189)
proving that $LB_{1-\alpha}\leq\tau$ with probability converging to at least $1-\alpha$. ∎
Proof of Lemmas 8 and 9
Proof of Lemma 8.
Let $h_{i}$ be defined as in Lemma 9, and let $\sigma^{2}={\operatorname{Var}\,}\sum_{i}h_{i}$. Divide the sequence of experiments into two subsequences, depending on whether $\sigma^{2}\geq N^{2}/M^{4/3-\epsilon}$.
For the subsequence that satisfies $\sigma^{2}\geq N^{2}/M^{4/3-\epsilon}$, it holds by linearization arguments made in the proof of Theorem 6 that
$\frac{\hat{\tau}_{k}^{{\textup{Haj}}}-\tau}{\sigma}$ $=\frac{\nabla h(\mathbb{E}\xi)^{T}(\xi-\mathbb{E}\xi)}{\sigma}+O_{P}\left(\frac{1}{\sigma}\right)$ (190)
$=\frac{\sum_{i=1}^{N}h_{i}}{\sigma}+O_{P}\left(\frac{1}{\sigma}\right)$ (191)
where (190) holds by (123) and (127), while (191) holds by (112)-(122) and (124). By Lemma 9 and (191) it follows that
$\frac{\hat{\tau}_{k}^{{\textup{Haj}}}-\tau}{\sigma}\rightarrow N(0,1)$
from which it follows by Lemma 10 that
$\liminf\limits_{N\rightarrow\infty}{\operatorname{Var}\,}\left[\frac{\hat{\tau}_{k}^{\textup{Haj}}-\tau}{\sigma}\right]\geq 1$ (192)
Since by Theorem 7 it holds that
$\frac{\hat{V}_{k}}{\sigma^{2}}$ $=\frac{\mathbb{E}\hat{V}_{k}}{\sigma^{2}}+\frac{1}{\sigma^{2}}O_{P}\left(\frac{N^{2}\log M}{M^{3/2}}\right)$
$\geq{\operatorname{Var}\,}\left(\frac{\hat{\tau}_{k}^{\textup{Haj}}}{\sigma}\right)+o_{P}\left(1\right)$ (193)
Combining (192) and (193) yields
$\liminf\limits_{N\rightarrow\infty}\frac{\hat{V}_{k}}{\sigma^{2}}\geq 1$
which implies that
$P\left(\hat{\tau}_{k}^{\textup{Haj}}-\tau\geq z_{1-\frac{\alpha}{2}}\sqrt{\hat{V}_{k}}\right)\leq\frac{\alpha}{2}+o(1)$ (194)
for the subsequence of experiments that satisfies $\sigma^{2}\geq N^{2}/M^{4/3-\epsilon}$.
For the other subsequence, Chebychev inequality implies that
$P\left(\left|\sum_{i}h_{i}\right|\geq\frac{N}{M^{2/3-\epsilon}\sqrt{\alpha/2}}\right)\leq\alpha/2$
where we have used that $\frac{N}{M^{2/3-\epsilon}}=\omega(\sigma)$ for $\epsilon>0$. In turn, this implies that
$P\left(\hat{\tau}_{k}-\tau\geq\frac{N}{M^{2/3-\epsilon}\sqrt{\alpha/2}}\right)\leq 1-\alpha+o(1)$ (195)
where we have used that $\hat{\tau}_{k}-\tau=\sum_{i}h_{i}+O_{P}(1)$, which holds by (190). Applying (195) yields
$P\left(\hat{\tau}_{k}-\tau\geq z_{1-\frac{\alpha}{2}}\sqrt{\Delta}\right)\leq\alpha/2+o(1)$ (196)
where $\Delta=N/(M^{2/3-\epsilon}z_{1-\frac{\alpha}{2}}^{2}\cdot\alpha/2)$. Combining (194) and (196) yields that for all experiments, it holds that
$P\left(\hat{\tau}_{k}-\tau\geq z_{1-\frac{\alpha}{2}}\sqrt{\max\left(\hat{V}_{k},\Delta\right)}\right)\leq\alpha/2+o(1),$
As $\widetilde{V}_{k}$ can be seen to equal $\max\left(\hat{V}_{k},\Delta\right)$, this proves the lemma.
Proof of Lemma 9.
Let $U_{m}$ for $m\in[M]$ be given by
$U_{m}=\sum_{i\in G_{m}}h_{i}$
Since $|h_{i}|\leq\frac{1}{\rho}$ and $|G_{m}|\leq CN/M$, it follows that $|U_{m}|\leq\frac{CN}{M\rho}$. Additionally, it can be seen that the dependency neighborhoods of $U_{1},\ldots,U_{M}$ have maximum size $d^{2}$, since each $U_{m}$ depends on $X$ only through $\{X_{G_{\ell}}:\ell\in\eta_{m}\}$. As a result, applying Theorem 9 yields
$d_{W}\left(\frac{\sum_{i}h_{i}}{\sigma}\ ,\ N(0,1)\right)$ $\leq\frac{d^{4}}{\sigma^{3}}\cdot M\left(\frac{CN}{M\rho}\right)^{3}+\frac{\sqrt{28}}{\sqrt{\pi}}\cdot\frac{d^{3}}{\sigma^{2}}\sqrt{M\left(\frac{CN}{M\rho}\right)^{4}}$
$=O\left(\frac{N^{3}}{\sigma^{3}M^{2}}\right)+O\left(\frac{N^{2}}{M^{3/2}\sigma^{2}}\right)$ (197)
$=o(1)$ (198)
where (198) uses the assumption of the lemma that $\sigma^{2}\geq\frac{N}{M^{2/3-\epsilon}}$ for $\epsilon>0$. ∎